### v1

# Standard Gaussian Process is All You Need for High-Dimensional Bayesian Optimization

Abstract
There has been a long-standing and widespread belief that Bayesian Optimization (BO) with standard Gaussian process (GP), referred to as standard BO, is ineffective in high-dimensional optimization problems. This perception may partly stem from the intuition that GPs struggle with high-dimensional inputs for covariance modeling and function estimation. While these concerns seem reasonable, empirical evidence supporting this belief is lacking. In this paper, we systematically investigated BO with standard GP regression across a variety of synthetic and real-world benchmark problems for high-dimensional optimization. Surprisingly, the performance with standard GP consistently ranks among the best, often outperforming existing BO methods specifically designed for high-dimensional optimization by a large margin. Contrary to the stereotype, we found that standard GP can serve as a capable surrogate for learning high-dimensional target functions. Without strong structural assumptions, BO with standard GP not only excels in high-dimensional optimization but also proves robust in accommodating various structures within the target functions. Furthermore, with standard GP, achieving promising optimization performance is possible by only using maximum likelihood estimation, eliminating the need for expensive Markov-Chain Monte Carlo (MCMC) sampling that might be required by more complex surrogate models. We thus advocate for a re-evaluation and in-depth study of the potential of standard BO in addressing high-dimensional problems.
1 Introduction
Many applications require optimizing complex functions, allowing queries only for the function values, possibly with noises, yet without any gradient information. Bayesian Optimization (BO) (Snoek et al., 2012; Mockus, 2012) is a popular approach to address such optimization challenges. BO typically employs Gaussian process (GP) (Williams and Rasmussen, 2006) as a probabilistic surrogate. It iteratively approximates the target function, integrates the posterior information to optimize an acquisition function for generating new inputs for querying, then updates the GP model with new examples, and concurrently approaches the optimum.
Despite numerous successes, there has been a widespread belief that BO with standard GP regression, referred to as standard BO, is limited to low-dimensional optimization problems (Frazier, 2018; Nayebi et al., 2019; Eriksson and Jankowiak, 2021; Moriconi et al., 2020; Letham et al., 2020; Wang et al., 2016; Li et al., 2016). It is commonly thought that the number of optimization variables should not exceed 15 or 20, as beyond this threshold, BO is prone to failure (Frazier, 2018; Nayebi et al., 2019). This belief may partly stem from the intuition that, as a kernel method, GPs struggle to handle high-dimensional inputs for covariance modeling, rending them incapable of providing an expressive prior for high-dimensional functions.
A continuous line of research is dedicated to developing novel high-dimensional Bayesian Optimization (BO) methods. The fundamental strategy is to impose strong structural assumptions into surrogate modeling so as to avoid directly dealing with high-dimensional inputs. One class of methods assumes a decomposition structure within the functional space (Kandasamy et al., 2016; Rolland et al., 2018; Han et al., 2020; Ziomek and Bou-Ammar, 2023). In other words, the target function can be expressed as the summation of a group of low-dimensional functions, each operating over a small number of variables. Another family of methods assumes that the inputs are intrinsically low-rank. These methods either project the original inputs into a low-dimensional space (Wang et al., 2016; Nayebi et al., 2019; Letham et al., 2020) or use sparse-inducing priors to trim down massive input variables (Eriksson and Jankowiak, 2021). Subsequently, GP surrogates are built with the reduced input dimensions.
While the aforementioned beliefs and concerns may appear intuitive and reasonable, there lacks strong empirical evidence confirming that standard BO is indeed incapable of high-dimensional optimization. To bridge this gap, we systematically investigated BO with standard GP regression across a variety of synthetic and real-world benchmark problems. Surprisingly, the performance of standard BO consistently ranks among the best and often surpasses existing methods by a substantial margin. These methods are specifically designed for high-dimensional optimization This leads us to call for a re-evaluation and in-depth study of the potential of standard BO for high-dimensional optimization. The major discoveries are as follows.
  • Optimization performance. We conducted an investigation using BO with standard GP regression across six synthetic and six real-world high-dimensional optimization tasks. Our selection includes eleven benchmarks widely used in literature, and a new real-world benchmark task on neural network pruning. The number of optimization variables varies from 30 to 388. In comparison with six state-of-the-art high-dimensional BO methods, standard BO consistently achieves the best (or close to the best) optimization performance across all the tasks. In contrast, the competing methods demonstrate strong performance when the target function aligns with their structural assumption. However, in cases of mismatch or when the structure of the target function is agnostic (e.g., in real-world problems), the performance of the competing methods often experiences a significant drop. This underscores that BO with standard GP not only can excel in high-dimensional optimization but also is robust in accommodating various structures within the target functions.
  • Surrogate Learning. We investigated the task of learning high-dimensional functions using standard GP regression. Across twelve benchmark datasets, our findings reveal that the prediction of standard GP aligns well with the ground-truth, mostly performing better than more complex GP models that induce low-rank structures. This observation is in line with the recent results in kernel operator learning (Batlle et al., 2024), where inputs are discretized functions — typically high-dimensional. It confirms that standard GP can serve as a capable surrogate for high-dimensional functions.
  • Computational Efficiency. We discovered that, with standard GP, achieving promising optimization performance is possible by using maximum likelihood estimation. The marginalization of the parameters (e.g., kernel parameters and noise variance) via Markov-Chain Monte Carlo (MCMC) provides only minor benefits while significantly increasing the training cost. In contrast, due to the usage of more complex surrogate models, training for alternative high-dimensional BO methods is often more expensive. Some of these methods heavily rely on extensive MCMC sampling to prevent overfitting, making the training of the surrogate model a computational bottleneck.
2 Standard Bayesian Optimization
Consider maximizing a $d$-dimensional black-box objective function $f:\mathcal{X}\subset\mathbb{R}^{d}\rightarrow\mathbb{R}$. The form of $f(\cdot)$ is unknown. We can only query the function value at any input ${\bf x}\in\mathcal{X}$, possibly with noises, and no gradient information is accessible. We aim to find
${\bf x}^{\dagger}=\operatornamewithlimits{argmax}_{{\bf x}\in\mathcal{X}}f({\bf x}).$ (1)
To achieve this, Bayesian Optimization (BO) employs a probabilistic surrogate model to learn and predict $f(\cdot)$ across the domain $\mathcal{X}$, while also quantifying the uncertainty of the prediction. This information is then integrated to compute an acquisition function, which measures the utility of querying at any new input location given the current function estimate. The utility typically encapsulates an exploration-exploitation trade-off. By maximizing the acquisition function, BO identifies a new input location at which to query, ideally closer to the optimum. Concurrently, the acquired new example is incorporated into the training dataset, and the surrogate model is retrained to improve the accuracy. BO begins by querying at a few (randomly selected) input locations, and trains an initial surrogate model. Then the iterative procedure repeats until BO converges to the optimum or a predefined stopping criterion is met.
The standard BO adopts GP regression (Williams and Rasmussen, 2006) for surrogate modeling. Specifically, a GP prior is placed over the target function,
$f\sim\mathcal{GP}\left(m({\bf x}),\kappa({\bf x},{\bf x}^{\prime})\right)$ (2)
where $m({\bf x})$ is the mean function, which is often set as a constant, and $\kappa(\cdot,\cdot)$ is the covariance function, which is often chosen as a Mercer kernel function. For example, a popular kernel used in BO is the Matérn-5/2 Automatic Relevance Determination (ARD) kernel,
$\kappa({\bf x},{\bf x}^{\prime})=a\left(1+{\sqrt{5}\tau}+\frac{5\tau^{2}}{3}\right)\exp\left(-{\sqrt{5}\tau}\right)$ (3)
where
$\tau=\sqrt{({\bf x}-{\bf x}^{\prime})^{\top}{\rm diag}(\frac{1}{\bm{\eta}})({\bf x}-{\bf x}^{\prime})},$ (4)
$a>0$ is the amplitude, and $\bm{\eta}>{\bf 0}$ are the length-scale parameters. We refer to this kernel as ARD because each input dimension has a distinct length-scale parameter. Alternatively, one can use a single length-scale parameter shared by all the input dimensions.
Suppose we have collected a set of training inputs ${\bf X}=[{\bf x}_{1},\ldots,{\bf x}_{N}]^{\top}$ and (noisy) outputs ${\bf y}=[y_{1},\ldots,y_{N}]^{\top}$. Due to the GP prior, the function values at $\bf X$}, denoted as ${\bf f}=[f({\bf x}_{1}),\ldots,f({\bf x}_{N})]^{\top}$, follow a multi-variate Gaussian distribution, $p({\bf f})=\mathcal{N}({\bf f}|{\bf m},{\bf K})$, where ${\bf m}=[m({\bf x}_{1}),\ldots,m({\bf x}_{N})]^{\top}$, $\bf K$} is the kernel matrix on $\bf X$}, and each ${\bf K}]_{ij}=\kappa({\bf x}_{i},{\bf x}_{j})$. Next, a Gauss noise model is used to model the likelihood of $\bf y$}, namely, $p({\bf y}|{\bf f})=\mathcal{N}({\bf y}|{\bf f},\sigma^{2}{\bf I})$ where $\sigma^{2}$ is the noise variance. One can marginalize $\bf f$} out,
$p({\bf y}|{\bf X})=\mathcal{N}({\bf y}|{\bf m},{\bf K}+\sigma^{2}{\bf I}).$ (5)
To estimate the GP parameters (e.g., $a$, $\bm{\eta}$, and $\sigma^{2}$), one can use MCMC or simply maximize (5) to obtain a point estimate. The predictive distribution (given the GP parameters) has a closed form. At any new input $\bf x$$^{*}$, since $f({\bf x}^{*})$ and $\bf y$} jointly follow a multi-variate Gaussian distribution, the predictive distribution is a conditional Gaussian distribution,
$p\left(f({\bf x}^{*})|{\bf y}\right)=\mathcal{N}\left(f({\bf x}^{*})|\mu({\bf x}^{*}),v({\bf x}^{*})\right),$ (6)
where
$\mu({\bf x}^{*})$ $=m({\bf x}^{*})+\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}({\bf y}-{\bf m})$ (7)
$v({\bf x}^{*})$ $=\kappa({\bf x}^{*},{\bf x}^{*})-\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}\kappa({\bf X},{\bf x}^{*}).$ (8)
In each iteration, given the current GP surrogate, we maximize an acquisition function to identify the next input location at which to query. One commonly used acquisition function is the upper confidence bound (UCB), defined as
$\text{UCB}({\bf x})=\mu({\bf x})+\lambda\sqrt{v({\bf x})},\;\;{\bf x}\in\mathcal{X},$ (9)
where $\lambda$ represents the exploration level. There have been other acquisition functions, such as Expected Improvement (EI) (Jones et al., 1998), Thompson sampling (Russo et al., 2018), among others.
3 High Dimensional Bayesian Optimization
An enduring and widespread belief is that when the input dimension $d$ is high (e.g., a few hundreds), the standard BO is prone to failure. This belief might partly arise from an intuition that commonly used (stationary) kernels, such as in (3), could encounter challenges in dealing with high-dimensional inputs, making GP struggle in capturing the target function. As a result, there is a dedicated line of research aimed at developing novel high-dimensional BO methods. The key idea of these methods is to introduce strong structural assumptions in surrogate modeling to sidestep directly handling the high-dimensional inputs in kernels.
3.1 Structural Assumption in Functional Space
The first class of methods assumes a decomposition structure within the functional space. Specifically, the target function $f$ is modeled as
$f({\bf x})=\sum_{j=1}^{M}f_{j}({\bf x}^{j}),\;\;f_{j}\sim\mathcal{GP}\left(m_{j}({\bf x}^{j}),\kappa_{j}(\cdot,\cdot)\right),$ (10)
where each ${\bf x}^{j}\subset{\bf x}$ is a small group of input variables, and ${\bf x}={\bf x}^{1}\cup\ldots\cup{\bf x}^{M}$. There can be a variety of choices for the number of group $M$ and each variable group ${\bf x}^{j}$. In (Kandasamy et al., 2016), the input $\bf x$} is partitioned into non-overlapped groups. After every a few BO iterations, a new partition is selected from a set of random partitions. The selection is based on the model evidence. In (Rolland et al., 2018), the variable groups can overlap, and are represented as the maximum cliques on a dependency graph. A Gibbs sampling method was developed to learn the structure of the dependency graph from data. Han et al. (2020) proposed Tree-UCB, which restricted the dependency graph to be tree-structured so as to boost the efficiency of structure learning. However, the most recent work (Ziomek and Bou-Ammar, 2023) points out that, learning the group structure through relatively small data can be misleading. A wrongly learned structure can cause queries to be stuck at local optima. The work proposes RDUCB that randomly decomposed $\bf x$} into a tree structure, and surprisingly works better than those learned structures through various methods.
3.2 Structural Assumption in Input Space
Another class of methods assumes a low-rank structure in the input space.
Low-dimensional Embedding. Wang et al. (2016) proposed a method named REMBO, building the GP surrogate in a low-dimensional embedding space, $\mathcal{Z}\subset\mathbb{R}^{d_{\text{emb}}}$ where $d_{\text{emb}}\ll d$. The acquisition function is optimized in the embedding space. When querying the function value, the input is recovered by ${\bf x}={\bf A}{\bf z}$ where ${\bf z}\in\mathcal{Z}$, and $\bf A$} is a random matrix. If the recovered $\bf x$} is out of the domain $\mathcal{X}$, it will be clipped as a boundary point. To avoid clipping, Nayebi et al. (2019) proposed HESBO, which randomly duplicates $\bf z$} (and/or flip the signs) to form $\bf x$}. This way, the optimization of $\bf z$} can respect the same bounding constraints as in $\mathcal{X}$. This essentially impose further constraints on the embedding space. A more flexible solution, ALEBO, was proposed in  (Letham et al., 2020). ALEBO also uses a random matrix $\bf A$} to recover the input $\bf x$}. But when maximizing the acquisition function, ALEBO incorporates a constraint that ${\bf A}{\bf z}$ must be in $\mathcal{X}$, thereby avoiding the clipping issue.
Partial Variable Dependency. Another strategy is to trim down the input variables. The recent start-of-the-art, SaasBO (Eriksson and Jankowiak, 2021), uses a horse-shoe prior (Carvalho et al., 2009) over the length-scale parameter for each variable. As the horse-shoe prior induces strong sparsity, a massive number of variables can be pruned, substantially reducing the input dimension. For effective training with the horse-shoe prior, SaasBO used a No-U-Turn Sampler (NUTS) (Hoffman et al., 2014). While one can also use MAP estimation, the performance is often much inferior.
4 Comprehensive Evaluation
We now examine and analyze the performance of standard BO in high-dimensional optimization. To this end, we conducted a series of systematic investigations.
4.1 Experimental Settings
We employed twelve benchmarks, of which six are synthetic and the other six are real-world problems. For clarity in the discussion, all the tasks aim for function maximization.
Synthetic Benchmarks. We considered four popular synthetic functions: Ackley, Rosenbrock, Hartmann6, and Stybtang. Definitions of these functions are provided in Section A.1 of the Appendix. Each task is represented by $\text{Fun}(d,d^{\prime})$, where $d$ is the input dimension that the BO methods optimize for, and $d^{\prime}$ is the number of effective variables ($d^{\prime}\leq d$). The target function is computed using the first $d^{\prime}$ variables. The benchmarks and the structures within the target functions are summarized in Table 1.
$\text{Fun}(d,d^{\prime})$ Structure
Ackley(300, 150) Partial variable dependency
Rosenbrok(300, 100) Partial variable dependency
Nonoverlap additive decomposition
Hartmann6(300,6) Partial variable dependency
Rosenbrock(100, 100) Nonoverlap additive decomposition
Stybtang(200, 200)
Ackley(150, 150) None
Table 1: Synthetic benchmarks and the structures within the target function: $d$ is the input dimension, and $d^{\prime}$ is the number of effective variables to compute the target function.
Real-World Benchmarks. We employed the following real-world benchmark problems.
  • Mopta08 (124)  (Jones, 2008), a vehicle design problem. The objective is to minimize the mass of a vehicle with respect to 124 design variables, subject to 68 performance constraints. We followed (Eriksson and Jankowiak, 2021) to encode the constraints as a soft penalty, which is added into the objective.
  • SVM (388) (Eriksson and Jankowiak, 2021), a hyper-parameter tuning problem that optimizes 3 regularization parameters and 385 length-scale parameters for a kernel support vector machine regression model.
  • Rover (60), a rover trajectory planning problem from (Wang et al., 2018). The goal is to find an optimal trajectory determined by the location of 30 waypoints in a 2D environment.
  • DNA (180) (Šehić et al., 2022), a hyper-parameter optimization problem that optimizes 180 regularization parameters for weighted lasso regression on an DNA dataset (Mills, 2020). Prior analysis (Šehić et al., 2022) shows that only around 43 variables are relevant to the target function.
  • NAS201 (30)  (Dong and Yang, 2020), a neural architecture search problem on CIFAR-100 dataset.
  • NN-Prune (151). We provided an additional real-world benchmark on neural network pruning (Blalock et al., 2020). We considered ResNet-152 (He et al., 2015), a deep residual network with 151 convolution layers and one linear layer, including more than 60 million parameters in total. We introduced a pruning rate $x_{i}$ for each convolution layer $i$. Given a global pruning ratio set to 0.5, the objective is to optimize the per-layer pruning rates to maximize the prediction accuracy on a test dataset after pruning and fine tuning. The details are given in Section A.2 of the Appendix.
Method Structural Assumption
ALEBO, HESBO Low-dim embedding
SaasBO Partial variable dependency
RDUCB, Tree UCB Additive function decomposition
GP-ARD, GP, TURBO None
Table 2: BO methods and their structural assumption.
Methods. To execute the standard BO, we used GP regression with Matérn-5/2 ARD kernel (see (3)). We also tested with Matérn-5/2 kernel with a single length-scale parameter. We denoted these two choices as GP-ARD and GP, respectively. Throughout the experiments, we used UCB as the acquisition function where the exploration level $\lambda$ was set to 1.5. We used GPyTorch (Gardner et al., 2021) for surrogate training and BOTorch (Balandat et al., 2020) for acquisition function computation.
We compared with six state-of-the-art high-dimensional BO methods, including Tree UCB (Han et al., 2020), RDUCB (Ziomek and Bou-Ammar, 2023), HESBO (Nayebi et al., 2019), ALEBO (Letham et al., 2020), SaasBO (Eriksson and Jankowiak, 2021), and TURBO (Eriksson et al., 2020). The first five have been introduced in Section 3. TURBO is designed for a special scenario where the target function can be extensively evaluated (e.g., tens of thousands evaluations). This is not common in BO applications, as the function evaluation is typically deemed expensive, and one aims to evaluate the target function as few times as possible. TURBO searches for trust-regions in the input domain, learns a set of local GP models over each region, and then uses an implicit bandit strategy to decide which local optimization runs to continue. Table 2 summarizes the structural assumptions made by all the methods. We used the original implementation and default (recommended) settings for the competing methods. See details in Appendix Section A.3.
For each optimization task, we randomly queried the target function for 20 times to generate the initial data. We then ran each method for 400 steps for synthetic tasks and for 300 steps for real-world tasks. We tested TURBO with a single trust region and five trust regions, denoted by TURBO-1 and TURBO-5 respectively. For HESBO and ALEBO, we varied $d_{\text{emb}}\in\{10,20\}$ and denote the choice as HESBO/ALEBO-{10, 20}. For SaasBO, we performed tests using both NUTS and MAP estimation for surrogate training, denoted by SaasBO-NUTS and SaasBO-MAP, respectively. For standard BO, we trained the GP regression by maximizing the marginal likelihood using ADAM optimization (Kingma and Ba, 2014) (learning rate $0.1$ and 500 steps, as suggested in GPyTorch documents). We treated the mean function as a constant and jointly estimated it with the kernel parameters and noise variance.
4.2 Results and Discussion
4.2.1 Optimization Performance
We present the ultimate optimization outcomes for synthetic benchmarks in Fig. 1 and for real-world benchmarks in Fig. 2. These are violin plots that illustrate the average of the obtained maximum function value over ten runs (the middle bar), the best (the top bar), the worst (the bottom bar), and the distribution of the obtained maximum (the shaded region). We can observe that on average, the standard BO with GP-ARD outperforms the competing high-dimensional BO methods in optimization every target function, often by a large margin, except that in the cases of Hartman6(300, 6) and NN-Prune, the performance with GP-ARD is slight worse (but still very close to) the best baseline(s).
It is noteworthy that the performance of the competing methods appears to be sensitive to the structure within the target function. For example, SaasBO imposes a strong inductive bias that only a small number of variables is related to the objective (via the usage of the horse-shoe prior). The target function in Hartman6(300, 6) and DNA aligns well with this assumption. As a result, SaasBO exhibits outstanding performance in both tasks. However, when the objective deviates from this assumption, as seen in cases like Ackley(150, 150), StybTang(200, 200), NN-Prune, and even Rosenbrock(300, 100), the performance of SaasBO experiences a dramatic drop. Another example is RDUCB and Tree-UCB, which assume that the target function can be decomposed as a summation of a group of functions, each over a small number of variables. RDUCB and Tree-UCB perform well in synthetic benchmarks that match this assumption, such as Rosenbrock(100, 100), Rosenbrock(300, 100), and StybTang(200, 200). However, their performance drops significantly in cases that do not align with their assumption, such as Ackley(150, 150) and Ackley(300, 150).
In contrast, GP-ARD exhibits much greater robustness to the structure of the target function. The standard BO with GP-ARD consistently ranks among the best in all tasks, regardless of whether the target function depends only on a subset of variables, can be decomposed, etc. Interestingly, the standard BO shows superior performance even when using a single length-scale parameter for GP, which closely matches GP-ARD across most tasks, except for Ackley(300, 150), GP struggled in finding right locations to query, leading to relative poor performance.
We also examined the runtime performance of each method. We report the maximum function value obtained throughout the optimization process in Fig. 5 and Fig. 6 of the Appendix. It is noticeable that after a few initial steps, the standard BO with GP-ARD (and also with GP in most cases) consistently generated superior queries, finding larger function values. Its curves are mostly above those of the competing methods, indicating a more efficient optimization of the target function. This trend is particularly evident in examples such as Rosenbrock(100, 100), StybTang(200, 200), Mopta08, Rover, SVM, and more.
The collective results demonstrate that the standard GP with ARD kernels is already effective for high-dimensional optimization. In comparison with using more complex surrogate models with additional assumptions, GP-ARD exhibits greater flexibility in adapting to the target function whose intrinsic structures are often agnostic in practice, and provides superior optimization results.
Figure 1: Optimization performance in synthetic benchmarks. The results were summarized from 10 runs.
Figure 2: Optimization performance in real-world problems. The results were summarized from 10 runs.
4.2.2 Surrogate Learning
Next, we explicitly evaluated the performance of the standard GP in learning high-dimensional functions to confirm if it can serve as a reliable surrogate model for high-dimensional BO. We sampled 300 training examples from all the six synthetic benchmark functions and four real-world benchmark tasks, including Mopta08 (124), Rover (60), DNA (180) and SVM (388). We sampled another 100 examples for testing. We compared with the surrogate models used in SaasBO and ALEBO (with $d_{\text{emb}}=10$ and $d_{\text{emb}}=20$). SaasBO was trained by NUTS. We repeated the experiment for five times, and evaluated the average test log-likelihood (LL), mean-square-error (MSE), and the standard deviation.
The results are reported in Table 3 and Appendix Table 5. We can see that in the vast majority of cases, GP-ARD achieves the largest test LL and the smallest MSE. GP-ARD and GP together nearly always outperforms SaasBO, ALEBO-10 and ALEBO-20. Notably, GP-ARD surpasses ALEBO by a substantial margin, while SaasBO exhibits a closer prediction accuracy. In addition, we examined the alignment between the prediction and the ground-truth across the test examples from Mopta and DNA. As illustrated in Fig. 4, the prediction (as well the predictive standard deviation) of GP-ARD matches the ground-truth equally as well as SaasBO-NUTS, with a slight improvement over GP (using a single length-scale parameter).
Together these results demonstrate that directly incorporating high-dimensional inputs into the kernel function does not necessarily compromise the learning performance. In other words, this implies that, even when the input space is high dimensional, commonly used kernels (e.g., the Matérn-5/2) remains expressive enough to capture the similarities in the space, allowing them to perform well in covariance modeling. On the contrary, enforcing (unnecessary) low-rank structures or sparsity might cause information loss, potentially detrimental to the learning performance. Our observation is consistent with the recent work in kernel operator learning (Batlle et al., 2024), where the standard kernel (GP) regression has demonstrated the state-of-the-art performance in learning function-to-function mappings. In this context, the input comprises of a discretized function, typically of very high-dimensionality (e.g., several thousands).
4.2.3 Computational Efficiency
So far, in our BO loop, the GP model is trained via maximum likelihood estimation (MLE). Although the performance has been promising, it is worth investigating whether a full Bayesian approach would yield substantial improvements. In particular, we can use MCMC to perform posterior inference of the GP parameters, and then integrate them out during the prediction. We conducted another round of BO experiments. For each step, we used NUTS to train GP-ARD. We adopted the same settings as used in SaasBO-NUTS (Eriksson and Jankowiak, 2021), where the mean parameter is assigned the standard Gaussian prior $\mathcal{N}(0,1)$, amplitude $\text{Gamma}(2,0.15)$, and noise variance $\text{Gamma}(0.9,10)$. We then assigned a uniform prior over each length-scale, $\text{Uniform}(0.001,30)$. The number of warm up steps was set to 512, and after the warm up stage, we ran 256 steps, and selected the samples at every 16-th step as the posterior samples. We show the running and final optimization performance in Fig 3 , and Appendix Fig. 7. It can be observed that while NUTS contributes to a slight improvement in several tasks, the enhancement is very marginal. However, running NUTS is much more computationally costly compared to MLE. For example, on NAS201 and DNA, NUTS takes 68x and 40x running time than MLE in each optimization step. Therefore, for practical purposes, using maximum likelihood estimation (MLE) for Bayesian Optimization with GP-ARD appears sufficient. It is computationally efficient, and the sacrifice of optimization performance is minimal compared to much more costly training methods like NUTS.
By contrast, for more complex surrogate models, such as SaasBO, the MCMC based training might be more crucial for achieving good performance. As depicted in Fig. 2, SaasBO-NUTS demonstrates a substantial improvement over SaasBO-MAP across all real-world benchmarks, excluding NAS201. However, the computational cost of running SaasBO-NUTS is considerably higher, ranging from 5.4x (NAS201) to 33.9x (SVM, DNA) compared to the running time of SaasBO-MAP. This substantial increase in computational requirements may pose a severe bottleneck for practical usage. Detailed per-optimization-step running time for these methods are provided in Table 4 and Appendix Table 6.
Task GP-ARD GP SaasBO ALEBO-10 ALEBO-20
[…]
Table 3: Test log likelihood of each surrogate model. The results were averaged from five runs.
Mopta08 (124) Rover (60) NAS201 (30) DNA (180) SVM (388)
GP-ARD 11.23 9.84 5.76 14.34 32.16
GP-ARD-NUTS 492.62 404.17 393.62 575.03 768.29
SaasBO-MAP 46.53 41.36 42.35 51.33 67.37
SaasBO-NUTS 985.76 375.06 229.07 1739.56 2269.50
Table 4: Average running time (in seconds) per optimization step in real-world benchmark tasks.
Figure 3: Optimization performance when using GP-ARD trained with maximum likelihood estimation (MLE) and NUTS on SVM (388) and DNA (180). The results were summarized from 10 runs.
(a) GP-ARD
(b) GP
(c) SaasBO-NUTS
Figure 4: Alignment between the prediction (y-axis) and ground-truth function values (x-axis). The top row is the results for Mopta08 (124) and the second row DNA (180).
5 Conclusion
A longstanding belief suggests that standard Gaussian process is insufficient to model high-dimensional functions, and Bayesian optimization with standard GP is expected to exhibit poor performance when optimizing functions with more than 20 variables. However, through a comprehensive investigation across various benchmark tasks, this paper challenges and overturns the prevailing notion.
Our findings reveal that standard GP excels in function estimation and uncertainty calibration, even when dealing with hundreds of input variables. Bayesian Optimization (BO) conducted with standard GP demonstrates remarkable performance in high-dimensional optimization, frequently outperforming existing high-dimensional BO methods by a substantial margin. Those methods, incorporating strong structural assumptions and employing more complex surrogate modeling, exhibit less robustness to various target functions, whose intrinsic structural properties are often agnostic in practice. Furthermore, the standard GP, benefiting from a parsimonious parameter space, can achieve satisfactory performance using only maximum likelihood estimation, thus eliminating the need for more computationally expensive Monte-Carlo methods in practical applications. Given these revelations, we advocate for a comprehensive exploration of the potential of the standard GP in high-dimensional optimization tasks.
Appendix
A More Experiment Details
A.1 Definition of Synthetic Functions
Stybtang. We used the following slightly-modified Stybtang function,
$f({\bf x})=\frac{1}{2}\sum_{i=1}^{D}\left((x_{i}-c_{i})^{4}-16(x_{i}-c_{i})^{2}+5(x_{i}-c_{i})\right),$
where ${\bf x}\in[-5,5]^{D}$, and we set $c_{1},\dots,c_{D}]$ as an evenly spaced sequence in $[0,7.5]$ ($c_{1}=0$ and $c_{D}=7.5$). The optimum is at ${\bf x}^{\dagger}=[c_{1}-2.903534,\ldots,c_{D}-2.903534]$.
Rosenbrock. We used the following Rosenbrock function,
$f({\bf x})=\sum_{i=1}^{D-1}\left[100\left((x_{i+1}-c_{i+1})-(x_{i}-c_{i})^{2}\right)^{2}+\left(1-(x_{i}-c_{i})\right)^{2}\right],$
where ${\bf x}\in[-2.048,2.048]^{D}$. We set $c_{1},...,c_{D}]$ as an evenly spaced sequence in [2,2][-2,2], where $c_{1}=-2$ and $c_{D}=2$. The optimum is at $x^{\dagger}=[c_{1}+1,...,c_{D}+1]$.
Ackley. We used the Ackley function defined in (Surjanovic and Bingham, ),
$f({\bf x})=-20\text{ exp}\left(-0.2\sqrt{\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}}\right)-\text{exp}\left(\frac{1}{d}\sum_{i=1}^{d}\text{cos}(2{\pi}x_{i})\right)+20+\text{exp}(1),$
where each $x_{i}\in[-32.768,32.768]$ and the (global) optimum is at $x^{\dagger}=[0,\ldots,0]$.
Hartmann6. The function is given by
$f({\bf x})=-\sum_{i=1}^{4}{\alpha}_{i}{\exp}\left(-\sum_{j=1}^{6}A_{ij}(x_{j}-P_{ij})^{2}\right),$
where each $x_{i}\in[0,1]$, ${\bf A}=[A_{ij}]$ and ${\bf P}=[P_{ij}]$ are two pre-defined matrices as in (Surjanovic and Bingham, ). The global optimum is at ${\bf x}^{\dagger}=[0.20169,0.150011,0.476874,0.275332,0.6573]$.
Figure 5: Maximum Function Value Obtained vs. Number of Steps in synthetic benchmarks. The results were averaged from 10 runs.
Figure 6: Maximum Function Value Obtained vs. Number of Steps in real-world benchmarks. The results were averaged from 10 runs.
A.2 NN-Prune Details
Given specific per-layer pruning rates ${\bf x}=[x_{1},\ldots,x_{151}]$, the pruning procedure is as follows. The weights in each layer $i$ are first ranked according to their magnitudes, and the bottom $x_{i}$ portion of the weights are removed. The pruned network is then fine tuned on Tiny ImageNet (Le and Yang, 2015). The prediction accuracy after pruning and fine tuning on a separate test dataset is used as the target function. To respect the global pruning rate constraint $\alpha$, we add to the objective a soft penalty term $\lambda\cdot\max(\alpha-\beta,0)$, where $\beta$ is the global pruning rate reflected from $\bf x$}. Note that the penalty is only incurred when $\alpha>\beta$. When $\alpha<\beta$, namely, the network is pruned to be sparser than the expected level, the test accuracy typically experienced a significant decline. Consequently, the optimization process automatically adjusts to increase $\beta$. We set $\lambda=8.0$ to introduce a strong penalty. In our experiment, we employed $\alpha=0.5$.
A.3 Implementation
A.4 Computing Environment
We conducted the experimental investigations on a large computer cluster equipped with Intel Cascade Lake Platinum 8268 chips. For the NN-Prun problem, we ran the experiments on a Linux workstation, with a 16-core Intel 13900K CPU and one NVIDIA RTX 4090 GPU.
Task GP-ARD GP SaasBO ALEBO-10 ALEBO-20
[…]
Table 5: Mean Squared Error (MSE) of each surrogate model. The results were averaged over five runs.
Rosenbrock Ackley Stybtang Rosenbrock Ackley Hartmann6
(100,100) (150, 150) (200, 200) (300, 100) (300, 150) (300, 6)
GP-ARD 8.85 12.94 13.67 12.63 9.99 8.81
GP-ARD-NUTS 857.21 850.23 1120.28 1276.42 1049.78 948.25
SaasBO-MAP 70.17 70.69 66.35 100.79 85.86 85.72
SaasBO-NUTS 1294.90 1621.60 1511.99 1938.53 925.80 2204.54
Table 6: Average running time (in seconds) per optimization step on synthetic benchmark tasks.
Figure 7: Optimization performance when using GP-ARD trained with maximum likelihood estimation (MLE) and NUTS on the rest of benchmarks. The results were summarized from 10 runs.

### v2

@@
 We thus advocate for a re-evaluation and in-depth study of the potential of standard BO in addressing high-dimensional problems.
+The code for our experiment is available at https://github.com/XZT008/Standard-GP-is-all-you-need-for-HDBO.
 1 Introduction
@@
 Suppose we have collected a set of training inputs ${\bf X}=[{\bf x}_{1},\ldots,{\bf x}_{N}]^{\top}$ and (noisy) outputs ${\bf y}=[y_{1},\ldots,y_{N}]^{\top}$.
-Due to the GP prior, the function values at $\bf X$}, denoted as ${\bf f}=[f({\bf x}_{1}),\ldots,f({\bf x}_{N})]^{\top}$, follow a multi-variate Gaussian distribution, $p({\bf f})=\mathcal{N}({\bf f}|{\bf m},{\bf K})$, where ${\bf m}=[m({\bf x}_{1}),\ldots,m({\bf x}_{N})]^{\top}$, $\bf K$} is the kernel matrix on $\bf X$}, and each ${\bf K}]_{ij}=\kappa({\bf x}_{i},{\bf x}_{j})$.
+Due to the GP prior, the function values at ${\bf X}$, denoted as ${\bf f}=[f({\bf x}_{1}),\ldots,f({\bf x}_{N})]^{\top}$, follow a multi-variate Gaussian distribution, $p({\bf f})=\mathcal{N}({\bf f}|{\bf m},{\bf K})$, where ${\bf m}=[m({\bf x}_{1}),\ldots,m({\bf x}_{N})]^{\top}$, ${\bf K}$ is the kernel matrix on ${\bf X}$, and each $[{\bf K}]_{ij}=\kappa({\bf x}_{i},{\bf x}_{j})$.
 Next, a Gauss noise model is used to model the likelihood of ${\bf y}$, namely, $p({\bf y}|{\bf f})=\mathcal{N}({\bf y}|{\bf f},\sigma^{2}{\bf I})$ where $\sigma^{2}$ is the noise variance.
@@
   • Rover (60), a rover trajectory planning problem from (Wang et al., 2018). The goal is to find an optimal trajectory determined by the location of 30 waypoints in a 2D environment.
-  • DNA (180) (Šehić et al., 2022), a hyper-parameter optimization problem that optimizes 180 regularization parameters for weighted lasso regression on an DNA dataset (Mills, 2020).
-Prior analysis (Šehić et al., 2022) shows that only around 43 variables are relevant to the target function.
+  • DNA (180) (vSehić et al., 2022), a hyper-parameter optimization problem that optimizes 180 regularization parameters for weighted lasso regression on an DNA dataset (Mills, 2020).
+Prior analysis (vSehić et al., 2022) shows that only around 43 variables are relevant to the target function.
   • NAS201 (30)  (Dong and Yang, 2020), a neural architecture search problem on CIFAR-100 dataset.
@@
 Task GP-ARD GP SaasBO ALEBO-10 ALEBO-20
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+Mopta08(124) 0.613$\pm$0.156 -0.198$\pm$0.233 0.584$\pm$0.158 -1.176$\pm$0.080 -1.177$\pm$0.327
+Rover(60) -1.209$\pm$0.194 -0.908$\pm$0.068 -0.909$\pm$0.108 -1.270$\pm$0.094 -1.687$\pm$0.251
+DNA(180) 0.005$\pm$0.175 -0.459$\pm$0.142 -0.066$\pm$0.225 -1.346$\pm$0.108 -1.572$\pm$0.432
+SVM(388) 0.296$\pm$0.097 -0.090$\pm$0.065 0.160$\pm$0.113 -1.175$\pm$0.116 -1.077$\pm$0.148
+Ackley(150, 150) 0.720$\pm$0.092 0.899$\pm$0.058 0.809$\pm$0.053 -0.852$\pm$0.091 -0.266$\pm$0.091
+Ackley150(300, 150) 0.617$\pm$0.191 0.367$\pm$0.081 0.536$\pm$0.095 -0.870$\pm$0.097 -0.478$\pm$0.204
+Hartmann6 (300, 6) 1.681$\pm$0.110 -1.031$\pm$0.117 2.152$\pm$0.971 -1.433$\pm$0.249 -1.689$\pm$0.411
+Rosenbrock(100, 100) 0.602$\pm$0.069 0.172$\pm$0.099 0.553$\pm$0.084 -1.007$\pm$0.121 -0.854$\pm$0.355
+Rosenbrock(300, 100) 0.436$\pm$0.098 -0.256$\pm$0.071 0.265$\pm$0.115 -1.118$\pm$0.065 -1.380$\pm$0.329
+Stybtang(200, 200) 0.558$\pm$0.122 0.215$\pm$0.076 0.463$\pm$0.098 -1.050$\pm$0.099 -1.238$\pm$0.436
 Table 3: Test log likelihood of each surrogate model. The results were averaged from five runs.
@@
 $f({\bf x})=\frac{1}{2}\sum_{i=1}^{D}\left((x_{i}-c_{i})^{4}-16(x_{i}-c_{i})^{2}+5(x_{i}-c_{i})\right),$
-where ${\bf x}\in[-5,5]^{D}$, and we set $c_{1},\dots,c_{D}]$ as an evenly spaced sequence in $[0,7.5]$ ($c_{1}=0$ and $c_{D}=7.5$).
+where ${\bf x}\in[-5,5]^{D}$, and we set $[c_{1},\dots,c_{D}]$ as an evenly spaced sequence in $[0,7.5]$ ($c_{1}=0$ and $c_{D}=7.5$).
 The optimum is at ${\bf x}^{\dagger}=[c_{1}-2.903534,\ldots,c_{D}-2.903534]$.
@@
 $f({\bf x})=\sum_{i=1}^{D-1}\left[100\left((x_{i+1}-c_{i+1})-(x_{i}-c_{i})^{2}\right)^{2}+\left(1-(x_{i}-c_{i})\right)^{2}\right],$
-where ${\bf x}\in[-2.048,2.048]^{D}$. We set $c_{1},...,c_{D}]$ as an evenly spaced sequence in [2,2][-2,2], where $c_{1}=-2$ and $c_{D}=2$. The optimum is at $x^{\dagger}=[c_{1}+1,...,c_{D}+1]$.
+where ${\bf x}\in[-2.048,2.048]^{D}$. We set $[c_{1},...,c_{D}]$ as an evenly spaced sequence in $[-2,2]$, where $c_{1}=-2$ and $c_{D}=2$. The optimum is at $x^{\dagger}=[c_{1}+1,...,c_{D}+1]$.
 Ackley. We used the Ackley function defined in (Surjanovic and Bingham, ),
@@
 The prediction accuracy after pruning and fine tuning on a separate test dataset is used as the target function.
-To respect the global pruning rate constraint $\alpha$, we add to the objective a soft penalty term $\lambda\cdot\max(\alpha-\beta,0)$, where $\beta$ is the global pruning rate reflected from $\bf x$}.
+To respect the global pruning rate constraint $\alpha$, we add to the objective a soft penalty term $-\lambda\cdot\max(\alpha-\beta,0)$, where $\beta$ is the global pruning rate reflected from ${\bf x}$.
 Note that the penalty is only incurred when $\alpha>\beta$.
@@
 Task GP-ARD GP SaasBO ALEBO-10 ALEBO-20
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+Mopta08(124) 0.048$\pm$0.015 0.126$\pm$0.041 0.052$\pm$0.018 0.654$\pm$0.146 0.659$\pm$0.096
+Rover(60) 0.493$\pm$0.126 0.449$\pm$0.108 0.453$\pm$0.121 0.813$\pm$0.168 1.266$\pm$0.352
+DNA(180) 0.098$\pm$0.028 0.175$\pm$0.070 0.115$\pm$0.039 0.827$\pm$0.173 0.955$\pm$0.335
+SVM(388) 0.042$\pm$0.010 0.098$\pm$0.012 0.061$\pm$0.017 0.624$\pm$0.159 0.493$\pm$0.132
+Ackley(150, 150) 0.010$\pm$0.003 0.004$\pm$0.001 0.004$\pm$0.001 0.367$\pm$0.074 0.102$\pm$0.015
+Ackley(300, 150) 0.010$\pm$0.003 0.018$\pm$0.004 0.010$\pm$0.002 0.360$\pm$0.109 0.177$\pm$0.075
+Hartmann6(300, 6) 0.020$\pm$0.009 0.769$\pm$0.124 0.077$\pm$0.045 1.012$\pm$0.154 1.286$\pm$0.210
+Rosenbrock(100, 100) 0.049$\pm$0.006 0.124$\pm$0.020 0.062$\pm$0.008 0.601$\pm$0.095 0.807$\pm$0.248
+Rosenbrock(300, 100) 0.048$\pm$0.012 0.202$\pm$0.035 0.073$\pm$0.021 0.568$\pm$0.057 1.290$\pm$0.882
+Stybtang(200, 200) 0.040$\pm$0.014 0.078$\pm$0.012 0.044$\pm$0.014 0.502$\pm$0.066 0.861$\pm$0.329
 Table 5: Mean Squared Error (MSE) of each surrogate model. The results were averaged over five runs.

### v3

@@
-Standard Gaussian Process is All You Need for High-Dimensional Bayesian Optimization
+Standard Gaussian Process Can Be Excellent for High-Dimensional Bayesian Optimization
 Abstract
 There has been a long-standing and widespread belief that Bayesian Optimization (BO) with standard Gaussian process (GP), referred to as standard BO, is ineffective in high-dimensional optimization problems.
-This perception may partly stem from the intuition that GPs struggle with high-dimensional inputs for covariance modeling and function estimation.
-While these concerns seem reasonable, empirical evidence supporting this belief is lacking.
+While this belief sounds reasonable, strong empirical evidence is lacking.
 In this paper, we systematically investigated BO with standard GP regression across a variety of synthetic and real-world benchmark problems for high-dimensional optimization.
-Surprisingly, the performance with standard GP consistently ranks among the best, often outperforming existing BO methods specifically designed for high-dimensional optimization by a large margin.
-Contrary to the stereotype, we found that standard GP can serve as a capable surrogate for learning high-dimensional target functions.
-Without strong structural assumptions, BO with standard GP not only excels in high-dimensional optimization but also proves robust in accommodating various structures within the target functions.
-Furthermore, with standard GP, achieving promising optimization performance is possible by only using maximum likelihood estimation, eliminating the need for expensive Markov-Chain Monte Carlo (MCMC) sampling that might be required by more complex surrogate models.
+We found that, surprisingly, when using Matérn kernels and Upper Confidence Bound (UCB), standard BO consistently achieves top-tier performance, often outperforming other BO methods specifically designed for high-dimensional optimization.
+Contrary to the stereotype, we found that standard GP equipped with Matérn kernels can serve as a capable surrogate for learning high-dimensional functions.
+Without strong structural assumptions, BO with standard GP not only excels in high-dimensional optimization but also is robust in accommodating various structures within target functions.
+Furthermore, with standard GP, achieving promising optimization performance is possible via maximum a posterior (MAP) estimation with diffuse priors or merely maximum likelihood estimation, eliminating the need for expensive Markov-Chain Monte Carlo (MCMC) sampling that might be required by more complex surrogate models.
+In parallel, we also investigated and analyzed alternative popular settings in running standard BO, which, however, often fail in high-dimensional optimization.
+This might link to the a few failure cases reported in literature.
 We thus advocate for a re-evaluation and in-depth study of the potential of standard BO in addressing high-dimensional problems.
@@
 Many applications require optimizing complex functions, allowing queries only for the function values, possibly with noises, yet without any gradient information.
-Bayesian Optimization (BO) (Snoek et al., 2012; Mockus, 2012) is a popular approach to address such optimization challenges.
-BO typically employs Gaussian process (GP) (Williams and Rasmussen, 2006) as a probabilistic surrogate.
-It iteratively approximates the target function, integrates the posterior information to optimize an acquisition function for generating new inputs for querying, then updates the GP model with new examples, and concurrently approaches the optimum.
-Despite numerous successes, there has been a widespread belief that BO with standard GP regression, referred to as standard BO, is limited to low-dimensional optimization problems (Frazier, 2018; Nayebi et al., 2019; Eriksson and Jankowiak, 2021; Moriconi et al., 2020; Letham et al., 2020; Wang et al., 2016; Li et al., 2016).
-It is commonly thought that the number of optimization variables should not exceed 15 or 20, as beyond this threshold, BO is prone to failure (Frazier, 2018; Nayebi et al., 2019).
+Bayesian Optimization (BO) [28, 23] is a popular approach for addressing such challenges.
+BO typically employs Gaussian process (GP) [33] as a probabilistic surrogate.
+It iteratively approximates the target function, integrates the posterior information to maximize an acquisition function for generating new inputs at which to query, then updates the GP model with new examples, and concurrently approach the optimum.
+Despite numerous successes, there has been a widespread belief that BO with standard GP regression, referred to as standard BO, is limited to low-dimensional optimization problems [8, 25, 6, 24, 18, 32, 19].
+It is commonly thought that the number of optimization variables should not exceed 15 or 20, as beyond this threshold, BO is prone to failure [8, 25].
 This belief may partly stem from the intuition that, as a kernel method, GPs struggle to handle high-dimensional inputs for covariance modeling, rending them incapable of providing an expressive prior for high-dimensional functions.
-A continuous line of research is dedicated to developing novel high-dimensional Bayesian Optimization (BO) methods.
+Kernel GP Training Acquisition Function
+Matérn ✓ L-BFGS/MAP/Diffuse prior ✓ UCB ✓
+ADAM/MAP/Diffuse prior ✓
+L-BFGS/MAP/Concentrated prior ✗ EI ✗
+SE ✗ ADAM/MAP/Concentrated prior ✗
+L-BFGS/MLE ✗ TS ✗
+ADAM/MLE ✓
+Table 1: Favorable and unfavorable settings of running standard BO for high-dimensional optimization.
+SE: square exponential; EI: expected improvement; TS: Thompson sampling; MLE: maximum likelihood estimation.
+A continuous line of research is dedicated to developing novel high-dimensional BO methods.
 The fundamental strategy is to impose strong structural assumptions into surrogate modeling so as to avoid directly dealing with high-dimensional inputs.
-One class of methods assumes a decomposition structure within the functional space (Kandasamy et al., 2016; Rolland et al., 2018; Han et al., 2020; Ziomek and Bou-Ammar, 2023).
+One class of methods assumes a decomposition structure within the functional space [16, 26, 10, 34].
 In other words, the target function can be expressed as the summation of a group of low-dimensional functions, each operating over a small number of variables.
 Another family of methods assumes that the inputs are intrinsically low-rank.
-These methods either project the original inputs into a low-dimensional space (Wang et al., 2016; Nayebi et al., 2019; Letham et al., 2020) or use sparse-inducing priors to trim down massive input variables (Eriksson and Jankowiak, 2021).
+These methods either project the original inputs into a low-dimensional space [32, 25, 18] or use sparse-inducing priors to trim down massive input variables [6].
 Subsequently, GP surrogates are built with the reduced input dimensions.
@@
 To bridge this gap, we systematically investigated BO with standard GP regression across a variety of synthetic and real-world benchmark problems.
-Surprisingly, the performance of standard BO consistently ranks among the best and often surpasses existing methods by a substantial margin.
-These methods are specifically designed for high-dimensional optimization This leads us to call for a re-evaluation and in-depth study of the potential of standard BO for high-dimensional optimization.
-The major discoveries are as follows.
+Surprisingly, we found that, with ARD Matérn kernels and UCB as the acquisition function, the performance of standard BO consistently ranks among the best and often surpasses existing methods by a substantial margin.
+Those methods are specifically designed for high-dimensional optimization.
+This leads us to call for a re-evaluation and in-depth study of the potential of standard BO for high-dimensional optimization.
+The major discoveries are summarized as follows.
   • Optimization performance.
-We conducted an investigation using BO with standard GP regression across six synthetic and six real-world high-dimensional optimization tasks.
-Our selection includes eleven benchmarks widely used in literature, and a new real-world benchmark task on neural network pruning.
+We investigated BO with standard GP regression across eleven benchmarks widely used in literature, including six synthetic and five real-world high-dimensional optimization tasks.
 The number of optimization variables varies from 30 to 388.
-In comparison with six state-of-the-art high-dimensional BO methods, standard BO consistently achieves the best (or close to the best) optimization performance across all the tasks.
+In comparison with six state-of-the-art high-dimensional BO methods, standard BO with ARD Matérn kernel and UCB consistently achieves the best (or close to the best) optimization performance across all the tasks.
 In contrast, the competing methods demonstrate strong performance when the target function aligns with their structural assumption.
-However, in cases of mismatch or when the structure of the target function is agnostic (e.g., in real-world problems), the performance of the competing methods often experiences a significant drop.
+However, in cases of mismatch or when the structure of the target function is agnostic, such as those in real-world problems, the performance of the competing methods often experiences a significant drop.
 This underscores that BO with standard GP not only can excel in high-dimensional optimization but also is robust in accommodating various structures within the target functions.
   • Surrogate Learning.
-We investigated the task of learning high-dimensional functions using standard GP regression.
-Across twelve benchmark datasets, our findings reveal that the prediction of standard GP aligns well with the ground-truth, mostly performing better than more complex GP models that induce low-rank structures.
-This observation is in line with the recent results in kernel operator learning (Batlle et al., 2024), where inputs are discretized functions — typically high-dimensional.
+We investigated the task of learning high-dimensional functions with standard GP regression.
+Across ten benchmark datasets, our findings reveal that the prediction of standard GP equipped with Matérn kernels aligns well with the ground-truth, mostly performing better than more complex GP models that induce low-rank structures.
+This observation is in line with the recent results in kernel operator learning [3], where inputs are discretized functions — typically high-dimensional (hundreds or thousands of dimensions).
 It confirms that standard GP can serve as a capable surrogate for high-dimensional functions.
   • Computational Efficiency.
-We discovered that, with standard GP, achieving promising optimization performance is possible by using maximum likelihood estimation.
-The marginalization of the parameters (e.g., kernel parameters and noise variance) via Markov-Chain Monte Carlo (MCMC) provides only minor benefits while significantly increasing the training cost.
+We discovered that, with standard GP, achieving promising optimization performance is possible by using MAP estimation with diffuse priors or just maximum likelihood estimation.
+The marginalization of the parameters (kernel parameters and noise variance) via Markov-Chain Monte Carlo (MCMC) provides minor benefits while significantly increasing the training cost.
 In contrast, due to the usage of more complex surrogate models, training for alternative high-dimensional BO methods is often more expensive.
 Some of these methods heavily rely on extensive MCMC sampling to prevent overfitting, making the training of the surrogate model a computational bottleneck.
+  • Failure Modes with Other Settings.
+We found failure modes of standard BO with other popular settings, such as using vanilla Square-Exponential (SE) kernels, Expected-Improvement (EI) and Thompson sampling (TS) as acquisition functions, and MAP estimation with concentrated Gamma priors over length-scale parameters.
+We conducted a careful ablation study to reveal the failure of these choices in both optimization and function learning, providing analysis and discussions.
+In Table 1, we list the discovered favorable and unfavorable settings for high-dimensional optimization.
 2 Standard Bayesian Optimization
@@
 We can only query the function value at any input ${\bf x}\in\mathcal{X}$, possibly with noises, and no gradient information is accessible.
-We aim to find
-${\bf x}^{\dagger}=\operatornamewithlimits{argmax}_{{\bf x}\in\mathcal{X}}f({\bf x}).$ (1)
+We aim to find ${\bf x}^{\dagger}=\operatornamewithlimits{argmax}_{{\bf x}\in\mathcal{X}}f({\bf x})$.
 To achieve this, Bayesian Optimization (BO) employs a probabilistic surrogate model to learn and predict $f(\cdot)$ across the domain $\mathcal{X}$, while also quantifying the uncertainty of the prediction.
@@
 Concurrently, the acquired new example is incorporated into the training dataset, and the surrogate model is retrained to improve the accuracy.
-BO begins by querying at a few (randomly selected) input locations, and trains an initial surrogate model.
+BO begins by querying at a few (typically randomly selected) input locations, and trains an initial surrogate model.
 Then the iterative procedure repeats until BO converges to the optimum or a predefined stopping criterion is met.
-The standard BO adopts GP regression (Williams and Rasmussen, 2006) for surrogate modeling. Specifically, a GP prior is placed over the target function,
-$f\sim\mathcal{GP}\left(m({\bf x}),\kappa({\bf x},{\bf x}^{\prime})\right)$ (2)
-where $m({\bf x})$ is the mean function, which is often set as a constant, and $\kappa(\cdot,\cdot)$ is the covariance function, which is often chosen as a Mercer kernel function.
-For example, a popular kernel used in BO is the Matérn-5/2 Automatic Relevance Determination (ARD) kernel,
-$\kappa({\bf x},{\bf x}^{\prime})=a\left(1+{\sqrt{5}\tau}+\frac{5\tau^{2}}{3}\right)\exp\left(-{\sqrt{5}\tau}\right)$ (3)
-where
-$\tau=\sqrt{({\bf x}-{\bf x}^{\prime})^{\top}{\rm diag}(\frac{1}{\bm{\eta}})({\bf x}-{\bf x}^{\prime})},$ (4)
-$a>0$ is the amplitude, and $\bm{\eta}>{\bf 0}$ are the length-scale parameters.
+The standard BO adopts GP regression [33] for surrogate modeling.
+Specifically, a GP prior is placed over the target function, $f\sim\mathcal{GP}\left(m({\bf x}),\kappa({\bf x},{\bf x}^{\prime})\right)$, where $m({\bf x})$ is the mean function, which is often set as a constant, and $\kappa(\cdot,\cdot)$ is the covariance function, which is often chosen as a Mercer kernel function.
+For example, a popular kernel used in BO is the Matérn-5/2 kernel with Automatic Relevance Determination (ARD),
+[…]
+where $\tau=\sqrt{({\bf x}-{\bf x}^{\prime})^{\top}{\rm diag}(\frac{1}{\bm{\eta}})({\bf x}-{\bf x}^{\prime})}$, $a>0$ is the amplitude, and $\bm{\eta}>{\bf 0}$ are the length-scale parameters.
 We refer to this kernel as ARD because each input dimension has a distinct length-scale parameter.
@@
 Next, a Gauss noise model is used to model the likelihood of ${\bf y}$, namely, $p({\bf y}|{\bf f})=\mathcal{N}({\bf y}|{\bf f},\sigma^{2}{\bf I})$ where $\sigma^{2}$ is the noise variance.
-One can marginalize ${\bf f}$ out,
+One can marginalize ${\bf f}$ out to obtain
 $p({\bf y}|{\bf X})=\mathcal{N}({\bf y}|{\bf m},{\bf K}+\sigma^{2}{\bf I}).$ (2)
-To estimate the GP parameters (e.g., $a$, $\bm{\eta}$, and $\sigma^{2}$), one can use MCMC or simply maximize (5) to obtain a point estimate.
-The predictive distribution (given the GP parameters) has a closed form.
-At any new input ${\bf x}^{*}$, since $f({\bf x}^{*})$ and ${\bf y}$ jointly follow a multi-variate Gaussian distribution, the predictive distribution is a conditional Gaussian distribution,
+To estimate the GP parameters (e.g., $a$, $\bm{\eta}$, and $\sigma^{2}$), one can place a prior over these parameters and use MCMC to compute posterior samples or perform MAP estimation.
+One can also simply maximize (2) to perform MLE.
+The predictive distribution (given the GP parameters) has a closed form: at any new input ${\bf x}^{*}$, the predictive distribution is conditional Gaussian,
 $p\left(f({\bf x}^{*})|{\bf y}\right)=\mathcal{N}\left(f({\bf x}^{*})|\mu({\bf x}^{*}),v({\bf x}^{*})\right),$ (3)
-where
-$\mu({\bf x}^{*})$ $=m({\bf x}^{*})+\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}({\bf y}-{\bf m})$ (7)
-$v({\bf x}^{*})$ $=\kappa({\bf x}^{*},{\bf x}^{*})-\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}\kappa({\bf X},{\bf x}^{*}).$ (8)
+where $\mu({\bf x}^{*})=m({\bf x}^{*})+\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}({\bf y}-{\bf m})$, and $v({\bf x}^{*})=\kappa({\bf x}^{*},{\bf x}^{*})-\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}\kappa({\bf X},{\bf x}^{*})$.
 In each iteration, given the current GP surrogate, we maximize an acquisition function to identify the next input location at which to query.
@@
 $\text{UCB}({\bf x})=\mu({\bf x})+\lambda\sqrt{v({\bf x})},\;\;{\bf x}\in\mathcal{X},$ (4)
-where $\lambda$ represents the exploration level.
-There have been other acquisition functions, such as Expected Improvement (EI) (Jones et al., 1998), Thompson sampling (Russo et al., 2018), among others.
+where $\lambda$ represents the exploration level. There have been other popular acquisition functions, such as Expected Improvement (EI) [14], Thompson sampling (TS) [27], among others.
 3 High Dimensional Bayesian Optimization
 An enduring and widespread belief is that when the input dimension $d$ is high (e.g., a few hundreds), the standard BO is prone to failure.
-This belief might partly arise from an intuition that commonly used (stationary) kernels, such as in (3), could encounter challenges in dealing with high-dimensional inputs, making GP struggle in capturing the target function.
+This belief might partly arise from an intuition that commonly used kernels, such as in (1), could encounter challenges in dealing with high-dimensional inputs, making GP struggle in capturing the target function.
 As a result, there is a dedicated line of research aimed at developing novel high-dimensional BO methods.
@@
 The first class of methods assumes a decomposition structure within the functional space. Specifically, the target function $f$ is modeled as
-$f({\bf x})=\sum_{j=1}^{M}f_{j}({\bf x}^{j}),\;\;f_{j}\sim\mathcal{GP}\left(m_{j}({\bf x}^{j}),\kappa_{j}(\cdot,\cdot)\right),$ (10)
+$f({\bf x})=\sum\nolimits_{j=1}^{M}f_{j}({\bf x}^{j}),\;\;f_{j}\sim\mathcal{GP}\left(m_{j}({\bf x}^{j}),\kappa_{j}(\cdot,\cdot)\right),$ (5)
 where each ${\bf x}^{j}\subset{\bf x}$ is a small group of input variables, and ${\bf x}={\bf x}^{1}\cup\ldots\cup{\bf x}^{M}$.
-There can be a variety of choices for the number of group $M$ and each variable group ${\bf x}^{j}$.
-In (Kandasamy et al., 2016), the input ${\bf x}$ is partitioned into non-overlapped groups.
+There can be a variety of choices for the group number $M$ and each variable group ${\bf x}^{j}$.
+In [16], the input ${\bf x}$ is partitioned into non-overlapped groups.
 After every a few BO iterations, a new partition is selected from a set of random partitions.
 The selection is based on the model evidence.
-In (Rolland et al., 2018), the variable groups can overlap, and are represented as the maximum cliques on a dependency graph.
-A Gibbs sampling method was developed to learn the structure of the dependency graph from data.
-Han et al.
-(2020) proposed Tree-UCB, which restricted the dependency graph to be tree-structured so as to boost the efficiency of structure learning.
-However, the most recent work (Ziomek and Bou-Ammar, 2023) points out that, learning the group structure through relatively small data can be misleading.
+In [26], the variable groups can overlap, and are represented as the maximum cliques on a dependency graph.
+A Gibbs sampling method was developed to learn the structure of the dependency graph.
+Han et al., [10] proposed Tree-UCB, which restricted the dependency graph to be tree-structured so as to boost the efficiency of structure learning.
+However, the most recent work [34] points out that, learning the group structure through relatively small data can be misleading.
 A wrongly learned structure can cause queries to be stuck at local optima.
@@
 Low-dimensional Embedding.
-Wang et al.
-(2016) proposed a method named REMBO, building the GP surrogate in a low-dimensional embedding space, $\mathcal{Z}\subset\mathbb{R}^{d_{\text{emb}}}$ where $d_{\text{emb}}\ll d$.
+Wang et al., [32] proposed a method named REMBO, building the GP surrogate in a low-dimensional embedding space, $\mathcal{Z}\subset\mathbb{R}^{d_{\text{emb}}}$ where $d_{\text{emb}}\ll d$.
 The acquisition function is optimized in the embedding space.
@@
 If the recovered ${\bf x}$ is out of the domain $\mathcal{X}$, it will be clipped as a boundary point.
-To avoid clipping, Nayebi et al.
-(2019) proposed HESBO, which randomly duplicates ${\bf z}$ (and/or flip the signs) to form ${\bf x}$.
+To avoid clipping, Nayebi et al., [25] proposed HESBO, which randomly duplicates ${\bf z}$ (and/or flip the signs) to form ${\bf x}$.
 This way, the optimization of ${\bf z}$ can respect the same bounding constraints as in $\mathcal{X}$.
-This essentially impose further constraints on the embedding space.
-A more flexible solution, ALEBO, was proposed in  (Letham et al., 2020).
+This essentially imposes further constraints on the embedding space.
+A more flexible solution, ALEBO, was proposed in  [18].
 ALEBO also uses a random matrix ${\bf A}$ to recover the input ${\bf x}$.
@@
 Another strategy is to trim down the input variables.
-The recent start-of-the-art, SaasBO (Eriksson and Jankowiak, 2021), uses a horse-shoe prior (Carvalho et al., 2009) over the length-scale parameter for each variable.
+The recent start-of-the-art, SaasBO [6], uses a horse-shoe prior [4] over the length-scale parameter for each variable.
 As the horse-shoe prior induces strong sparsity, a massive number of variables can be pruned, substantially reducing the input dimension.
-For effective training with the horse-shoe prior, SaasBO used a No-U-Turn Sampler (NUTS) (Hoffman et al., 2014).
-While one can also use MAP estimation, the performance is often much inferior.
+For effective training with the horse-shoe prior, SaasBO used the No-U-Turn Sampler (NUTS) [11].
+While one can also use MAP estimation, the performance is often much worse.
 4 Comprehensive Evaluation
-We now examine and analyze the performance of standard BO in high-dimensional optimization. To this end, we conducted a series of systematic investigations.
+We now examine and analyze the performance of standard BO in high-dimensional optimization. To this end, we conducted a series of systematic experiments.
 4.1 Experimental Settings
-We employed twelve benchmarks, of which six are synthetic and the other six are real-world problems. For clarity in the discussion, all the tasks aim for function maximization.
-Synthetic Benchmarks.
-We considered four popular synthetic functions: Ackley, Rosenbrock, Hartmann6, and Stybtang.
-Definitions of these functions are provided in Section A.1 of the Appendix.
-Each task is represented by $\text{Fun}(d,d^{\prime})$, where $d$ is the input dimension that the BO methods optimize for, and $d^{\prime}$ is the number of effective variables ($d^{\prime}\leq d$).
-The target function is computed using the first $d^{\prime}$ variables.
-The benchmarks and the structures within the target functions are summarized in Table 1.
+We employed eleven benchmarks, of which six are synthetic and five are real-world problems. For clarity in the discussion, all the tasks aim for function maximization.
 $\text{Fun}(d,d^{\prime})$ Structure
@@
 Table 2: Synthetic benchmarks and the structures within the target function: $d$ is the input dimension, and $d^{\prime}$ is the number of effective variables to compute the target function.
+Synthetic Benchmarks.
+We considered four popular synthetic functions: Ackley, Rosenbrock, Hartmann6, and Stybtang.
+Definitions of these functions are provided in Section A.1 of the Appendix.
+Each task is represented by “$\text{Fun}(d,d^{\prime})$”, where $d$ is the input dimension that BO methods optimize for, and $d^{\prime}$ is the number of effective variables ($d^{\prime}\leq d$).
+The ground-truth of the target function values is computed with the first $d^{\prime}$ variables.
+The benchmarks and the structures within the target functions are summarized in Table 2.
 Real-World Benchmarks. We employed the following real-world benchmark problems.
-  • Mopta08 (124)  (Jones, 2008), a vehicle design problem.
+  • Mopta08 (124)  [13], a vehicle design problem.
 The objective is to minimize the mass of a vehicle with respect to 124 design variables, subject to 68 performance constraints.
-We followed (Eriksson and Jankowiak, 2021) to encode the constraints as a soft penalty, which is added into the objective.
-  • SVM (388) (Eriksson and Jankowiak, 2021), a hyper-parameter tuning problem that optimizes 3 regularization parameters and 385 length-scale parameters for a kernel support vector machine regression model.
-  • Rover (60), a rover trajectory planning problem from (Wang et al., 2018). The goal is to find an optimal trajectory determined by the location of 30 waypoints in a 2D environment.
-  • DNA (180) (vSehić et al., 2022), a hyper-parameter optimization problem that optimizes 180 regularization parameters for weighted lasso regression on an DNA dataset (Mills, 2020).
-Prior analysis (vSehić et al., 2022) shows that only around 43 variables are relevant to the target function.
-  • NAS201 (30)  (Dong and Yang, 2020), a neural architecture search problem on CIFAR-100 dataset.
-  • NN-Prune (151).
-We provided an additional real-world benchmark on neural network pruning (Blalock et al., 2020).
-We considered ResNet-152 (He et al., 2015), a deep residual network with 151 convolution layers and one linear layer, including more than 60 million parameters in total.
-We introduced a pruning rate $x_{i}$ for each convolution layer $i$.
-Given a global pruning ratio set to 0.5, the objective is to optimize the per-layer pruning rates to maximize the prediction accuracy on a test dataset after pruning and fine tuning.
-The details are given in Section A.2 of the Appendix.
+We followed [6] to encode the constraints as a soft penalty, which is added into the objective.
+  • SVM (388) [6], a hyper-parameter tuning problem that optimizes 3 regularization parameters and 385 length-scale parameters for a kernel support vector machine regression model.
+  • Rover (60), a rover trajectory planning problem from [31]. The goal is to find an optimal trajectory determined by the location of 30 waypoints in a 2D environment.
+  • DNA (180) [35], a hyper-parameter optimization problem that optimizes 180 regularization parameters for weighted lasso regression on an DNA dataset [22].
+Prior analysis [35] shows that only around 43 variables are relevant to the target function.
+  • NAS201 (30)  [5], a neural architecture search problem on CIFAR-100 dataset.
 Method Structural Assumption
@@
 RDUCB, Tree UCB Additive function decomposition
-GP-ARD, GP, TURBO None
-Table 2: BO methods and their structural assumption.
+SBO, TURBO, BNN None
+Table 3: BO methods and their structural assumption about the target function.
 Methods.
-To execute the standard BO, we used GP regression with Matérn-5/2 ARD kernel (see (3)).
-We also tested with Matérn-5/2 kernel with a single length-scale parameter.
-We denoted these two choices as GP-ARD and GP, respectively.
-Throughout the experiments, we used UCB as the acquisition function where the exploration level $\lambda$ was set to 1.5.
-We used GPyTorch (Gardner et al., 2021) for surrogate training and BOTorch (Balandat et al., 2020) for acquisition function computation.
-We compared with six state-of-the-art high-dimensional BO methods, including Tree UCB (Han et al., 2020), RDUCB (Ziomek and Bou-Ammar, 2023), HESBO (Nayebi et al., 2019), ALEBO (Letham et al., 2020), SaasBO (Eriksson and Jankowiak, 2021), and TURBO (Eriksson et al., 2020).
+To execute the standard BO (SBO), we used GP regression with ARD Matérn-5/2 kernel (see (1)).
+We used UCB as the acquisition function where the exploration level $\lambda$ was set to 1.5.
+We used GPyTorch [9] for surrogate training.
+To compare with other high-dimensional BO methods, we considered two training schemes.
+One performs MAP estimation with LBFGS.
+The maximum number of iterations was 15K and tolerance level 2E-9.
+We employed a diffuse prior over each length-scale, i.e., Uniform(0.001, 30).
+The noise variance was assigned a diffuse Gamma prior, Gamma(1.1, 0.05), and the amplitude the prior Gamma(2.0, 0.15).
+We denote this method as SBO-1.
+The second training scheme simply uses ADAM optimization [17] to perform MLE, with learning rate $0.1$ and the maximum number of epochs was set to $500$.
+We denote this method as SBO-2.
+We also tested with Matérn-5/2 kernel with a single length-scale parameter, which we denoted as SBO-3.
+We used BOTorch [2] for acquisition function maximization, which uses LBFGS.
+Additional evaluations on alternative choices of the kernel, training scheme, and acquisition function are provided in Appendix Section B.
+We compared with six state-of-the-art high-dimensional BO methods, including Tree UCB [10], RDUCB [34], HESBO [25], ALEBO [18], SaasBO [6], and TURBO [7].
 The first five have been introduced in Section 3.
-TURBO is designed for a special scenario where the target function can be extensively evaluated (e.g., tens of thousands evaluations).
-This is not common in BO applications, as the function evaluation is typically deemed expensive, and one aims to evaluate the target function as few times as possible.
+TURBO is designed for a special scenario where the target function can be extensively evaluated, e.g., tens of thousands evaluations.
+This is uncommon in BO applications, as the function evaluation is typically deemed expensive, and one aims to evaluate the target function as few as possible.
 TURBO searches for trust-regions in the input domain, learns a set of local GP models over each region, and then uses an implicit bandit strategy to decide which local optimization runs to continue.
+We used the original implementation and default (recommended) settings for the competing methods.
+In addition, we compared with two BO methods based on Bayesian neural networks (BNN).
+One uses Laplace approximation and the other Hamiltonian Monte-Carlo (HMC) sampling for posterior inference.
+We used a high-quality open-source implementation for each BNN based method.
 Table 3 summarizes the structural assumptions made by all the methods.
-We used the original implementation and default (recommended) settings for the competing methods.
-See details in Appendix Section A.3.
+The details of all the competing methods are given in Appendix Section A.2.
 For each optimization task, we randomly queried the target function for 20 times to generate the initial data.
@@
 For SaasBO, we performed tests using both NUTS and MAP estimation for surrogate training, denoted by SaasBO-NUTS and SaasBO-MAP, respectively.
-For standard BO, we trained the GP regression by maximizing the marginal likelihood using ADAM optimization (Kingma and Ba, 2014) (learning rate $0.1$ and 500 steps, as suggested in GPyTorch documents).
-We treated the mean function as a constant and jointly estimated it with the kernel parameters and noise variance.
-4.2 Results and Discussion
-4.2.1 Optimization Performance
-We present the ultimate optimization outcomes for synthetic benchmarks in Fig.
-1 and for real-world benchmarks in Fig.
-2.
-These are violin plots that illustrate the average of the obtained maximum function value over ten runs (the middle bar), the best (the top bar), the worst (the bottom bar), and the distribution of the obtained maximum (the shaded region).
-We can observe that on average, the standard BO with GP-ARD outperforms the competing high-dimensional BO methods in optimization every target function, often by a large margin, except that in the cases of Hartman6(300, 6) and NN-Prune, the performance with GP-ARD is slight worse (but still very close to) the best baseline(s).
-It is noteworthy that the performance of the competing methods appears to be sensitive to the structure within the target function.
+For a reliable comparison, we ran every method for ten times.
+4.2 Optimization Performance
+Ultimate Outcomes.
+We first examined the ultimate optimization outcomes of each method.
+The results are reported in Fig.
+1 for synthetic benchmarks and in Fig.
+2 for real-world benchmarks.
+Following [6], we used violin plots that illustrate the average of the obtained maximum function value over the ten runs (the middle bar), the best (the top bar), the worst (the bottom bar), and the distribution of the obtained maximum (the shaded region).
+We can observe that on average, the standard BO, including both SBO-1 and SBO-2, outperforms the competing methods at every target function, often by a large margin, except that in the cases of Hartman6(300, 6) and NAS201, the performance of SBO is slight worse than (but still very close to) the best baseline(s).
+It is noteworthy that the performance of the competing high-dimensional BO methods appears to be sensitive to the structure within the target function.
 For example, SaasBO imposes a strong inductive bias that only a small number of variables is related to the objective (via the usage of the horse-shoe prior).
@@
 As a result, SaasBO exhibits outstanding performance in both tasks.
-However, when the objective deviates from this assumption, as seen in cases like Ackley(150, 150), StybTang(200, 200), NN-Prune, and even Rosenbrock(300, 100), the performance of SaasBO experiences a dramatic drop.
+However, when the objective deviates from this assumption, as seen in cases like Ackley(150, 150), StybTang(200, 200) and even Rosenbrock(300, 100), the performance of SaasBO experiences a dramatic drop.
 Another example is RDUCB and Tree-UCB, which assume that the target function can be decomposed as a summation of a group of functions, each over a small number of variables.
@@
 However, their performance drops significantly in cases that do not align with their assumption, such as Ackley(150, 150) and Ackley(300, 150).
-In contrast, GP-ARD exhibits much greater robustness to the structure of the target function.
-The standard BO with GP-ARD consistently ranks among the best in all tasks, regardless of whether the target function depends only on a subset of variables, can be decomposed, etc.
-Interestingly, the standard BO shows superior performance even when using a single length-scale parameter for GP, which closely matches GP-ARD across most tasks, except for Ackley(300, 150), GP struggled in finding right locations to query, leading to relative poor performance.
-We also examined the runtime performance of each method.
-We report the maximum function value obtained throughout the optimization process in Fig.
+Across all the benchmarks, the BNN based methods performed worse than most GP based methods.
+This might be due to that under low data regimes, it is more challenging for a neural network to capture the shape of the target function and quantify the uncertainty, especially for high-dimensional targets.
+In contrast, the standard BO exhibits much greater robustness to the structure of the target function.
+Using standard GP equipped with Matérn kernels, both SBO-1 and SBO-2 consistently rank among the best in all tasks, regardless of whether the target function depends only on a subset of variables, can be decomposed, etc.
+Interestingly, as shown in Appendix Fig.7, even using a single length-scale parameter, the standard BO, i.e., SBO-3, still consistently shows superior performance, which closely matches ARD Matérn across most tasks, except for Ackley(300, 150), SBO-3 struggled in finding right locations to query, leading to poor performance.
+Runtime Performance.
+Next, we examined the runtime performance.
+For each method, we report the average maximum function value obtained at each step in Fig.
 5 and Fig.
 6 of the Appendix.
-It is noticeable that after a few initial steps, the standard BO with GP-ARD (and also with GP in most cases) consistently generated superior queries, finding larger function values.
-Its curves are mostly above those of the competing methods, indicating a more efficient optimization of the target function.
+We supplement the runtime standard deviation and the result of SBO-3 in Fig.
+8 of the Appendix.
+Note that since we have twelve baseline methods, reporting the standard deviations of all the methods makes the figure very cluttered.
+To ensure clarity, in Fig.
+8, we show the standard deviation of the best, worst, and median performed baselines so far at each step.
+One can see that after a few initial steps, SBO-1 and SBO-2 (and SBO-3 in most cases) consistently generated superior queries, finding larger function values.
+Their curves are mostly above those of the competing methods, indicating a more efficient optimization of the target function.
 This trend is particularly evident in examples such as Rosenbrock(100, 100), StybTang(200, 200), Mopta08, Rover, SVM, and more.
-The collective results demonstrate that the standard GP with ARD kernels is already effective for high-dimensional optimization.
-In comparison with using more complex surrogate models with additional assumptions, GP-ARD exhibits greater flexibility in adapting to the target function whose intrinsic structures are often agnostic in practice, and provides superior optimization results.
+The collective results demonstrate that the standard GP (with Matérn kernels) is effective for high-dimensional optimization.
+In comparison with more complex surrogate models using additional assumptions, the standard GP exhibits greater flexibility in adapting to the target function whose intrinsic structures are often agnostic in practice, and enables superior optimization results.
 Figure 1: Optimization performance in synthetic benchmarks. The results were summarized from 10 runs.
 Figure 2: Optimization performance in real-world problems. The results were summarized from 10 runs.
-4.2.2 Surrogate Learning
+4.3 Surrogate Learning
 Next, we explicitly evaluated the performance of the standard GP in learning high-dimensional functions to confirm if it can serve as a reliable surrogate model for high-dimensional BO.
-We sampled 300 training examples from all the six synthetic benchmark functions and four real-world benchmark tasks, including Mopta08 (124), Rover (60), DNA (180) and SVM (388).
+To this end, from the optimization trajectories of all the methods, we randomly sampled 300 training examples from the six synthetic benchmark functions and four real-world benchmark tasks, including Mopta08 (124), Rover (60), DNA (180) and SVM (388).
 We sampled another 100 examples for testing.
-We compared with the surrogate models used in SaasBO and ALEBO (with $d_{\text{emb}}=10$ and $d_{\text{emb}}=20$).
+We compared with the surrogate models used in SaasBO, namely, placing a horse-shoe prior over the the length-scale parameters to prune the variables.
+We also compared with a random projection GP, denoted as ProjGP, to align with the baselines employing low-dimensional embeddings.
+ProjGP employs the kernel function in the form of $\vartheta({\bf x},{\bf x}^{\prime})=\kappa({\bf A}{\bf x},{\bf A}{\bf x}^{\prime})$, where ${\bf A}$ is a random projection matrix.
+We varied the rank of ${\bf A}$ from 10 and 20.
+Each element is sampled from the standard normal distribution.
+In addition, we compared with an additive GP, denoted as AddGP, where the kernel takes the form $\vartheta({\bf x},{\bf x}^{\prime})=\sum_{j}\kappa({\bf x}^{j},{\bf x}^{\prime j})$ where each ${\bf x}^{j}\subset{\bf x}$ and ${\bf x}^{\prime j}\subset{\bf x}^{\prime}$, ${\bf x}$ and ${\bf x}^{\prime}$ are the union of ${\bf x}^{j}$’s and ${\bf x}^{\prime j}$’s, respectively.
+AddGP imposes a decomposition structure in the functional space; see (5).
+We tested AddGP with randomly partitioning the input into 10 variables per group.
+We tested two versions: one using non-overlapping partitions, and the other overlapping partitions with one variable shared by successive groups.
+We trained the standard GP with ARD Matérn-5/2 kernel and with a single length-scale Matérn-5/2 kernel using MLE.
 SaasBO was trained by NUTS.
-We repeated the experiment for five times, and evaluated the average test log-likelihood (LL), mean-square-error (MSE), and the standard deviation.
-The results are reported in Table 3 and Appendix Table 5.
-We can see that in the vast majority of cases, GP-ARD achieves the largest test LL and the smallest MSE.
-GP-ARD and GP together nearly always outperforms SaasBO, ALEBO-10 and ALEBO-20.
-Notably, GP-ARD surpasses ALEBO by a substantial margin, while SaasBO exhibits a closer prediction accuracy.
-In addition, we examined the alignment between the prediction and the ground-truth across the test examples from Mopta and DNA.
+For a fair comparison, ProjGP and AddGP were based on ARD Matérn-5/2 and trained via MLE as well.
+(a) ARD Matérn
+(b) Matérn
+(c) SaasBO-NUTS
+Figure 3: Alignment between the prediction (y-axis) and ground-truth function values (x-axis). The top row includes the results for Mopta08 (124) and the second row DNA (180).
+We repeated the experiment for five times, and evaluated mean-square-error (MSE), the average test log-likelihood (LL), and the standard deviation.
+Table 4 and Appendix Table 5 report the results for all the GP models.
+The best and second best results are highlighted with a bold font.
+As we can see, in every dataset, standard GP with Mateŕ kernels achieves either the best or the second best result, demonstrating excellent and robust performance in learning high-dimensional functions.
+In addition, we examined the alignment between the prediction and the ground-truth across the test examples from Mopta08 (124) and DNA (180).
 As illustrated in Fig.
-4, the prediction (as well the predictive standard deviation) of GP-ARD matches the ground-truth equally as well as SaasBO-NUTS, with a slight improvement over GP (using a single length-scale parameter).
+3, the prediction (and the predictive standard deviation) of ARD Matérn matches the ground-truth equally as well as SaasBO-NUTS, with a slight improvement over Matérn (using a single length-scale parameter).
 Together these results demonstrate that directly incorporating high-dimensional inputs into the kernel function does not necessarily compromise the learning performance.
-In other words, this implies that, even when the input space is high dimensional, commonly used kernels (e.g., the Matérn-5/2) remains expressive enough to capture the similarities in the space, allowing them to perform well in covariance modeling.
-On the contrary, enforcing (unnecessary) low-rank structures or sparsity might cause information loss, potentially detrimental to the learning performance.
-Our observation is consistent with the recent work in kernel operator learning (Batlle et al., 2024), where the standard kernel (GP) regression has demonstrated the state-of-the-art performance in learning function-to-function mappings.
+This implies that, even when the input space is high dimensional, the popular Matérn kernel remains expressive enough to capture the similarities in the space, allowing them to perform well in covariance modeling.
+On the contrary, enforcing (unnecessary) low-rank structures or sparsity might cause information loss, potentially detrimental to optimization performance.
+Our observation is consistent with the recent work in kernel operator learning [3], where the standard kernel (GP) regression has demonstrated the state-of-the-art performance in learning function-to-function mappings.
 In this context, the input comprises of a discretized function, typically of very high-dimensionality (e.g., several thousands).
-4.2.3 Computational Efficiency
-So far, in our BO loop, the GP model is trained via maximum likelihood estimation (MLE).
+4.4 Computational Efficiency
+So far, in our BO loop, the GP model is trained via MAP or MLE.
 Although the performance has been promising, it is worth investigating whether a full Bayesian approach would yield substantial improvements.
@@
 We conducted another round of BO experiments.
-For each step, we used NUTS to train GP-ARD.
-We adopted the same settings as used in SaasBO-NUTS (Eriksson and Jankowiak, 2021), where the mean parameter is assigned the standard Gaussian prior $\mathcal{N}(0,1)$, amplitude $\text{Gamma}(2,0.15)$, and noise variance $\text{Gamma}(0.9,10)$.
-We then assigned a uniform prior over each length-scale, $\text{Uniform}(0.001,30)$.
-The number of warm up steps was set to 512, and after the warm up stage, we ran 256 steps, and selected the samples at every 16-th step as the posterior samples.
+For each step, we used NUTS to train GP with ARD Matérn.
+We adopted the same settings as used in SaasBO-NUTS [6]: The number of warm up steps was set to 512, and after the warm up stage, we ran 256 steps, and selected the samples at every 16-th step as the posterior samples.
 We show the running and final optimization performance in Fig 4 , and Appendix Fig.
-7.
-It can be observed that while NUTS contributes to a slight improvement in several tasks, the enhancement is very marginal.
-However, running NUTS is much more computationally costly compared to MLE.
-For example, on NAS201 and DNA, NUTS takes 68x and 40x running time than MLE in each optimization step.
-Therefore, for practical purposes, using maximum likelihood estimation (MLE) for Bayesian Optimization with GP-ARD appears sufficient.
+9.
+It can be observed that while NUTS contributes to a slight improvement in several tasks, the enhancement is marginal.
+However, running NUTS is much more computationally costly compared to doing MAP and MLE.
+For example, on DNA, NUTS takes 59x and 40x running time of MAP and MLE, respectively, on average in each step.
+Therefore, for practical purposes, using MAP or MLE for GP training appears sufficient.
 It is computationally efficient, and the sacrifice of optimization performance is minimal compared to much more costly training methods like NUTS.
@@
 However, the computational cost of running SaasBO-NUTS is considerably higher, ranging from 5.4x (NAS201) to 33.9x (SVM, DNA) compared to the running time of SaasBO-MAP.
-This substantial increase in computational requirements may pose a severe bottleneck for practical usage.
-Detailed per-optimization-step running time for these methods are provided in Table 4 and Appendix Table 6.
-Task GP-ARD GP SaasBO ALEBO-10 ALEBO-20
-Mopta08(124) 0.613$\pm$0.156 -0.198$\pm$0.233 0.584$\pm$0.158 -1.176$\pm$0.080 -1.177$\pm$0.327
-Rover(60) -1.209$\pm$0.194 -0.908$\pm$0.068 -0.909$\pm$0.108 -1.270$\pm$0.094 -1.687$\pm$0.251
-DNA(180) 0.005$\pm$0.175 -0.459$\pm$0.142 -0.066$\pm$0.225 -1.346$\pm$0.108 -1.572$\pm$0.432
-SVM(388) 0.296$\pm$0.097 -0.090$\pm$0.065 0.160$\pm$0.113 -1.175$\pm$0.116 -1.077$\pm$0.148
-Ackley(150, 150) 0.720$\pm$0.092 0.899$\pm$0.058 0.809$\pm$0.053 -0.852$\pm$0.091 -0.266$\pm$0.091
-Ackley150(300, 150) 0.617$\pm$0.191 0.367$\pm$0.081 0.536$\pm$0.095 -0.870$\pm$0.097 -0.478$\pm$0.204
-Hartmann6 (300, 6) 1.681$\pm$0.110 -1.031$\pm$0.117 2.152$\pm$0.971 -1.433$\pm$0.249 -1.689$\pm$0.411
-Rosenbrock(100, 100) 0.602$\pm$0.069 0.172$\pm$0.099 0.553$\pm$0.084 -1.007$\pm$0.121 -0.854$\pm$0.355
-Rosenbrock(300, 100) 0.436$\pm$0.098 -0.256$\pm$0.071 0.265$\pm$0.115 -1.118$\pm$0.065 -1.380$\pm$0.329
-Stybtang(200, 200) 0.558$\pm$0.122 0.215$\pm$0.076 0.463$\pm$0.098 -1.050$\pm$0.099 -1.238$\pm$0.436
-Table 3: Test log likelihood of each surrogate model. The results were averaged from five runs.
-Mopta08 (124) Rover (60) NAS201 (30) DNA (180) SVM (388)
-GP-ARD 11.23 9.84 5.76 14.34 32.16
-GP-ARD-NUTS 492.62 404.17 393.62 575.03 768.29
-SaasBO-MAP 46.53 41.36 42.35 51.33 67.37
-SaasBO-NUTS 985.76 375.06 229.07 1739.56 2269.50
-Table 4: Average running time (in seconds) per optimization step in real-world benchmark tasks.
-Figure 3: Optimization performance when using GP-ARD trained with maximum likelihood estimation (MLE) and NUTS on SVM (388) and DNA (180). The results were summarized from 10 runs.
-(a) GP-ARD
-(b) GP
-(c) SaasBO-NUTS
-Figure 4: Alignment between the prediction (y-axis) and ground-truth function values (x-axis). The top row is the results for Mopta08 (124) and the second row DNA (180).
+The substantial increase in computational costs may pose a bottleneck for practical usage.
+Detailed per-optimization-step running time for these methods are provided in Appendix Table 6 and 7.
+Dataset ARD Matérn Matérn SaasBO Proj-GP Add-GP
+Mopta08(124) 0.034$\pm$0.006 0.145$\pm$0.057 0.038$\pm$0.008 0.461$\pm$0.157 0.107$\pm$0.033
+Rover(60) 0.327$\pm$0.078 0.284$\pm$0.09 0.284$\pm$0.09 0.548$\pm$0.097 0.321$\pm$0.097
+DNA(180) 0.149$\pm$0.038 0.332$\pm$0.053 0.152$\pm$0.049 0.747$\pm$0.121 0.312$\pm$0.061
+SVM(388) 0.034$\pm$0.01 0.103$\pm$0.017 0.034$\pm$0.014 0.392$\pm$0.071 0.077$\pm$0.012
+Ackley(150, 150) 0.017$\pm$0.008 0.007$\pm$0.001 0.007$\pm$0.001 0.034$\pm$0.007 0.007$\pm$0.001
+Ackley150(300, 150) 0.023$\pm$0.006 0.012$\pm$0.001 0.011$\pm$0.001 0.080$\pm$0.024 0.021$\pm$0.003
+Hartmann6 (300, 6) 0.017$\pm$0.013 0.419$\pm$0.048 0.024$\pm$0.023 0.583$\pm$0.094 0.416$\pm$0.084
+Rosenbrock(100, 100) 0.084$\pm$0.013 0.118$\pm$0.023 0.124$\pm$0.025 0.378$\pm$0.069 0.097$\pm$0.015
+Rosenbrock(300, 100) 0.090$\pm$0.064 0.160$\pm$0.09 0.131$\pm$0.075 0.533$\pm$0.214 0.138$\pm$0.075
+Stybtang(200, 200) 0.043$\pm$0.018 0.036$\pm$0.014 0.039$\pm$0.012 0.253$\pm$0.040 0.038$\pm$0.009
+Table 4: Mean Squared Error(MSE) of each surrogate model. The results were averaged from five runs.
+Figure 4: Optimization performance when using ARD Matérn trained with MAP and MLE, and NUTS on SVM (388) and DNA (180). The results were summarized from 10 runs.
+4.5 Ablation Study and Failure Modes with Other Settings
+Finally, we conducted a careful ablation study on the alternative settings to carry out standard BO.
+We found, however, those settings, such as using SE kernel and EI acquisition, often lead to inferior performance in high-dimensional optimization, which align with the a few failure cases reported in the literature [15, 20].
+The detailed results and analysis are given in Section B of the Appendix.
+We summarize the favorable and unfavorable settings for high-dimensional optimization in Table 1.
 5 Conclusion
-A longstanding belief suggests that standard Gaussian process is insufficient to model high-dimensional functions, and Bayesian optimization with standard GP is expected to exhibit poor performance when optimizing functions with more than 20 variables.
-However, through a comprehensive investigation across various benchmark tasks, this paper challenges and overturns the prevailing notion.
-Our findings reveal that standard GP excels in function estimation and uncertainty calibration, even when dealing with hundreds of input variables.
-Bayesian Optimization (BO) conducted with standard GP demonstrates remarkable performance in high-dimensional optimization, frequently outperforming existing high-dimensional BO methods by a substantial margin.
-Those methods, incorporating strong structural assumptions and employing more complex surrogate modeling, exhibit less robustness to various target functions, whose intrinsic structural properties are often agnostic in practice.
-Furthermore, the standard GP, benefiting from a parsimonious parameter space, can achieve satisfactory performance using only maximum likelihood estimation, thus eliminating the need for more computationally expensive Monte-Carlo methods in practical applications.
-Given these revelations, we advocate for a comprehensive exploration of the potential of the standard GP in high-dimensional optimization tasks.
-Appendix
-A More Experiment Details
+We have presented a comprehensive investigation on standard BO (SBO) in high-dimensional optimization.
+We discovered that simply using Matérn kernel and UCB, SBO has already achieved the state-of-the-art performance.
+There seems no need for imposing additional structure assumptions for function learning and/or placing strong priors over the model parameters.
+In the future, we will continue applying SBO in more tasks, and seek for theoretical justifications and guarantees.
+Appendix A More Experiment Details
 A.1 Definition of Synthetic Functions
@@
 where ${\bf x}\in[-2.048,2.048]^{D}$. We set $[c_{1},...,c_{D}]$ as an evenly spaced sequence in $[-2,2]$, where $c_{1}=-2$ and $c_{D}=2$. The optimum is at $x^{\dagger}=[c_{1}+1,...,c_{D}+1]$.
-Ackley. We used the Ackley function defined in (Surjanovic and Bingham, ),
+Ackley. We used the Ackley function defined in [30],
 $f({\bf x})=-20\text{ exp}\left(-0.2\sqrt{\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}}\right)-\text{exp}\left(\frac{1}{d}\sum_{i=1}^{d}\text{cos}(2{\pi}x_{i})\right)+20+\text{exp}(1),$
@@
 $f({\bf x})=-\sum_{i=1}^{4}{\alpha}_{i}{\exp}\left(-\sum_{j=1}^{6}A_{ij}(x_{j}-P_{ij})^{2}\right),$
-where each $x_{i}\in[0,1]$, ${\bf A}=[A_{ij}]$ and ${\bf P}=[P_{ij}]$ are two pre-defined matrices as in (Surjanovic and Bingham, ).
-The global optimum is at ${\bf x}^{\dagger}=[0.20169,0.150011,0.476874,0.275332,0.6573]$.
-Figure 5: Maximum Function Value Obtained vs. Number of Steps in synthetic benchmarks. The results were averaged from 10 runs.
-Figure 6: Maximum Function Value Obtained vs. Number of Steps in real-world benchmarks. The results were averaged from 10 runs.
-A.2 NN-Prune Details
-Given specific per-layer pruning rates ${\bf x}=[x_{1},\ldots,x_{151}]$, the pruning procedure is as follows.
-The weights in each layer $i$ are first ranked according to their magnitudes, and the bottom $x_{i}$ portion of the weights are removed.
-The pruned network is then fine tuned on Tiny ImageNet (Le and Yang, 2015).
-The prediction accuracy after pruning and fine tuning on a separate test dataset is used as the target function.
-To respect the global pruning rate constraint $\alpha$, we add to the objective a soft penalty term $-\lambda\cdot\max(\alpha-\beta,0)$, where $\beta$ is the global pruning rate reflected from ${\bf x}$.
-Note that the penalty is only incurred when $\alpha>\beta$.
-When $\alpha<\beta$, namely, the network is pruned to be sparser than the expected level, the test accuracy typically experienced a significant decline.
-Consequently, the optimization process automatically adjusts to increase $\beta$.
-We set $\lambda=8.0$ to introduce a strong penalty.
-In our experiment, we employed $\alpha=0.5$.
-A.3 Implementation
-A.4 Computing Environment
+where each $x_{i}\in[0,1]$, ${\bf A}=[A_{ij}]$ and ${\bf P}=[P_{ij}]$ are two pre-defined matrices as in [30]. The global optimum is at ${\bf x}^{\dagger}=[0.20169,0.150011,0.476874,0.275332,0.6573]$.
+Figure 5: Runtime performance: Maximum Function Value Obtained vs. Number of Steps in synthetic benchmarks. The results were averaged from 10 runs.
+Figure 6: Runtime performance: Maximum Function Value Obtained vs. Number of Steps in real-world benchmarks. The results were averaged from 10 runs.
+Figure 7: Ultimate optimization performance: Maximum Function Value obtained at the end of optimization loop, including SBO-3, namely training with a single length-scale parameter in Matérn-5/2 kernel.
+The results were collect from 10 runs.
+Figure 8: Runtime performance: Maximum Function Value obtained vs.
+Number of Steps in real-world benchmarks, including SBO-3, namely training with a single length-scale parameter in Matérn-5/2 kernel.
+The results were averaged from 10 runs.
+The shaded region depicts the standard deviation.
+A.2 Implementation
+  • ALEBO. We used ALEBO implementation shared by the Adaptive Experimentation (AX) Platform (version 0.2.2). The source code is at https://github.com/facebook/Ax/blob/main/ax/models/torch/alebo.py.
+  • HESBO. We used HESBO implementation of the original authors (https://github.com/aminnayebi/HesBO).
+  • TURBO. We used the original TURBO implementation (https://github.com/uber-research/TuRBO)
+  • SaasBO-NUTS. We used the original implementation of the authors (https://github.com/martinjankowiak/saasbo).
+  • SaasBO-MAP.
+The SaasBO implementation available to the public does not include the version using MAP estimation.
+We therefore implemented this method based on the original paper [6].
+All the hyperparameter settings follow exactly the same as the original paper.
+  • RDUCB and Tree UCB. The implementation of both methods is publicly available at https://github.com/huawei-noah/HEBO/tree/master/RDUCB.
+[…]
+A.3 Computing Environment
 We conducted the experimental investigations on a large computer cluster equipped with Intel Cascade Lake Platinum 8268 chips.
-For the NN-Prun problem, we ran the experiments on a Linux workstation, with a 16-core Intel 13900K CPU and one NVIDIA RTX 4090 GPU.
-Task GP-ARD GP SaasBO ALEBO-10 ALEBO-20
-Mopta08(124) 0.048$\pm$0.015 0.126$\pm$0.041 0.052$\pm$0.018 0.654$\pm$0.146 0.659$\pm$0.096
-Rover(60) 0.493$\pm$0.126 0.449$\pm$0.108 0.453$\pm$0.121 0.813$\pm$0.168 1.266$\pm$0.352
-DNA(180) 0.098$\pm$0.028 0.175$\pm$0.070 0.115$\pm$0.039 0.827$\pm$0.173 0.955$\pm$0.335
-SVM(388) 0.042$\pm$0.010 0.098$\pm$0.012 0.061$\pm$0.017 0.624$\pm$0.159 0.493$\pm$0.132
-Ackley(150, 150) 0.010$\pm$0.003 0.004$\pm$0.001 0.004$\pm$0.001 0.367$\pm$0.074 0.102$\pm$0.015
-Ackley(300, 150) 0.010$\pm$0.003 0.018$\pm$0.004 0.010$\pm$0.002 0.360$\pm$0.109 0.177$\pm$0.075
-Hartmann6(300, 6) 0.020$\pm$0.009 0.769$\pm$0.124 0.077$\pm$0.045 1.012$\pm$0.154 1.286$\pm$0.210
-Rosenbrock(100, 100) 0.049$\pm$0.006 0.124$\pm$0.020 0.062$\pm$0.008 0.601$\pm$0.095 0.807$\pm$0.248
-Rosenbrock(300, 100) 0.048$\pm$0.012 0.202$\pm$0.035 0.073$\pm$0.021 0.568$\pm$0.057 1.290$\pm$0.882
-Stybtang(200, 200) 0.040$\pm$0.014 0.078$\pm$0.012 0.044$\pm$0.014 0.502$\pm$0.066 0.861$\pm$0.329
-Table 5: Mean Squared Error (MSE) of each surrogate model. The results were averaged over five runs.
+Task ARD Matérn Matérn SaasBO Proj-GP Add-GP
+Mopta08(124) 0.570$\pm$0.08 -0.277$\pm$0.101 0.560$\pm$0.064 -0.962$\pm$0.183 -0.089$\pm$0.105
+Rover(60) -0.975$\pm$0.216 -0.647$\pm$0.158 -0.638$\pm$0.14 -0.995$\pm$0.116 -0.703$\pm$0.125
+DNA(180) -0.256$\pm$0.261 -0.666$\pm$0.102 -0.183$\pm$0.158 -1.137$\pm$0.143 -0.679$\pm$0.143
+SVM(388) 0.581$\pm$0.119 -0.179$\pm$0.042 0.650$\pm$0.12 -0.837$\pm$0.095 0.062$\pm$0.129
+Ackley(150, 150) 0.428$\pm$0.222 0.633$\pm$0.067 0.574$\pm$0.068 -0.006$\pm$0.268 0.569$\pm$0.066
+Ackley(300, 150) 0.232$\pm$0.155 0.252$\pm$0.057 0.279$\pm$0.077 -0.372$\pm$0.185 0.220$\pm$0.074
+Hartmann6(300, 6) 1.200$\pm$0.169 -0.625$\pm$0.079 1.528$\pm$0.564 -0.927$\pm$0.084 -0.617$\pm$0.118
+Rosenbrock(100, 100) 0.640$\pm$0.081 0.445$\pm$0.062 0.526$\pm$0.08 -0.466$\pm$0.182 0.521$\pm$0.057
+Rosenbrock(300, 100) 0.518$\pm$0.225 0.121$\pm$0.155 0.357$\pm$0.21 -0.706$\pm$0.209 0.189$\pm$0.2
+Stybtang(200, 200) 0.574$\pm$0.164 0.563$\pm$0.073 0.543$\pm$0.067 -0.458$\pm$0.069 0.582$\pm$0.072
+Table 5: Test log likelihood of each surrogate model. The results were averaged over five runs. Bolded are the best and second best results.
+Mopta08 (124) Rover (60) NAS201 (30) DNA (180) SVM (388)
+ARD Matérn-MAP 7.39 9.99 7.81 9.77 13.55
+ARD Matérn-MLE 11.23 9.84 5.76 14.34 32.16
+ARD Matérn-NUTS 492.62 404.17 393.62 575.03 768.29
+SaasBO-MAP 46.53 41.36 42.35 51.33 67.37
+SaasBO-NUTS 985.76 375.06 229.07 1739.56 2269.50
+Table 6: Average running time (in seconds) per optimization step in real-world benchmark tasks.
 Rosenbrock Ackley Stybtang Rosenbrock Ackley Hartmann6
 (100,100) (150, 150) (200, 200) (300, 100) (300, 150) (300, 6)
-GP-ARD 8.85 12.94 13.67 12.63 9.99 8.81
-GP-ARD-NUTS 857.21 850.23 1120.28 1276.42 1049.78 948.25
+ARD Matérn-MAP 8.52 36.26 21.42 16.36 25.90 51.56
+ARD Matérn-MLE 8.85 12.94 13.67 12.63 9.99 8.81
+ARD Matérn-NUTS 857.21 850.23 1120.28 1276.42 1049.78 948.25
 SaasBO-MAP 70.17 70.69 66.35 100.79 85.86 85.72
@@
 Table 7: Average running time (in seconds) per optimization step on synthetic benchmark tasks.
-Figure 7: Optimization performance when using GP-ARD trained with maximum likelihood estimation (MLE) and NUTS on the rest of benchmarks. The results were summarized from 10 runs.
+Figure 9: Optimization performance using MLE training, MAP training with a diffuse prior over the length-scale parameters, and NUTS on the rest of benchmarks. The results were summarized from 10 runs.
+Appendix B Ablation Study and Failure Modes With Other Settings
+In this section, we conducted a careful ablation study to investigate alternative popular settings in executing standard BO.
+From these settings, we identified a few failure modes during high-dimensional optimization.
+We provided analysis and discussions accordingly.
+B.1 GP Training
+First, we examined MLE training with LBFGS.
+We found that, however, LBFGS often incurs numerical issues, and cannot reliably finish model training.
+This is consistent with the report in other works [12].
+The issue might be because that without constraints (or regularization) over the parameters (such as the length-scales and amplitude), the exploration of LBFGS tends to be overly aggressive, which easily cause numerical problems.
+Hence, using stochastic optimization, like ADAM, might be a wise choice for MLE training.
+Second, we examined MAP training with alternative prior choices over the length-scale parameters.
+In the main paper, we use a diffuse prior, Uniform(0.001, 30).
+In practice, another commonly used choice is a concentrated Gamma prior, e.g., Gamma(3, 6), which is the default choice in the popular BO library BoTorch.
+To see how it is concentrated, in Fig.
+10 we compare its probability density curve with other priors, including the aforementioned uniform prior and the diffuse Gamma priors over the amplitude and noise variance.
+We found that employing a concentrated prior over the length-scale parameters will severely deteriorate the high-dimensional optimization performance.
+As shown in Fig 12 and 13, after placing the prior Gamma(3,6) over each length-scale parameter, both the ultimate and runtime optimization performance drops substantially across many tasks, such as Mopta08 (124), SVM (388) and Ackley(150, 150), no matter using ADAM or LBFGS to conduct training.
+To figure out the reason, we further investigated the surrogate learning performance.
+Following the same setting as in Fig.
+3, we examined the prediction alignment after applying the prior Gamma(3,6) over the length-scale parameters.
+As depicted in Fig.
+11 (third and fourth column), this concentrated prior leads to a severe prediction deviation from the ground-truth, indicating failure in learning the target function.
+This is consistent with the report in the SaasBO paper [6] (see Fig.1 of that paper).
+The quantitative results shown in Table 8 and 9 further confirm the substantially worse prediction accuracy as compared with GP training with a diffuse prior.
+In addition, we notice that the previous work[21] have also tested standard BO using Gamma(3,6) as the prior over the length-scale parameters.
+In their high-dimensional optimization experiments (see Fig.
+5 of that paper), standard BO performs badly.
+We used our setting (SBO-1 and SBO-2) to re-run one experiment that maximizes a high-dimensional polynomial function.
+It turns out that both SBO-1 and SBO-2 can surpass the best method I-BNN reported by that paper.
+Together these results have shown that a concentrated prior can severely limit GP’s capacity for high-dimensional function estimation, leading to inferior performance in high-dimensional optimization.
+Figure 10: Probability density curves. Gamma(2, 0.15) and Gamma(1.1, 0.5) are the priors for the amplitude and noise variance, respectively.
+Figure 11: Alignment between the prediction (y-axis) and ground-truth function values (x-axis).
+Task MLE MAP-Uniform(1E-3, 30) MAP-Gamma(3, 6) MAP-Gamma(3, 6)
+LBFGS ADAM
+Mopta08(124) 0.034$\pm$0.006 0.035$\pm$0.005 0.837$\pm$0.112 0.856$\pm$0.114
+Rover(60) 0.327$\pm$0.078 0.323$\pm$0.075 0.78$\pm$0.154 0.819$\pm$0.115
+DNA(180) 0.149$\pm$0.038 0.166$\pm$0.049 1.014$\pm$0.097 1.026$\pm$0.101
+SVM(388) 0.034$\pm$0.01 0.039$\pm$0.01 1.057$\pm$0.066 1.06$\pm$0.067
+Ackley(150, 150) 0.017$\pm$0.008 0.026$\pm$0.007 0.019$\pm$0.004 0.019$\pm$0.005
+Ackley(300, 150) 0.023$\pm$0.006 0.07$\pm$0.047 0.431$\pm$0.09 0.459$\pm$0.092
+Hartmann6(300, 6) 0.017$\pm$0.013 0.178$\pm$0.325 0.881$\pm$0.045 0.906$\pm$0.034
+Rosenbrock(100, 100) 0.084$\pm$0.013 0.081$\pm$0.012 0.9$\pm$0.156 0.925$\pm$0.172
+Rosenbrock(300, 100) 0.09$\pm$0.064 0.116$\pm$0.078 0.971$\pm$0.222 1.001$\pm$0.217
+Stybtang(200, 200) 0.043$\pm$0.018 0.054$\pm$0.024 0.698$\pm$0.05 0.732$\pm$0.034
+Table 8: Mean Squared Error(MSE) averaged from five runs. The best results are bolded.
+Task MLE MAP-Uniform(1E-3, 30) MAP-Gamma(3, 6) MAP-Gamma(3, 6)
+LBFGS ADAM
+Mopta08(124) 0.57$\pm$0.08 0.542$\pm$0.104 -1.337$\pm$0.058 -1.715$\pm$0.212
+Rover(60) -0.975$\pm$0.216 -1.062$\pm$0.286 -$1$1.322$\pm$0.073
+DNA(180) -0.256$\pm$0.261 -0.388$\pm$0.35 -1.426$\pm$0.054 -2.026$\pm$0.183
+SVM(388) 0.581$\pm$0.119 0.504$\pm$0.128 -1.453$\pm$0.038 -2.095$\pm$0.129
+Ackley(150, 150) 0.428$\pm$0.222 0.14$\pm$0.218 0.451$\pm$0.088 0.445$\pm$0.087
+Ackley(300, 150) 0.232$\pm$0.155 -0.257$\pm$0.528 -1.006$\pm$0.091 -1.037$\pm$0.087
+Hartmann6(300, 6) 1.2$\pm$0.169 0.67$\pm$0.867 -1.344$\pm$0.029 -1.618$\pm$0.128
+Rosenbrock(100, 100) 0.64$\pm$0.081 0.606$\pm$0.1 -1.308$\pm$0.083 -1.598$\pm$0.42
+Rosenbrock(300, 100) 0.518$\pm$0.225 0.347$\pm$0.325 -1.411$\pm$0.134 -1.967$\pm$0.413
+Stybtang(200, 200) 0.574$\pm$0.164 0.421$\pm$0.228 -1.251$\pm$0.017 -1.41$\pm$0.067
+Table 9: Test log likelihood averaged from five runs. The best results are bolded.
+Figure 12: Ultimate optimization performance using a diffuse prior, Uniform(0.001, 30), vs.
+a concentrated Gamma prior, Gamma(3,6), over the length scale parameters.
+SBO-1 and SBO-2 use Uniform(0.001, 30).
+Figure 13: Runtime optimization performance using a diffuse prior, Uniform(0.001, 30), vs.
+a concentrated Gamma prior, Gamma(3,6), over the length scale parameters.
+SBO-1 and SBO-2 use Uniform(0.001, 30).
+B.2 Alternative Acquisition Functions
+Next, we examined two other popular acquisition functions used in BO: Expected Improvement (EI) and Thompson sampling (TS).
+At each step, EI computes the expected improvement upon the best sample so far,
+$\text{EI}({\bf x})=(\mu({\bf x})-f({\bf x}^{*}))\Psi\left(\frac{\mu({\bf x})-f({\bf x}^{*})}{\sqrt{v({\bf x})}}\right)+\sqrt{v({\bf x})}\Phi\left(\frac{\mu({\bf x})-f({\bf x}^{*})}{\sqrt{v({\bf x})}}\right),$ (6)
+where $\Psi(\cdot)$ and $\Phi(\cdot)$ are the CDF and PDF of the standard Gaussian distribution, ${\bf x}^{*}$ is the best sample that gives the largest function value among all the queries up to the current step, $\mu({\bf x})$ and $\sqrt{v({\bf x})}$ are the posterior mean and standard deviation of $f({\bf x})$ given current training data.
+We used BOTorch to maximize EI through L-BFGS.
+To conduct TS at each step, we used Sobol Sequence [29] to sample 3K inputs in the domain, and jointly sampled the 3K function values from the GP posterior.
+We then selected the input with the largest function value as the next query.
+We denote this method by TS-3K.
+To explore more input candidates, we also employed the Lancos and conjugate gradient descent to approximate the posterior covariance matrix, and sampled 10K input candidates from Sobol Sequence.
+We denote this method as TS-10K.
+We used BOTorch implementation of both methods.
+We used ARD Matérn kernel and MLE training at each step.
+The ultimate and runtime optimization performance for using EI, TS-3K and TS-10K, are shown in Fig.
+14 and 15, respectively.
+As we can see, in most cases, such as Mopta08 (128), SVM (388), DNA (180) and Rosenbrock (100, 100), both EI and TS perform worse than UCB, and also worse than the best baseline.
+On Ackley(150, 150) and Ackley(300, 150), the performance of EI and TS are even among the worst.
+The limitation of EI might arise from the gradient vanishing during the optimization, as pointed by the recent work [1].
+A remedy is to use log-EI as the acquisition function.
+The failure of the TS might be due to the high-dimension of the input space.
+A few thousand candidates might not be sufficient to cover the whole space; as a result, the selected queries are inferior in indicating the best direction to optimize the target function.
+Figure 14: Ultimate performance with different acquisition functions. Note that SBO-1 and SBO-2 use UCB.
+Figure 15: Runtime performance with different acquisition functions. Note that SBO-1 and SBO-2 use UCB.
+B.3 Using Square Exponential (SE) kernels
+Third, we considered another widely used kernel in BO, namely, the SE kernel.
+We tested with ARD SE and SE kernel using a single length-scale parameter.
+We used MLE training and UCB $(\lambda=1.5)$ to run BO.
+As depicted in Fig.
+17 and 18, with smaller optimization dimensions, such as Mopta08 (124), NAS201 (3) and DNA (180), the performance of SE is among the best.
+However, with higher dimensions, e.g., above 200, the performance of SE drops substantially, such as in SVM (388) and Ackley(300, 150).
+We argue that the inferior performance might be due to that the vanilla SE kernel is not powerful enough to enable GP to well capture high-dimensional functions.
+Consider a single length-scale parameter.
+we observe that the value of the SE kernel is determined by the squared distance between the inputs,
+$\kappa_{\text{SE}}({\bf x},{\bf x}^{\prime})=a\exp\left(-\eta\|{\bf x}-{\bf x}^{\prime}\|^{2}\right).$ (7)
+With the growth of the input dimension $d$, the kernel function value decays much faster than the increase of the distance: a squared decay coupling with the exponential decay.
+The learning can be easily driven to adapt to a sparse, diagonal-dominant covariance matrix, which roughly treats each data point as independent.
+This can make the learning fail to capture the correlation between function values, and result in inferior learning performance.
+By contrast, the value of the Matérn kernel (see (1)) is dominated by the distance rather than the squared distance.
+Therefore, with the growth of the distance, the kernel function value decays much slower than the SE kernel.
+It is easier for the optimization to adjust the length-scale parameter so as to avoid “diagonal-like” covariance matrices.
+Consequently, the correlation between function values can be better captured in the high dimensional input space.
+To verify our argument, we examined the GP learning performance with ARD SE kernel and ARD Matérn kernel.
+We used Rosenbrock(d,d) and Hartmann6(d, 6) as test cases.
+The former is a “full-rank” function, namely, all the $d$ input variables are relevant to the function value.
+The latter has a “low-rank” structure in that only the first 6 variables out of the $d$ input variables determines the function value.
+For each test case, we increased the input dimension from 50 to 300.
+We randomly generated 500 training examples and 100 test examples.
+We then trained each method with MLE, and examined the mean square error (MSE) and the test log likelihood.
+We repeated the training and test procedure for five times, and report the average MSE, average test log likelihood, and the standard deviations in Fig.
+16.
+As we can see, in both test cases, when the input dimension $d$ does not exceed 150, the performance of using SE and Matérn-5/2 is close, in terms of both MSE and log likelihood.
+The two curves almost overlap.
+However, for $d>150$ in both test scenarios, the MSE significantly escalates when employing the SE kernel, accompanied by a substantial decrease in the test log likelihood.
+These observations suggest a failure of the SE kernel in effectively learning the target function.
+In comparison, employing the Matérn kernel maintains a consistently favorable prediction performance even when $d>150$.
+On the Rosenbrock dataset, the MSE gradually increases while the test log likelihood decreases, following a roughly linear trend.
+This outcome is expected as, with the rise in input dimension $d$, more training examples are necessary to sustain comparable performance.
+Furthermore, on the Hartmann6 dataset, little variation is observed in both MSE and test log likelihood with increasing the input dimension when using the Matérn kernel.
+This suggests that the GP has effectively discerned the pertinent input variables and mitigated the influence of irrelevant ones.
+In summary, these results collectively underscore the advantage of employing Matérn kernels for learning high-dimensional functions.
+It is worth noting that the most recent work [12] designs a log-normal prior over each length-scale parameter to overcome the learning issue with the ARD SE kernel.
+Figure 16: Learning high-dimension functions with different kernels. We trained on 500 points, and tested on 100 points. Results shown with mean and standard deviation over 5 runs.
+Figure 17: Ultimate optimization performance with different kernels. Note that SBO-1 and SBO-2 use Matérn-5/2 kernel.
+Figure 18: Runtime optimization performance with different kernels. Note that SBO-1 and SBO-2 use Matérn-5/2 kernel.
+B.4 Different Choice of $\lambda$ in UCB
+Finally, we examined how the exploration parameter $\lambda$ in UCB (see (4)) influences the performance in high-dimensional optimization.
+To this end, we used ARD Maérn kernel and varied $\lambda$ from {0.5, 1.0, 1.5, 2.0}.
+We conducted MLE training via ADAM optimization.
+We show the ultimate and runtime optimization performance in Fig.
+19 and 20, respectively.
+It can be seen that, in most benchmarks (Mopta08, Rover, NAS201, DNA, etc.), the performance with different $\lambda$’s is close.
+However, on DNA, SVM and Ackley, the performance of $\lambda=0.5$ is typically worse than the bigger choice of $\lambda$.
+This might be due to that those problems are more challenging (e.g., Ackley has many local minima), a larger exploration level ($\lambda$) is needed to identify better optimization outcomes.
+Figure 19: Ultimate optimization performance with different exploration level $\lambda$ in UCB. Note that SBO-1 and SBO-2 use $\lambda=1.5$.
+Figure 20: Runtime optimization performance with different exploration level $\lambda$ in UCB. Note that SBO-1 and SBO-2 use $\lambda=1.5$.

### v4

@@
 Abstract
-There has been a long-standing and widespread belief that Bayesian Optimization (BO) with standard Gaussian process (GP), referred to as standard BO, is ineffective in high-dimensional optimization problems.
-While this belief sounds reasonable, strong empirical evidence is lacking.
-In this paper, we systematically investigated BO with standard GP regression across a variety of synthetic and real-world benchmark problems for high-dimensional optimization.
-We found that, surprisingly, when using Matérn kernels and Upper Confidence Bound (UCB), standard BO consistently achieves top-tier performance, often outperforming other BO methods specifically designed for high-dimensional optimization.
-Contrary to the stereotype, we found that standard GP equipped with Matérn kernels can serve as a capable surrogate for learning high-dimensional functions.
-Without strong structural assumptions, BO with standard GP not only excels in high-dimensional optimization but also is robust in accommodating various structures within target functions.
-Furthermore, with standard GP, achieving promising optimization performance is possible via maximum a posterior (MAP) estimation with diffuse priors or merely maximum likelihood estimation, eliminating the need for expensive Markov-Chain Monte Carlo (MCMC) sampling that might be required by more complex surrogate models.
-In parallel, we also investigated and analyzed alternative popular settings in running standard BO, which, however, often fail in high-dimensional optimization.
-This might link to the a few failure cases reported in literature.
-We thus advocate for a re-evaluation and in-depth study of the potential of standard BO in addressing high-dimensional problems.
-The code for our experiment is available at https://github.com/XZT008/Standard-GP-is-all-you-need-for-HDBO.
+A long-standing belief holds that Bayesian Optimization (BO) with standard Gaussian processes (GP) — referred to as standard BO — underperforms in high-dimensional optimization problems.
+While this belief seems plausible, it lacks both robust empirical evidence and theoretical justification.
+To address this gap, we present a systematic investigation.
+First, through a comprehensive evaluation across eleven widely used benchmarks, we found that while the popular Square Exponential (SE) kernel often leads to poor performance, using Matérn kernels enables standard BO to consistently achieve top-tier results, frequently surpassing methods specifically designed for high-dimensional optimization.
+Second, our theoretical analysis reveals that the SE kernel’s failure primarily stems from improper initialization of the length-scale parameters, which are commonly used in practice but can cause gradient vanishing in training.
+We provide a probabilistic bound to characterize this issue, showing that Matérn kernels are less susceptible and can robustly handle much higher dimensions.
+Third, we propose a simple robust initialization strategy that dramatically improves the performance of the SE kernel, bringing it close to state-of-the-art methods, without requiring any additional priors or regularization.
+We prove another probabilistic bound that demonstrates how the gradient vanishing issue can be effectively mitigated with our method.
+Our findings advocate for a re-evaluation of standard BO’s potential in high-dimensional settings.
 1 Introduction
@@
 It is commonly thought that the number of optimization variables should not exceed 15 or 20, as beyond this threshold, BO is prone to failure [8, 23].
-This belief may partly stem from the intuition that, as a kernel method, GPs struggle to handle high-dimensional inputs for covariance modeling, rending them incapable of providing an expressive prior for high-dimensional functions.
-Kernel GP Training Acquisition Function
-Matérn ✓ L-BFGS/MAP/Diffuse prior ✓ UCB ✓
-ADAM/MAP/Diffuse prior ✓
-L-BFGS/MAP/Concentrated prior ✗ EI ✗
-SE ✗ ADAM/MAP/Concentrated prior ✗
-L-BFGS/MLE ✗ TS ✗
-ADAM/MLE ✓
-Table 1: Favorable and unfavorable settings of running standard BO for high-dimensional optimization.
-SE: square exponential; EI: expected improvement; TS: Thompson sampling; MLE: maximum likelihood estimation.
 A continuous line of research is dedicated to developing novel high-dimensional BO methods.
 The fundamental strategy is to impose strong structural assumptions into surrogate modeling so as to avoid directly dealing with high-dimensional inputs.
-One class of methods assumes a decomposition structure within the functional space [16, 26, 10, 34].
-In other words, the target function can be expressed as the summation of a group of low-dimensional functions, each operating over a small number of variables.
+One class of methods assumes a decomposition structure within the functional space [16, 24, 10, 33], where the target function is expressed as the summation of a group of low-dimensional functions, each operating over a small number of variables.
 Another family of methods assumes that the inputs are intrinsically low-rank.
@@
 Subsequently, GP surrogates are built with the reduced input dimensions.
-While the aforementioned beliefs and concerns may appear intuitive and reasonable, there lacks strong empirical evidence confirming that standard BO is indeed incapable of high-dimensional optimization.
-To bridge this gap, we systematically investigated BO with standard GP regression across a variety of synthetic and real-world benchmark problems.
-Surprisingly, we found that, with ARD Matérn kernels and UCB as the acquisition function, the performance of standard BO consistently ranks among the best and often surpasses existing methods by a substantial margin.
-Those methods are specifically designed for high-dimensional optimization.
-This leads us to call for a re-evaluation and in-depth study of the potential of standard BO for high-dimensional optimization.
-The major discoveries are summarized as follows.
-  • Optimization performance.
-We investigated BO with standard GP regression across eleven benchmarks widely used in literature, including six synthetic and five real-world high-dimensional optimization tasks.
-The number of optimization variables varies from 30 to 388.
-In comparison with six state-of-the-art high-dimensional BO methods, standard BO with ARD Matérn kernel and UCB consistently achieves the best (or close to the best) optimization performance across all the tasks.
-In contrast, the competing methods demonstrate strong performance when the target function aligns with their structural assumption.
-However, in cases of mismatch or when the structure of the target function is agnostic, such as those in real-world problems, the performance of the competing methods often experiences a significant drop.
-This underscores that BO with standard GP not only can excel in high-dimensional optimization but also is robust in accommodating various structures within the target functions.
-  • Surrogate Learning.
-We investigated the task of learning high-dimensional functions with standard GP regression.
-Across ten benchmark datasets, our findings reveal that the prediction of standard GP equipped with Matérn kernels aligns well with the ground-truth, mostly performing better than more complex GP models that induce low-rank structures.
-This observation is in line with the recent results in kernel operator learning [3], where inputs are discretized functions — typically high-dimensional (hundreds or thousands of dimensions).
-It confirms that standard GP can serve as a capable surrogate for high-dimensional functions.
-  • Computational Efficiency.
-We discovered that, with standard GP, achieving promising optimization performance is possible by using MAP estimation with diffuse priors or just maximum likelihood estimation.
-The marginalization of the parameters (kernel parameters and noise variance) via Markov-Chain Monte Carlo (MCMC) provides minor benefits while significantly increasing the training cost.
-In contrast, due to the usage of more complex surrogate models, training for alternative high-dimensional BO methods is often more expensive.
-Some of these methods heavily rely on extensive MCMC sampling to prevent overfitting, making the training of the surrogate model a computational bottleneck.
-  • Failure Modes with Other Settings.
-We found failure modes of standard BO with other popular settings, such as using vanilla Square-Exponential (SE) kernels, Expected-Improvement (EI) and Thompson sampling (TS) as acquisition functions, and MAP estimation with concentrated Gamma priors over length-scale parameters.
-We conducted a careful ablation study to reveal the failure of these choices in both optimization and function learning, providing analysis and discussions.
-In Table 1, we list the discovered favorable and unfavorable settings for high-dimensional optimization.
+While the aforementioned concerns may seem plausible, there is a lack of both strong empirical evidence and theoretical justification to confirm and explain why standard BO would be ineffective in high-dimensional optimization.
+To bridge this gap, we systematically investigated standard BO in this paper.
+The major contributions of our work are summarized as follows:
+  • Empirical Results.
+We investigated BO with standard GP across eleven widely-used benchmarks, including six synthetic and five real-world high-dimensional optimization tasks.
+The number of variables ranged from 30 to 388.
+We compared standard BO with nine state-of-the-art high-dimensional BO methods, and performed extensive evaluations and ablation studies.
+Surprisingly, while the popular SE kernel often led to poor performance, switching to the ARD Matérn kernel enabled standard BO to consistently achieve the best (or near-best) optimization performance across all the tasks.
+When successful, standard BO seems more flexible in accommodating various structures within the target functions.
+  • Theory.
+Through analyzing the gradient structure of the GP likelihood, we identified the primary failure mode of standard BO, in particular when using the SE kernel.
+Specifically, the commonly used initializations for the length-scale parameters, such as setting them to one, is improper in high-dimensional settings and can easily cause gradient vanishing, preventing effective training.
+We proved a probabilistic tail bound to characterize this issue under mild conditions.
+Applying the same analytical framework to the Matérn kernel, we show through a comparison of the tail bounds that the Matérn kernel is less prone to gradient vanishing with the same length-scale initialization, making it more effective for handling high-dimensional problems.
+  • Simple Robust Initialization.
+Based on our theoretical results, we proposed a simple yet robust length-scale initialization method, without requiring any additional priors or regularization.
+We proved a probabilistic bound showing that the probability of gradient vanishing decreases exponentially with the increase in dimensionality, implying that the gradient vanishing issue can be effectively mitigated in high-dimensional settings.
+Empirical evaluations demonstrate that our initialization method dramatically improves the performance of standard BO with SE kernels, enabling it to achieve, or come close to, the state-of-the-art performance in high-dimensional optimization.
 2 Standard Bayesian Optimization
-Consider maximizing a $d$-dimensional black-box objective function $f:\mathcal{X}\subset\mathbb{R}^{d}\rightarrow\mathbb{R}$.
+Consider maximizing a $d$-dimensional objective function $f:\mathcal{X}\subset\mathbb{R}^{d}\rightarrow\mathbb{R}$.
 The form of $f(\cdot)$ is unknown.
-We can only query the function value at any input ${\bf x}\in\mathcal{X}$, possibly with noises, and no gradient information is accessible.
-We aim to find ${\bf x}^{\dagger}=\operatornamewithlimits{argmax}_{{\bf x}\in\mathcal{X}}f({\bf x})$.
-To achieve this, Bayesian Optimization (BO) employs a probabilistic surrogate model to learn and predict $f(\cdot)$ across the domain $\mathcal{X}$, while also quantifying the uncertainty of the prediction.
+We can query the function value at any input ${\bf x}\in\mathcal{X}$, possibly with noises, but no gradient information is accessible.
+We aim to find ${\bf x}^{\dagger}=\mathrm{argmax}_{{\bf x}\in\mathcal{X}}f({\bf x})$.
+To achieve this, Bayesian Optimization (BO) employs a probabilistic surrogate model to predict $f$ across the domain $\mathcal{X}$, while also quantifying the uncertainty of the prediction.
 This information is then integrated to compute an acquisition function, which measures the utility of querying at any new input location given the current function estimate.
@@
 BO begins by querying at a few (typically randomly selected) input locations, and trains an initial surrogate model.
-Then the iterative procedure repeats until BO converges to the optimum or a predefined stopping criterion is met.
+Then the iterative procedure repeats until convergence or a predefined stopping criterion is met.
 The standard BO adopts GP regression [32] for surrogate modeling.
-Specifically, a GP prior is placed over the target function, $f\sim\mathcal{GP}\left(m({\bf x}),\kappa({\bf x},{\bf x}^{\prime})\right)$, where $m({\bf x})$ is the mean function, which is often set as a constant, and $\kappa(\cdot,\cdot)$ is the covariance function, which is often chosen as a Mercer kernel function.
-For example, a popular kernel used in BO is the Matérn-5/2 kernel with Automatic Relevance Determination (ARD),
+A GP prior is placed over the target function, $f\sim\mathcal{GP}\left(m({\bf x}),\kappa({\bf x},{\bf x}^{\prime})\right)$, where $m({\bf x})$ is the mean function, which is often set as a constant, and $\kappa(\cdot,\cdot)$ is the covariance function, which is usually chosen as a Mercer kernel function.
+One most popular kernel used in BO is the Square Exponential (SE) kernel with Automatic Relevance Determination (ARD),
+$\kappa_{\text{SE}}({\bf x},{\bf x}^{\prime})=a\exp(-\rho^{2}),$ (1)
+where $a>0$ is the amplitude, $\rho=\sqrt{({\bf x}-{\bf x}^{\prime})^{\top}{\rm diag}(\frac{1}{\bm{\ell}^{2}})({\bf x}-{\bf x}^{\prime})}$, and $\bm{\ell}=(\ell_{1},\ldots,\ell_{d})^{\top}>{\bf 0}$ are the length-scale parameters.
+We refer to this kernel as ARD because each input dimension has a distinct length-scale parameter.
+An alternative choice is the ARD Matérn-5/2 kernel,
 […]
-where $\tau=\sqrt{({\bf x}-{\bf x}^{\prime})^{\top}{\rm diag}(\frac{1}{\bm{\eta}})({\bf x}-{\bf x}^{\prime})}$, $a>0$ is the amplitude, and $\bm{\eta}>{\bf 0}$ are the length-scale parameters.
-We refer to this kernel as ARD because each input dimension has a distinct length-scale parameter.
-Alternatively, one can use a single length-scale parameter shared by all the input dimensions.
-Suppose we have collected a set of training inputs ${\bf X}=[{\bf x}_{1},\ldots,{\bf x}_{N}]^{\top}$ and (noisy) outputs ${\bf y}=[y_{1},\ldots,y_{N}]^{\top}$.
-Due to the GP prior, the function values at ${\bf X}$, denoted as ${\bf f}=[f({\bf x}_{1}),\ldots,f({\bf x}_{N})]^{\top}$, follow a multi-variate Gaussian distribution, $p({\bf f})=\mathcal{N}({\bf f}|{\bf m},{\bf K})$, where ${\bf m}=[m({\bf x}_{1}),\ldots,m({\bf x}_{N})]^{\top}$, ${\bf K}$ is the kernel matrix on ${\bf X}$, and each $[{\bf K}]_{ij}=\kappa({\bf x}_{i},{\bf x}_{j})$.
-Next, a Gauss noise model is used to model the likelihood of ${\bf y}$, namely, $p({\bf y}|{\bf f})=\mathcal{N}({\bf y}|{\bf f},\sigma^{2}{\bf I})$ where $\sigma^{2}$ is the noise variance.
-One can marginalize ${\bf f}$ out to obtain
-$p({\bf y}|{\bf X})=\mathcal{N}({\bf y}|{\bf m},{\bf K}+\sigma^{2}{\bf I}).$ (2)
-To estimate the GP parameters (e.g., $a$, $\bm{\eta}$, and $\sigma^{2}$), one can place a prior over these parameters and use MCMC to compute posterior samples or perform MAP estimation.
-One can also simply maximize (2) to perform MLE.
-The predictive distribution (given the GP parameters) has a closed form: at any new input ${\bf x}^{*}$, the predictive distribution is conditional Gaussian,
-$p\left(f({\bf x}^{*})|{\bf y}\right)=\mathcal{N}\left(f({\bf x}^{*})|\mu({\bf x}^{*}),v({\bf x}^{*})\right),$ (3)
-where $\mu({\bf x}^{*})=m({\bf x}^{*})+\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}({\bf y}-{\bf m})$, and $v({\bf x}^{*})=\kappa({\bf x}^{*},{\bf x}^{*})-\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}\kappa({\bf X},{\bf x}^{*})$.
-In each iteration, given the current GP surrogate, we maximize an acquisition function to identify the next input location at which to query.
-One commonly used acquisition function is the upper confidence bound (UCB), defined as
-$\text{UCB}({\bf x})=\mu({\bf x})+\lambda\sqrt{v({\bf x})},\;\;{\bf x}\in\mathcal{X},$ (4)
-where $\lambda$ represents the exploration level. There have been other popular acquisition functions, such as Expected Improvement (EI) [14], Thompson sampling (TS) [27], among others.
+Given training inputs ${\bf X}=[{\bf x}_{1},\ldots,{\bf x}_{N}]^{\top}$ and (noisy) outputs ${\bf y}=[y_{1},\ldots,y_{N}]^{\top}$, let us denote the function values at ${\bf X}$ as ${\bf f}=[f({\bf x}_{1}),\ldots,f({\bf x}_{N})]^{\top}$, which according to the GP prior, follow a multi-variate Gaussian distribution, $p({\bf f})=\mathcal{N}({\bf f}|{\bf m},{\bf K})$, where ${\bf m}=[m({\bf x}_{1}),\ldots,m({\bf x}_{N})]^{\top}$, ${\bf K}$ is the kernel matrix on ${\bf X}$, and each $[{\bf K}]_{ij}=\kappa({\bf x}_{i},{\bf x}_{j})$.
+Then one can employ a Gaussian likelihood for the observations ${\bf y}$, and obtain
+$p({\bf y}|{\bf X})=\mathcal{N}({\bf y}|{\bf m},{\bf K}+\sigma^{2}{\bf I}),$ (3)
+where $\sigma^{2}$ is the noise variance.
+To estimate the GP parameters, e.g., the length-scales $\bm{\ell}$, one can maximize the marginal likelihood or use MCMC sampling.
+The predictive distribution (given the GP parameters) at any new input ${\bf x}^{*}$, is a closed-form, conditional Gaussian distribution, $p\left(f({\bf x}^{*})|{\bf y}\right)=\mathcal{N}\left(f({\bf x}^{*})|\mu({\bf x}^{*}),v({\bf x}^{*})\right)$, where $\mu({\bf x}^{*})=m({\bf x}^{*})+\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}({\bf y}-{\bf m})$, and $v({\bf x}^{*})=\kappa({\bf x}^{*},{\bf x}^{*})-\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}\kappa({\bf X},{\bf x}^{*})$.
+In each iteration, given the current GP surrogate model, an acquisition function is maximized to identify the next input location at which to query.
+One commonly used acquisition function is the upper confidence bound (UCB), defined as $\text{UCB}({\bf x})=\mu({\bf x})+\lambda\sqrt{v({\bf x})},\;\;{\bf x}\in\mathcal{X}$, where $\lambda$ represents the exploration level.
+There have been other popular acquisition functions, such as Expected Improvement (EI) [15], Thompson sampling (TS) [25], the recently proposed log-EI [1], among others.
 3 High Dimensional Bayesian Optimization
-An enduring and widespread belief is that when the input dimension $d$ is high (e.g., a few hundreds), the standard BO is prone to failure.
-This belief might partly arise from an intuition that commonly used kernels, such as in (1), could encounter challenges in dealing with high-dimensional inputs, making GP struggle in capturing the target function.
-As a result, there is a dedicated line of research aimed at developing novel high-dimensional BO methods.
-The key idea of these methods is to introduce strong structural assumptions in surrogate modeling to sidestep directly handling the high-dimensional inputs in kernels.
-3.1 Structural Assumption in Functional Space
-The first class of methods assumes a decomposition structure within the functional space. Specifically, the target function $f$ is modeled as
+An enduring and widespread belief is that when the input dimension $d$ is high, e.g., a few hundreds, the standard BO is prone to failure.
+This belief might partly arise from an intuition that commonly used kernels, such as (1), could encounter challenges in dealing with high-dimensional inputs, making GP struggle in capturing the target function.
+A dedicated line of research is developing novel high-dimensional BO methods.
+The key idea of these methods is to impose some structural assumptions in surrogate modeling to sidestep directly handling the high-dimensional inputs in kernels.
+Structural Assumption in Functional Space. The first class of methods assumes a decomposition structure within the functional space. Specifically, the target function $f$ is modeled as
 $f({\bf x})=\sum\nolimits_{j=1}^{M}f_{j}({\bf x}^{j}),\;\;f_{j}\sim\mathcal{GP}\left(m_{j}({\bf x}^{j}),\kappa_{j}(\cdot,\cdot)\right),$ (4)
@@
 Han et al., [10] proposed Tree-UCB, which restricted the dependency graph to be tree-structured so as to boost the efficiency of structure learning.
-However, the most recent work [34] points out that, learning the group structure through relatively small data can be misleading.
+However, the more recent work [33] points out that, learning the group structure through relatively small data can be misleading.
 A wrongly learned structure can cause queries to be stuck at local optima.
-The work proposes RDUCB that randomly decomposed ${\bf x}$ into a tree structure, and surprisingly works better than those learned structures through various methods.
-3.2 Structural Assumption in Input Space
+The work proposes RDUCB that randomly decomposes ${\bf x}$ into a tree structure, and surprisingly works better than those learned structures through various methods.
+Structural Assumption in Input Space.
 Another class of methods assumes a low-rank structure in the input space.
-Low-dimensional Embedding.
+Many of them introduce Low-dimensional Embeddings.
 Wang et al., [31] proposed a method named REMBO, building the GP surrogate in a low-dimensional embedding space, $\mathcal{Z}\subset\mathbb{R}^{d_{\text{emb}}}$ where $d_{\text{emb}}\ll d$.
@@
 But when maximizing the acquisition function, ALEBO incorporates a constraint that ${\bf A}{\bf z}$ must be in $\mathcal{X}$, thereby avoiding the clipping issue.
-Partial Variable Dependency.
-Another strategy is to trim down the input variables.
+Another direction is to impose Partial Variable Dependency by triming down the input variables.
 The recent start-of-the-art, SaasBO [6], uses a horse-shoe prior [4] over the length-scale parameter for each variable.
@@
 While one can also use MAP estimation, the performance is often much worse.
+4 Theoretical Analysis
+To understand why standard BO can fail in high-dimensional cases, in particular with the SE kernel (see Section 6), we look into the gradient structure during the GP training.
+Specifically, we consider at the beginning, all the length-scale parameters are set to the same initial value, $\ell_{1}=\ldots=\ell_{d}=\ell_{0}$.
+From the marginal likelihood (3), we can derive the gradient w.r.t each length-scale,
+$\frac{\partial\log p({\bf y}|{\bf X})}{\partial\ell_{k}}=\frac{1}{2}{\rm tr}\left({\bf A}\cdot\frac{\partial{\bf K}}{\partial\ell_{k}}\right),$ (5)
+where ${\bf A}=\bm{\alpha}\bm{\alpha}^{\top}-({\bf K}+\sigma^{2}{\bf I})^{-1}$, and $\bm{\alpha}=({\bf K}+\sigma^{2}{\bf I})^{-1}({\bf y}-{\bf m})$.
+When using the SE kernel as specified in (1), each entry of $\frac{\partial{\bf K}}{\partial\ell_{k}}$ is as follows: $\left[\frac{\partial{\bf K}}{\partial\ell_{k}}\right]_{ii}=0$ and for $i\neq j$,
+$\left[\frac{\partial{\bf K}}{\partial\ell_{k}}\right]_{ij}=\frac{\partial\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})}{\partial\ell_{k}}=\frac{2a}{\ell_{k}}\exp(-\frac{\|{\bf x}_{i}-{\bf x}_{j}\|^{2}}{\ell_{0}^{2}})\frac{(x_{ik}-x_{jk})^{2}}{\ell_{k}^{2}}\leq\frac{2a}{\ell_{k}}\cdot\frac{\rho^{2}}{e^{\rho^{2}}},$ (6)
+where $x_{ik}$ and $x_{jk}$ are the $k$-th element in ${\bf x}_{i}$ and ${\bf x}_{j}$, respectively, and $\rho=\frac{\|{\bf x}_{i}-{\bf x}_{j}\|}{\ell_{0}}$.
+Intuitively, as the input dimension $d$ increases, the squared distance between input vectors grows rapidly.
+Consequently, the factor $\frac{\rho^{2}}{e^{\rho^{2}}}$ in (6) can quickly fall below the machine epsilon $\xi$, causing the gradient vanishing issue [3, 11].
+That is, the length-scale cannot be effectively updated according to $\partial\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})/\partial\ell_{k}$, due to the limitations of the numerical precision.
+Proposition 4.1.
+Given any $\xi>0$, $\frac{\rho^{2}}{e^{\rho^{2}}}<\xi$ when $\rho>\tau_{\text{SE}}=\frac{1}{2}+\sqrt{\frac{1}{4}-\log\xi}$.
+When using the double-precision floating-point format (float64), the rounding machine epsilon is $\xi=2^{-53}$, and the above threshold is $\tau_{\text{SE}}=6.58$.
+To analyze the likelihood of gradient vanishing, we provide a probabilistic bound under uniform data distribution.
+Lemma 4.2.
+Suppose the input domain is $[0,1]^{d}$ and the input vectors are sampled independently from the uniform distribution, then for any constant threshold $\tau>0$, when $d>6\ell_{0}^{2}\tau^{2}$,
+$p(\rho\geq\tau)>1-2\exp\left(-\frac{(d-6\ell_{0}^{2}\tau^{2})^{2}}{18d}\right).$ (7)
+The lower bound grows exponentially with the input dimension $d$ and converges to one.
+This implies that given any fixed choice of $\ell_{0}$ relevant to $d$ (e.g., $\ell_{0}=1$), as $d$ increases, the probability $p(\rho\geq\tau_{\text{SE}})$ will rapidly approach one, leading to vanishing gradient for every $\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})$.
+As a consequence, each length scale $\ell_{k}$ in (5) cannot be effectively updated, preventing meaningful training [^1] A union bound can be directly derived from (7) to lower bound the joint probability that all gradients $\{\partial\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})/\partial\ell_{k}\}_{i,j}$ across $N$ training instances are below machine epsilon.
+However, this bound is too loose to reflect our practical observation (see Fig.
+1 and  2 and Table 2).
+Therefore, we focus our discussion on a single pair of inputs.
+.
+The inferior prediction accuracy of the GP model further contributes to poor performance of the BO procedure.
+Since $\tau_{\text{SE}}$ is small, gradient vanishing can occur even when $d$ is not large.
+For example, for $\ell_{0}=0.5$ and $\ell_{0}=1$, we have $p(\rho>\tau_{\text{SE}})>0.99$ when $d\geq 205$ and $d\geq 473$, respectively.
+Why Matérn kernels perform much better?
+In our practical evaluations (see Section 6), standard BO with Matérn kernels (2) often perform much better than with the SE kernel in high-dimensional problems, frequently achieving state-of-the-art results.
+To understand the reasons behind this improvement, we apply the same framework to analyze the behavior of GP training with the Matérn kernel.
+First, we obtain the gradient,
+[…]
+As the distance increases, the factor $\rho^{2}/e^{\sqrt{5}\rho}$ will also converge to zero, implying that with the growth of $d$, the gradient will eventually fall below machine epsilon.
+However, since this factor $\rho^{2}/e^{\sqrt{5}\rho}$ decreases much slower than the corresponding factor to the SE kernel, namely $\rho^{2}/e^{\rho^{2}}$ in (6), the Matérn kernel is able to robustly handle much higher dimensions.
+Proposition 4.3.
+[…]
+[…]
+Lower bound 0.95 0.99 0.995 0.999 0.9995 0.9999
+SE ($d$) 172 205 219 250 264 294
+Matérn ($d$) 980 1040 1064 1116 1137 1185
+Table 1: The input dimension $d$ vs. lower bound probability of gradient vanishing under $\ell_{0}=0.5$.
+Numerical Verification.
+To verify whether gradient vanishing occurs during practical GP training and to access whether our analysis aligns with the empirical observations, we conducted numerical experiments using two synthetic functions, Hartmann6($d$, 6) and Rosenbrock(d,d) (see Section 6.1 for definitions).
+We varied the input dimension $d$ across {50, 100, 200, 300, 400, 500, 600}, with the domain set to $[0,1]^{d}$.
+For each value of $d$, we uniformly sampled $500$ training inputs and $100$ test inputs.
+We trained GP with both SE and Matérn kernels, using length-scale initialization $\ell_{0}$ from $\{0.1,0.5,0.693,1.0,\sqrt{d}\}$.
+Note $\ell_{0}=0.693=\text{SoftPlus}(0)$ is a popular initialization choice, used as the default choice in influential GP/BO libraries GPyTorch [9] and BOTorch [2].
+Training and testing were repeated 20 times for each $d$ and $\ell_{0}$.
+We then evaluated (1) the average relative $L_{2}$ difference between the length-scale vector before and after training, namely, $\|\bm{\ell}_{\text{trained}}-\bm{\ell}_{\text{init}}\|/\|\bm{\ell}_{\text{init}}\|$, (2) the average gradient norm at the first training step, and (3) the average test Mean-Squared-Error (MSE).
+As shown in Fig.
+1, gradient vanishing indeed occurred, leading to training failure.
+When using $\ell_{0}=0.1$ for GP training with the SE and Matérn kernels, both the length-scale relative $L_{2}$ difference and the gradient norm remained close to zero across all choices of $d$ (those values are not displayed because they are too small).
+Consequently, the MSE hovered around 1.0, indicating consistent training failure.
+When we increased $\ell_{0}$ to 0.5 and 1.0 for training with the SE kernel, both the length-scale difference and gradient norm became significantly larger and non-trivial for $d=50$ and $d=100$, resulting in MSE below 0.2.
+This demonstrates that gradient vanishing did not occur and training was successful.
+However, as $d$ increased, gradient vanishing reappeared at $d=200$ for $\ell_{0}=0.5$ and at $d=400$ for $\ell_{0}=1.0$, for both Hartmann6 and Rosenbrock.
+In these cases, the length-scale difference and gradient norm dropped near zero, while the MSE returned to around 1.0.
+A similar trend was observed with the Matérn kernel.
+However, GP training with the Matérn kernel can succeed with higher dimensions under the same initialization conditions.
+For example, with $\ell_{0}=0.5$, for both Hartmann6 and Rosenbrock, training with the Matérn kernel started to failed until $d=400$, compared to $d=200$ with the SE kernel.
+When using $\ell_{0}=1.0$, the Matérn kernel succeeded across all $d$ values, while the SE kernel started to fail at $d=400$.
+Table 2 summarizes the input dimensions at which gradient vanishing and training failure began to happen.
+Confirmation in BO Running.
+We next investigated whether gradient vanishing occurs during the BO process, potentially leading to poor optimization performance.
+To this end, we examined two benchmark BO tasks: SVM (D=388) and Rosenbrock (300,100), as detailed in Section 6.
+As shown in Fig.
+3 and 4, standard BO with the Matérn kernel (SBO-Matérn) achieved top-tier results, while BO with the SE kernel (SBO-SE) performed poorly.
+Both methods used the initialization $\ell_{0}=0.693$.
+We then extracted the BO trajectories for both methods and analyzed the relative $L_{2}$ difference in the length-scale vectors before and after training, and the gradient norm at the first training step.
+As illustrated in Fig.
+2, during the BO iterations, GP-SE ($\ell_{0}=0.693$) consistently experienced gradient vanishing and so training failure, whereas GP-Matérn ($\ell_{0}=0.693$) maintained non-trivial gradients.
+This difference directly correlates with their respective optimization performances — poor for SBO-SE and excellent for SBO-Matérn.
+Figure 1: Mean Square Error (MSE), Relative $L_{2}$ difference between the length-scale vectors before and after training, and $L_{2}$ norm of the length-scale gradient at the first training step, across different input dimensions ($x$-axis) and initializations (legend).
+The results were averaged from 20 runs.
+For $\ell_{0}=0.1$, the length-scale relative difference and gradient norm are not displayed, as they fall much below the minimum values on the $y$-axis.
+Figure 2: The relative $L_{2}$ difference of the length-scale vectors before and after training and the gradient norm at the first training iteration across every step of BO.
+The results for GP-SE ($\ell_{0}=0.693$) are not displayed, as they fall much below the minimum values on the y-axis.
+Hartmann6 $\ell_{0}=0.1$ 0.5 0.6931 1.0 $\sqrt{d}$
+SE 50 200 200 400
+Matérn 50 400 600
+Rosenbrock $\ell_{0}=0.1$ 0.5 0.6931 1.0 $\sqrt{d}$
+SE 50 200 200 400
+Matérn 50 400 600
+Table 2: The input dimension from which gradient vanishing started to occur and training started to fail; ✓means no gradient vanishing occurred.
+5 Robust Initialization
+From the analysis in Section 4, one straightforward approach to improving learning with the SE (and also Matérn) kernel is to initialize each length-scale with a larger $\ell_{0}$, which can loosen the lower bound in (7) and allow the GP model to better adapt to higher dimensions.
+However, this approach is not robust enough, as $d-6\ell_{0}^{2}\tau^{2}$ will continue to increase with $d$, and the bound will still rise toward one rapidly, causing the gradient to vanish.
+To address this issue, we propose setting $\ell_{0}$ to growing with $d$ rather than keeping it constant,
+$\ell_{0}=c\sqrt{d},\;\;\;c>0.$ (8)
+Lemma 5.1.
+Suppose the input domain is $[0,1]^{d}$ and each input vector is independently sampled from the uniform distribution.
+Given any constant threshold $\tau>0$, we set $\ell_{0}=c\sqrt{d}$ such that $c>\frac{1}{\sqrt{6}\tau}$, then
+$p(\rho\geq\tau)\leq 2\exp\left(-2(c^{2}\tau^{2}-\frac{1}{6})^{2}d\right)\propto\exp(-\mathcal{O}(d)).$ (9)
+With this initialization, the increase in $d$ exponentially reduces the upper bound on the probability of $\rho$ exceeding the given threshold $\tau$.
+As a result, the gradient vanishing issue can be fundamentally mitigated in the high-dimensional settings.
+In our experiments, we tested standard BO using SE kernel with our proposed length-scale initialization as specified in (8) where we set $c=1$.
+As shown in Fig.
+3 and Fig.
+4, the performance across all the benchmarks is dramatically improved, as compared to using popular initialization $\ell_{0}=0.693$.
+The performance matches that of using Matérn kernels, achieving state-of-the-art (or near-best) results.
+In addition, as shown in Fig.
+1 and Table 2, our numerical experiments confirm that gradient vanishing never occurred with our robust initialization method.
+Fig.
+2 further demonstrates that our method eliminates gradient vanishing during practical BO runs, enabling effective GP training and much better optimization results.
+Together this shows that our method can effectively adapt SE kernels to high-dimensional settings, and mitigate the failure mode of standard BO.
+Alternative method.
+Recent work [13] proposes an intuitive approach to address the learning challenge of the SE kernel in the high-dimensional spaces by constructing a log-normal prior over each length scale, $p(\ell_{k})=\text{LogNormal}(\ell_{k}|\mu_{0}+\frac{\log(d)}{2},\sigma_{0})$.
+This prior regularizes the length-scale to be on the order of $\sqrt{d}$ during training.
+Using our theoretical framework, this method can be justified as an alternative strategy to alleviate the gradient vanishing issue.
+However, our method, directly motivated by our analysis, is even simpler — requiring no additional prior construction or associated parameters, and has also shown effective in addressing the gradient vanishing issue.
+The comparison with  [13] is given in Section 6.
 6 Comprehensive Evaluation
-We now examine and analyze the performance of standard BO in high-dimensional optimization. To this end, we conducted a series of systematic experiments.
 6.1 Experimental Settings
-We employed eleven benchmarks, of which six are synthetic and five are real-world problems. For clarity in the discussion, all the tasks aim for function maximization.
+For a comprehensive evaluatio of the standard BO, we employed eleven benchmarks, of which six are synthetic and five are real-world problems. For clarity, all the tasks aim for function maximization.
 $\text{Fun}(d,d^{\prime})$ Structure
@@
 We considered four popular synthetic functions: Ackley, Rosenbrock, Hartmann6, and Stybtang.
-Definitions of these functions are provided in Section A.1 of the Appendix.
+Definitions of these functions are provided in Section B.1 of the Appendix.
 Each task is represented by “$\text{Fun}(d,d^{\prime})$”, where $d$ is the input dimension that BO methods optimize for, and $d^{\prime}$ is the number of effective variables ($d^{\prime}\leq d$).
@@
 The benchmarks and the structures within the target functions are summarized in Table 3.
-Real-World Benchmarks. We employed the following real-world benchmark problems.
-  • Mopta08 (124)  [13], a vehicle design problem.
+Real-World Benchmarks.
+We employed the following real-world benchmark problems: Mopta08 (124)  [14], a vehicle design problem.
 The objective is to minimize the mass of a vehicle with respect to 124 design variables, subject to 68 performance constraints.
@@
 SVM (388) [6], a hyper-parameter tuning problem that optimizes 3 regularization parameters and 385 length-scale parameters for a kernel support vector machine regression model.
-  • Rover (60), a rover trajectory planning problem from [31]. The goal is to find an optimal trajectory determined by the location of 30 waypoints in a 2D environment.
+Rover (60), a rover trajectory planning problem from [30].
+The goal is to find an optimal trajectory determined by the location of 30 waypoints in a 2D environment.
 DNA (180) [26], a hyper-parameter optimization problem that optimizes 180 regularization parameters for weighted lasso regression on an DNA dataset [20].
@@
 NAS201 (30)  [5], a neural architecture search problem on CIFAR-100 dataset.
-Method Structural Assumption
-ALEBO, HESBO Low-dim embedding
-SaasBO Partial variable dependency
-RDUCB, Tree UCB Additive function decomposition
-SBO, TURBO, BNN None
-Table 3: BO methods and their structural assumption about the target function.
-Methods.
-To execute the standard BO (SBO), we used GP regression with ARD Matérn-5/2 kernel (see (1)).
+Methods. We tested with three versions of standard BO:
+  • SBO-Matérn: GP regression with ARD Matérn-5/2 kernel as specified in (2), the length-scale initialization $\ell_{0}=\text{SoftPlus}(0)=0.693$, as the default choice in the influential GP/BO libraries GPyTorch and BOTorch.
+  • SBO-SE: GP regression with ARD SE kernel as specified in (1), the length-scale initialization $\ell_{0}=\text{SoftPlus}(0)=0.693$.
+  • SBO-SE(RI): GP regression with ARD SE kernel, using our proposed robust initialization $\ell_{0}=c\sqrt{d}$ where $d$ is the input dimension and $c=1$
 We used UCB as the acquisition function where the exploration level $\lambda$ was set to 1.5.
-We used GPyTorch [9] for surrogate training.
-To compare with other high-dimensional BO methods, we considered two training schemes.
-One performs MAP estimation with LBFGS.
+We used GPyTorch for GP training.
+To expose the possible gradient vanishing issue , we used Uniform(0.001, 30) over each length-scale.
+The noise variance was assigned a diffused Gamma prior, Gamma(1.1, 0.05), and the amplitude the prior Gamma(2.0, 0.15).
+We ran L-BFGS to obtain MAP estimation.
 The maximum number of iterations was 15K and tolerance level 2E-9.
-We employed a diffuse prior over each length-scale, i.e., Uniform(0.001, 30).
-The noise variance was assigned a diffuse Gamma prior, Gamma(1.1, 0.05), and the amplitude the prior Gamma(2.0, 0.15).
-We denote this method as SBO-1.
-The second training scheme simply uses ADAM optimization [17] to perform MLE, with learning rate $0.1$ and the maximum number of epochs was set to $500$.
-We denote this method as SBO-2.
-We also tested with Matérn-5/2 kernel with a single length-scale parameter, which we denoted as SBO-3.
-We used BOTorch [2] for acquisition function maximization, which uses LBFGS.
-Additional evaluations on alternative choices of the kernel, training scheme, and acquisition function are provided in Appendix Section B.
 We compared with six state-of-the-art high-dimensional BO methods, including Tree UCB [10], RDUCB [33], HESBO [23], ALEBO [18], SaasBO [6], and TURBO [7].
@@
 TURBO searches for trust-regions in the input domain, learns a set of local GP models over each region, and then uses an implicit bandit strategy to decide which local optimization runs to continue.
+In addition, we tested the most recent method VBO (Vanilla BO[13] that levels up the performance of the standard BO by constructing a log-normal prior over the length-scale parameters (see Section 5).
 We used the original implementation and default (recommended) settings for the competing methods.
-In addition, we compared with two BO methods based on Bayesian neural networks (BNN).
+Besides, we compared with two BO methods based on Bayesian neural networks (BNN).
 One uses Laplace approximation and the other Hamiltonian Monte-Carlo (HMC) sampling for posterior inference.
 We used a high-quality open-source implementation for each BNN based method.
-Table 3 summarizes the structural assumptions made by all the methods.
-The details of all the competing methods are given in Appendix Section A.2.
-For each optimization task, we randomly queried the target function for 20 times to generate the initial data.
-We then ran each method for 400 steps for synthetic tasks and for 300 steps for real-world tasks.
+Appendix Table 6 summarizes the structural assumptions made by all the methods.
+The details of all the competing methods are given in Appendix Section B.2.
+For each optimization task, we randomly queried the target function for 20 times to collect the initial data.
 We tested TURBO with a single trust region and five trust regions, denoted by TURBO-1 and TURBO-5 respectively.
@@
 4 for real-world benchmarks.
+Due to space limit, we supplement the BNN-based BO results in Appendix Section C.1.
 Following [6], we used violin plots that illustrate the average of the obtained maximum function value over the ten runs (the middle bar), the best (the top bar), the worst (the bottom bar), and the distribution of the obtained maximum (the shaded region).
-We can observe that on average, the standard BO, including both SBO-1 and SBO-2, outperforms the competing methods at every target function, often by a large margin, except that in the cases of Hartman6(300, 6) and NAS201, the performance of SBO is slight worse than (but still very close to) the best baseline(s).
-It is noteworthy that the performance of the competing high-dimensional BO methods appears to be sensitive to the structure within the target function.
-For example, SaasBO imposes a strong inductive bias that only a small number of variables is related to the objective (via the usage of the horse-shoe prior).
-The target function in Hartman6(300, 6) and DNA aligns well with this assumption.
-As a result, SaasBO exhibits outstanding performance in both tasks.
-However, when the objective deviates from this assumption, as seen in cases like Ackley(150, 150), StybTang(200, 200) and even Rosenbrock(300, 100), the performance of SaasBO experiences a dramatic drop.
-Another example is RDUCB and Tree-UCB, which assume that the target function can be decomposed as a summation of a group of functions, each over a small number of variables.
-RDUCB and Tree-UCB perform well in synthetic benchmarks that match this assumption, such as Rosenbrock(100, 100), Rosenbrock(300, 100), and StybTang(200, 200).
-However, their performance drops significantly in cases that do not align with their assumption, such as Ackley(150, 150) and Ackley(300, 150).
-Across all the benchmarks, the BNN based methods performed worse than most GP based methods.
-This might be due to that under low data regimes, it is more challenging for a neural network to capture the shape of the target function and quantify the uncertainty, especially for high-dimensional targets.
-In contrast, the standard BO exhibits much greater robustness to the structure of the target function.
-Using standard GP equipped with Matérn kernels, both SBO-1 and SBO-2 consistently rank among the best in all tasks, regardless of whether the target function depends only on a subset of variables, can be decomposed, etc.
-Interestingly, as shown in Appendix Fig.7, even using a single length-scale parameter, the standard BO, i.e., SBO-3, still consistently shows superior performance, which closely matches ARD Matérn across most tasks, except for Ackley(300, 150), SBO-3 struggled in finding right locations to query, leading to poor performance.
+On average, SBO-Matérn, SBO-SE(RI) and VBO consistently deliver top-tier results, often outperforming the other methods by a large margin, e.g., on benchmarks Ackley (300, 150), DNA, and SVM.
+It is noteworthy that SBO-SE also performed exceptionally well when the input dimension $d\leq 180$, as seen in benchmarks Rosenbrock (100, 100), Mopta08, NAS, and DNA.
+However, as $d$ increases, e.g., Ackley (300, 150), Hartman6 (300, 6), Rosenbrock (300, 100), StybTang(200, 200) and SVM (388), the performance of SBO-SE dropped dramatically due to the gradient vanishing issue (see our theoretical analysis and numerical experiments in Section 4).
+After applying our robust initialization, the performance of SE kernels, namely, SBO-SE (RI) immediately recovered to top-tier levels, closely matching both SBO-Matérn and VBO.
+On the other hand, though SBO-Matérn used a small length-scale initialization ($\ell_{0}=0.693$), the Matérn kernel’s structure has made it far less susceptible to the gradient vanishing issue, enabling it to handle much higher dimensions than the SE kernel.
+Free from this issue, SBO-Matérn consistently achieved top-tier results.
 Runtime Performance.
 Next, we examined the runtime performance.
-For each method, we report the average maximum function value obtained at each step in Fig.
-5 and Fig.
-6 of the Appendix.
-We supplement the runtime standard deviation and the result of SBO-3 in Fig.
-8 of the Appendix.
-Note that since we have twelve baseline methods, reporting the standard deviations of all the methods makes the figure very cluttered.
-To ensure clarity, in Fig.
-8, we show the standard deviation of the best, worst, and median performed baselines so far at each step.
-One can see that after a few initial steps, SBO-1 and SBO-2 (and SBO-3 in most cases) consistently generated superior queries, finding larger function values.
-Their curves are mostly above those of the competing methods, indicating a more efficient optimization of the target function.
-This trend is particularly evident in examples such as Rosenbrock(100, 100), StybTang(200, 200), Mopta08, Rover, SVM, and more.
-The collective results demonstrate that the standard GP (with Matérn kernels) is effective for high-dimensional optimization.
-In comparison with more complex surrogate models using additional assumptions, the standard GP exhibits greater flexibility in adapting to the target function whose intrinsic structures are often agnostic in practice, and enables superior optimization results.
+Fig.
+5 reports the average maximum function value obtained at each step and the standard deviation across all runs.
+Since we have 16 methods for comparison, showing the standard deviation for every method makes the figure cluttered.
+To ensure clarify, we show the curves and standard deviation of SBO, VBO, and the best, worst, and median performed remaining baselines so far at each step.
+After an initial stage, SBO-Matérn, SBO-SE (RI), and VBO consistently produced superior queries, achieving larger function values.
+Their performance curves are generally above those of competing methods, reflecting a more efficient optimization of the target function.
+This trend is particularly noticeable in benchmarks such as Ackley (300, 150), Rosenbrock (300, 100), and SVM.
+When $d>180$, however, the performance of SBO-SE declines, with its curve positioned toward the bottom due to the gradient vanishing issue.
+In benchmarks such as Rosenbrock (100, 100) and DNA (180), the curve for VBO is predominantly above those of SBO-Matérn and SBO-SE (RI), indicating even more efficient optimization.
+However, on Ackley (150, 150) and Ackley (300, 150), VBO’s curve is mostly below that of SBO-Matérn and SBO-SE (RI).
+For the remaining benchmarks, their curves are close and often overlap.
+Overall, in all cases, they tend to converge to similar function values.
+These results collectively demonstrate that standard BO can excel in high-dimensional optimization problems.
+The lack of additional structural assumptions in GP modeling may make it more flexible in capturing various correlations within high-dimensional spaces.
+The primary challenge for standard BO arises from the gradient vanishing issue, which is often due to improper initialization of the length-scales.
+By using our robust initialization method, this risk can be significantly mitigated.
+Additional Results.
+We conducted extensive additional evaluations and ablation studies to examine the impact of alternative GP training methods, exploration parameters in UCB, and other acquisition functions on SBO performance in high-dimensional settings.
+We found that MLE achieves optimization results comparable to the MAP training employed in our evaluation.
+However, using concentrated priors on the length-scale parameters (concentrated on small mean values) — such as the popular Gamma(3,6) prior that is the default choice in BOTorch — often leads to training failure and poor performance, aligning with our theoretical analysis.
+Both MLE and MAP perform similarly to Markov Chain Monte Carlo (MCMC) training in terms of optimization outcomes, but MCMC is significantly slower.
+UCB is notably robust to variations in the exploration parameter.
+Additionally, log-EI facilitates strong performance in high-dimensional optimization, whereas EI and TS often lead to subpar results.
+Detailed results and discussions are provided in Appendix Section C.
 Figure 3: Optimization performance in synthetic benchmarks. The results were summarized from 10 runs.
 Figure 4: Optimization performance in real-world problems. The results were summarized from 10 runs.
-4.3 Surrogate Learning
-Next, we explicitly evaluated the performance of the standard GP in learning high-dimensional functions to confirm if it can serve as a reliable surrogate model for high-dimensional BO.
-To this end, from the optimization trajectories of all the methods, we randomly sampled 300 training examples from the six synthetic benchmark functions and four real-world benchmark tasks, including Mopta08 (124), Rover (60), DNA (180) and SVM (388).
-We sampled another 100 examples for testing.
-We compared with the surrogate models used in SaasBO, namely, placing a horse-shoe prior over the the length-scale parameters to prune the variables.
-We also compared with a random projection GP, denoted as ProjGP, to align with the baselines employing low-dimensional embeddings.
-ProjGP employs the kernel function in the form of $\vartheta({\bf x},{\bf x}^{\prime})=\kappa({\bf A}{\bf x},{\bf A}{\bf x}^{\prime})$, where ${\bf A}$ is a random projection matrix.
-We varied the rank of ${\bf A}$ from 10 and 20.
-Each element is sampled from the standard normal distribution.
-In addition, we compared with an additive GP, denoted as AddGP, where the kernel takes the form $\vartheta({\bf x},{\bf x}^{\prime})=\sum_{j}\kappa({\bf x}^{j},{\bf x}^{\prime j})$ where each ${\bf x}^{j}\subset{\bf x}$ and ${\bf x}^{\prime j}\subset{\bf x}^{\prime}$, ${\bf x}$ and ${\bf x}^{\prime}$ are the union of ${\bf x}^{j}$’s and ${\bf x}^{\prime j}$’s, respectively.
-AddGP imposes a decomposition structure in the functional space; see (5).
-We tested AddGP with randomly partitioning the input into 10 variables per group.
-We tested two versions: one using non-overlapping partitions, and the other overlapping partitions with one variable shared by successive groups.
-We trained the standard GP with ARD Matérn-5/2 kernel and with a single length-scale Matérn-5/2 kernel using MLE.
-SaasBO was trained by NUTS.
-For a fair comparison, ProjGP and AddGP were based on ARD Matérn-5/2 and trained via MLE as well.
-(a) ARD Matérn
-(b) Matérn
-(c) SaasBO-NUTS
-Figure 3: Alignment between the prediction (y-axis) and ground-truth function values (x-axis). The top row includes the results for Mopta08 (124) and the second row DNA (180).
-We repeated the experiment for five times, and evaluated mean-square-error (MSE), the average test log-likelihood (LL), and the standard deviation.
-Table 4 and Appendix Table 5 report the results for all the GP models.
-The best and second best results are highlighted with a bold font.
-As we can see, in every dataset, standard GP with Mateŕ kernels achieves either the best or the second best result, demonstrating excellent and robust performance in learning high-dimensional functions.
-In addition, we examined the alignment between the prediction and the ground-truth across the test examples from Mopta08 (124) and DNA (180).
-As illustrated in Fig.
-3, the prediction (and the predictive standard deviation) of ARD Matérn matches the ground-truth equally as well as SaasBO-NUTS, with a slight improvement over Matérn (using a single length-scale parameter).
-Together these results demonstrate that directly incorporating high-dimensional inputs into the kernel function does not necessarily compromise the learning performance.
-This implies that, even when the input space is high dimensional, the popular Matérn kernel remains expressive enough to capture the similarities in the space, allowing them to perform well in covariance modeling.
-On the contrary, enforcing (unnecessary) low-rank structures or sparsity might cause information loss, potentially detrimental to optimization performance.
-Our observation is consistent with the recent work in kernel operator learning [3], where the standard kernel (GP) regression has demonstrated the state-of-the-art performance in learning function-to-function mappings.
-In this context, the input comprises of a discretized function, typically of very high-dimensionality (e.g., several thousands).
-4.4 Computational Efficiency
-So far, in our BO loop, the GP model is trained via MAP or MLE.
-Although the performance has been promising, it is worth investigating whether a full Bayesian approach would yield substantial improvements.
-In particular, we can use MCMC to perform posterior inference of the GP parameters, and then integrate them out during the prediction.
-We conducted another round of BO experiments.
-For each step, we used NUTS to train GP with ARD Matérn.
-We adopted the same settings as used in SaasBO-NUTS [6]: The number of warm up steps was set to 512, and after the warm up stage, we ran 256 steps, and selected the samples at every 16-th step as the posterior samples.
-We show the running and final optimization performance in Fig 4 , and Appendix Fig.
-9.
-It can be observed that while NUTS contributes to a slight improvement in several tasks, the enhancement is marginal.
-However, running NUTS is much more computationally costly compared to doing MAP and MLE.
-For example, on DNA, NUTS takes 59x and 40x running time of MAP and MLE, respectively, on average in each step.
-Therefore, for practical purposes, using MAP or MLE for GP training appears sufficient.
-It is computationally efficient, and the sacrifice of optimization performance is minimal compared to much more costly training methods like NUTS.
-By contrast, for more complex surrogate models, such as SaasBO, the MCMC based training might be more crucial for achieving good performance.
-As depicted in Fig.
-2, SaasBO-NUTS demonstrates a substantial improvement over SaasBO-MAP across all real-world benchmarks, excluding NAS201.
-However, the computational cost of running SaasBO-NUTS is considerably higher, ranging from 5.4x (NAS201) to 33.9x (SVM, DNA) compared to the running time of SaasBO-MAP.
-The substantial increase in computational costs may pose a bottleneck for practical usage.
-Detailed per-optimization-step running time for these methods are provided in Appendix Table 6 and 7.
-Dataset ARD Matérn Matérn SaasBO Proj-GP Add-GP
-Mopta08(124) 0.034$\pm$0.006 0.145$\pm$0.057 0.038$\pm$0.008 0.461$\pm$0.157 0.107$\pm$0.033
-Rover(60) 0.327$\pm$0.078 0.284$\pm$0.09 0.284$\pm$0.09 0.548$\pm$0.097 0.321$\pm$0.097
-DNA(180) 0.149$\pm$0.038 0.332$\pm$0.053 0.152$\pm$0.049 0.747$\pm$0.121 0.312$\pm$0.061
-SVM(388) 0.034$\pm$0.01 0.103$\pm$0.017 0.034$\pm$0.014 0.392$\pm$0.071 0.077$\pm$0.012
-Ackley(150, 150) 0.017$\pm$0.008 0.007$\pm$0.001 0.007$\pm$0.001 0.034$\pm$0.007 0.007$\pm$0.001
-Ackley150(300, 150) 0.023$\pm$0.006 0.012$\pm$0.001 0.011$\pm$0.001 0.080$\pm$0.024 0.021$\pm$0.003
-Hartmann6 (300, 6) 0.017$\pm$0.013 0.419$\pm$0.048 0.024$\pm$0.023 0.583$\pm$0.094 0.416$\pm$0.084
-Rosenbrock(100, 100) 0.084$\pm$0.013 0.118$\pm$0.023 0.124$\pm$0.025 0.378$\pm$0.069 0.097$\pm$0.015
-Rosenbrock(300, 100) 0.090$\pm$0.064 0.160$\pm$0.09 0.131$\pm$0.075 0.533$\pm$0.214 0.138$\pm$0.075
-Stybtang(200, 200) 0.043$\pm$0.018 0.036$\pm$0.014 0.039$\pm$0.012 0.253$\pm$0.040 0.038$\pm$0.009
-Table 4: Mean Squared Error(MSE) of each surrogate model. The results were averaged from five runs.
-Figure 4: Optimization performance when using ARD Matérn trained with MAP and MLE, and NUTS on SVM (388) and DNA (180). The results were summarized from 10 runs.
-4.5 Ablation Study and Failure Modes with Other Settings
-Finally, we conducted a careful ablation study on the alternative settings to carry out standard BO.
-We found, however, those settings, such as using SE kernel and EI acquisition, often lead to inferior performance in high-dimensional optimization, which align with the a few failure cases reported in the literature [15, 20].
-The detailed results and analysis are given in Section B of the Appendix.
-We summarize the favorable and unfavorable settings for high-dimensional optimization in Table 1.
 7 Conclusion
-We have presented a comprehensive investigation on standard BO (SBO) in high-dimensional optimization.
-We discovered that simply using Matérn kernel and UCB, SBO has already achieved the state-of-the-art performance.
-There seems no need for imposing additional structure assumptions for function learning and/or placing strong priors over the model parameters.
-In the future, we will continue applying SBO in more tasks, and seek for theoretical justifications and guarantees.
-Appendix A More Experiment Details
-A.1 Definition of Synthetic Functions
+We conducted a thorough investigation of standard BO in high-dimensional optimization, both empirically and theoretically.
+Our analysis identified gradient vanishing as a major failure mode of standard BO, and we proposed a simple yet robust initialization method to address this issue.
+Our empirical evaluation demonstrates that, once this problem is mitigated, standard BO can achieve state-of-the-art performance in high-dimensional optimization.
+Appendix A Proofs
+A.1 Proof of Proposition 4.1
+Proof.
+We intend to solve $\rho^{2}/e^{\rho^{2}}<\xi$, which is equivalent to solving
+$\log r-r<\log\xi,$
+where $r=\rho^{2}$.
+We then leverage the fact that $\log r<\sqrt{r}$ when $r>1$.
+Since $1/e>\xi$, we know the solution of $\rho$ must be larger than 1, implying that the solution for $r$ is larger than $1$.
+To meet the inequality, we can solve
+$\log r-r<\sqrt{r}-r<\log\xi.$
+It then converts to a quadratic form,
+$\sqrt{r}-\sqrt{r}^{2}<\log\xi.$ (10)
+Note that $\rho=\sqrt{r}$. It is straightforward to solve this inequality and we can obtain that the inequality holds when
+$\rho>\tau_{\text{SE}}=\frac{1}{2}+\sqrt{\frac{1}{4}-\log\xi}.$
+A.2 Proof of Lemma 4.2
+Proof.
+First, since each $x_{ik}$ and $x_{jk}$ independently follow $\text{Uniform}(0,1)$, it is straightforward to obtain $\mathbb{E}[(x_{ik}-x_{jk})^{2}]=\frac{1}{6}$.
+Let us define $\gamma=\|{\bf x}_{i}-{\bf x}_{j}\|^{2}$.
+Then we have $\mathbb{E}[\gamma]=\frac{1}{6}d$.
+According to Hoeffding’s inequality, for all $t>0$,
+$p(|\gamma-\frac{1}{6}d|\geq t)\leq 2\exp\left(-\frac{2t^{2}}{d}\right).$ (11)
+Since $\rho=\frac{\sqrt{\gamma}}{\ell_{0}}$, we have $p(\rho>\tau)=p(\gamma>\tau^{2}\ell_{0}^{2})$.
+Since $d>6\ell_{0}^{2}\tau^{2}$, we set $t=\frac{1}{6}d-\ell_{0}^{2}\tau^{2}>0$, and apply (11) to obtain
+$p(|\gamma-\frac{1}{6}d|\geq\frac{1}{6}d-\ell_{0}^{2}\tau^{2}|)\leq 2\exp\left(-\frac{2\left(\frac{1}{6}d-\ell_{0}^{2}\tau^{2}\right)^{2}}{d}\right)=2\exp\left(-\frac{\left(d-6\ell_{0}^{2}\tau^{2}\right)^{2}}{18d}\right),$
+which is equivalent to
+$p(|\gamma-\frac{1}{6}d|<\frac{1}{6}d-\ell_{0}^{2}\tau^{2})>1-2\exp\left(-\frac{\left(d-6\ell_{0}^{2}\tau^{2}\right)^{2}}{18d}\right).$ (12)
+Since $p(\gamma>\ell_{0}^{2}\tau^{2})\geq p(\ell_{0}^{2}\tau^{2}<\gamma<\frac{1}{3}d-\ell_{0}^{2}\tau^{2})=p(|\gamma-\frac{1}{6}d|<\frac{1}{6}d-\ell_{0}^{2}\tau^{2})$, combining with (12) we obtain the bound
+$p(\rho>\tau)=p(\gamma>\ell_{0}^{2}\tau^{2})>1-2\exp\left(-\frac{\left(d-6\ell_{0}^{2}\tau^{2}\right)^{2}}{18d}\right).$
+A.3 Proof of Proposition 4.3
+Proof.
+To solve $\rho^{2}/e^{\sqrt{5}\rho}<\xi$, it is equivalent to solving
+$\frac{1}{5}\frac{r^{2}}{e^{r}}<\xi,\;\;\;r=\sqrt{5}\rho,$
+which is further equivalent to
+$\log\frac{1}{5}+2\log r-r<\log\xi.$
+We leverage $\log r<\sqrt{r}$ for $r>1$, and solve the upper bound,
+$\log\frac{1}{5}+2\sqrt{r}-r<\log\xi,$
+which gives
+[…]
+A.4 Proof of Lemma 5.1
+Proof.
+First, similar to the proof in Section A.2, under the uniform distribution on $[0,1]^{d}$, from Hoeffding’s inequality, we have for all $t>0$,
+$p(|\gamma-\frac{1}{6}d|\geq t)\leq 2\exp\left(-\frac{2t^{2}}{d}\right).$ (14)
+When we set $\ell_{0}=c\sqrt{d}$ such that $c>1/\sqrt{6}\tau$, we have $\ell_{0}^{2}\tau^{2}>d/6$.
+Let us choose $t=\ell_{0}^{2}\tau^{2}-d/6=c^{2}\tau^{2}d-d/6>0$, and apply the Hoeffding’s inequality (14),
+$p\left(|\gamma-d/6|\geq\ell_{0}^{2}\tau^{2}-d/6\right)\leq 2\exp\left(-\frac{2(\ell_{0}^{2}\tau^{2}-d/6)^{2}}{d}\right)=2\exp\left(-2(c^{2}\tau^{2}-\frac{1}{6})^{2}d\right).$ (15)
+Since $p(\rho\geq\tau)=p(\gamma\geq\ell_{0}^{2}\tau^{2})\leq p(|\gamma-d/6|\geq\ell_{0}^{2}\tau^{2}-d/6)$, we obtain
+$p(\rho\geq\tau)\leq 2\exp\left(-2(c^{2}\tau^{2}-\frac{1}{6})^{2}d\right).$
+Appendix B More Experiment Details
+B.1 Definition of Synthetic Functions
 Stybtang. We used the following slightly-modified Stybtang function,
@@
 where each $x_{i}\in[0,1]$, ${\bf A}=[A_{ij}]$ and ${\bf P}=[P_{ij}]$ are two pre-defined matrices as in [29]. The global optimum is at ${\bf x}^{\dagger}=[0.20169,0.150011,0.476874,0.275332,0.6573]$.
-Figure 5: Runtime performance: Maximum Function Value Obtained vs. Number of Steps in synthetic benchmarks. The results were averaged from 10 runs.
-Figure 6: Runtime performance: Maximum Function Value Obtained vs. Number of Steps in real-world benchmarks. The results were averaged from 10 runs.
-Figure 7: Ultimate optimization performance: Maximum Function Value obtained at the end of optimization loop, including SBO-3, namely training with a single length-scale parameter in Matérn-5/2 kernel.
-The results were collect from 10 runs.
 Figure 5: Runtime performance: Maximum Function Value Obtained vs.
-Number of Steps in real-world benchmarks, including SBO-3, namely training with a single length-scale parameter in Matérn-5/2 kernel.
+Number of Steps in all real-world and synthetic benchmarks.
+The results were averaged from 10 runs.
+For cleaner view, we plotted the full results only for SBO and VBO.
+For the remaining methods, we show best, median, and worst result at each step.
+Figure 6: Runtime performance comparison with BNN-based BO methods and SBO with a single length-scale parameter.
+Matérn kernel was used to conduct SBO.
+For clarity, VBO was excluded for comparison.
 The results were averaged from 10 runs.
 The shaded region depicts the standard deviation.
-A.2 Implementation
+B.2 Implementation
   • ALEBO. We used ALEBO implementation shared by the Adaptive Experimentation (AX) Platform (version 0.2.2). The source code is at https://github.com/facebook/Ax/blob/main/ax/models/torch/alebo.py.
@@
 […]
-A.3 Computing Environment
+B.3 Computing Environment
 We conducted the experimental investigations on a large computer cluster equipped with Intel Cascade Lake Platinum 8268 chips.
-Task ARD Matérn Matérn SaasBO Proj-GP Add-GP
-Mopta08(124) 0.570$\pm$0.08 -0.277$\pm$0.101 0.560$\pm$0.064 -0.962$\pm$0.183 -0.089$\pm$0.105
-Rover(60) -0.975$\pm$0.216 -0.647$\pm$0.158 -0.638$\pm$0.14 -0.995$\pm$0.116 -0.703$\pm$0.125
-DNA(180) -0.256$\pm$0.261 -0.666$\pm$0.102 -0.183$\pm$0.158 -1.137$\pm$0.143 -0.679$\pm$0.143
-SVM(388) 0.581$\pm$0.119 -0.179$\pm$0.042 0.650$\pm$0.12 -0.837$\pm$0.095 0.062$\pm$0.129
-Ackley(150, 150) 0.428$\pm$0.222 0.633$\pm$0.067 0.574$\pm$0.068 -0.006$\pm$0.268 0.569$\pm$0.066
-Ackley(300, 150) 0.232$\pm$0.155 0.252$\pm$0.057 0.279$\pm$0.077 -0.372$\pm$0.185 0.220$\pm$0.074
-Hartmann6(300, 6) 1.200$\pm$0.169 -0.625$\pm$0.079 1.528$\pm$0.564 -0.927$\pm$0.084 -0.617$\pm$0.118
-Rosenbrock(100, 100) 0.640$\pm$0.081 0.445$\pm$0.062 0.526$\pm$0.08 -0.466$\pm$0.182 0.521$\pm$0.057
-Rosenbrock(300, 100) 0.518$\pm$0.225 0.121$\pm$0.155 0.357$\pm$0.21 -0.706$\pm$0.209 0.189$\pm$0.2
-Stybtang(200, 200) 0.574$\pm$0.164 0.563$\pm$0.073 0.543$\pm$0.067 -0.458$\pm$0.069 0.582$\pm$0.072
-Table 5: Test log likelihood of each surrogate model. The results were averaged over five runs. Bolded are the best and second best results.
 Mopta08 (124) Rover (60) NAS201 (30) DNA (180) SVM (388)
@@
 Table 5: Average running time (in seconds) per optimization step on synthetic benchmark tasks.
-Figure 9: Optimization performance using MLE training, MAP training with a diffuse prior over the length-scale parameters, and NUTS on the rest of benchmarks. The results were summarized from 10 runs.
-Appendix B Ablation Study and Failure Modes With Other Settings
-In this section, we conducted a careful ablation study to investigate alternative popular settings in executing standard BO.
-From these settings, we identified a few failure modes during high-dimensional optimization.
-We provided analysis and discussions accordingly.
-B.1 GP Training
-First, we examined MLE training with LBFGS.
-We found that, however, LBFGS often incurs numerical issues, and cannot reliably finish model training.
-This is consistent with the report in other works [12].
-The issue might be because that without constraints (or regularization) over the parameters (such as the length-scales and amplitude), the exploration of LBFGS tends to be overly aggressive, which easily cause numerical problems.
-Hence, using stochastic optimization, like ADAM, might be a wise choice for MLE training.
+Method Structural Assumption
+ALEBO, HESBO Low-dim embedding
+SaasBO Partial variable dependency
+RDUCB, Tree UCB Additive function decomposition
+SBO, TURBO, BNN None
+Table 6: BO methods and their structural assumption about the target function.
+Figure 7: Optimization performance using MLE training, MAP training with a uniform prior over the length-scale parameters, and NUTS on SVM (388) and DNA (180).
+The results were summarized from 10 runs.
+Figure 8: Optimization performance using MLE training, MAP training with a uniform prior over the length-scale parameters, and NUTS on the rest of benchmarks. The results were summarized from 10 runs.
+Appendix C Additional Results
+C.1 BNN-based BO and Standard BO with a Single Length-Scale
+Due to the space limit, we did not report the results of the BNN-based BO in the main paper.
+Here, we supplement the comparison with the aforementioned two BNN-based BO approaches, BNN Laplace and BNN HMC, as specified in Section B.2.
+In addition, we also tested SBO with a single length-scale.
+We used the Matérn kernel.
+The results are reported in Fig.
+6.
+Across all the benchmarks, the performance of the BNN based methods is worse than the median of GP based methods.
+In cases, such as SVM, Hartmann6 (300, 6) and Rosenbrock, the performance of the BNN based methods is close to or overlapping with the performance of the worst GP baselines.
+This might be due to that under low data regimes, it is more challenging for a neural network to capture the shape of the target function and quantify the uncertainty, especially for high-dimensional targets.
+Interestingly, even using a single length-scale parameter, the standard BO, namely SBO-Single — as long as it does not encounter gradient vanishing — still consistently shows superior performance, which closely matches ARD Matérn across most tasks, except for Ackley(300, 150), SBO-Single struggled in finding right locations to query, leading to poor performance.
+C.2 Alternative Ways to GP Training
+Figure 9: Final optimization performance using a diffuse prior, Uniform(0.001, 30), vs. a concentrated Gamma prior, Gamma(3,6), over the length scale parameters.
+Figure 10: Runtime optimization performance using a uniform prior, Uniform(0.001, 30), vs. a concentrated Gamma prior, Gamma(3,6), over the length scale parameters.
+C.2.1 Maximum Likelihood Estimation (MLE)
+In the main paper, we report the SBO performance when training GP via MAP estimation.
+We then examined another popular GP training approach, MLE training.
+We used ADAM optimization [17] [^2] We also experimented with using L-BFGS for MLE training.
+However, L-BFGS frequently encounters numerical issues and often fails to reliably complete model training.
+This is consistent with findings reported in [13].
+The issue likely arises because, without constraints or regularization on parameters like the length-scales and amplitude, L-BFGS tends to explore too aggressively, leading to numerical instability.
+to perform MLE, with learning rate $0.1$ and the maximum number of epochs was set to $500$.
+We employed ARD Matérn kernel.
+The ultimate and running optimization performance is shown in Fig.
+9 and 10.
+As we can see, the BO performance of using MAP and MLE training is close, and in many cases significantly overlaps, such as in DNA, SVM, Ackley (150, 150) and StybTang (200, 200).
+C.2.2 MAP Training with Alternative Priors
 Second, we examined MAP training with alternative prior choices over the length-scale parameters.
-In the main paper, we use a diffuse prior, Uniform(0.001, 30).
-In practice, another commonly used choice is a concentrated Gamma prior, e.g., Gamma(3, 6), which is the default choice in the popular BO library BoTorch.
+In the main paper, we use a uniform prior, Uniform(0.001, 30).
+In practice, another commonly used choice is a concentrated Gamma prior, e.g., Gamma(3, 6), which is the default choice in the popular GP/BO libraries GPyTorch and BOTorch.
 To see how it is concentrated, in Fig.
 11 we compare its probability density curve with other priors, including the aforementioned uniform prior and the diffuse Gamma priors over the amplitude and noise variance.
-We found that employing a concentrated prior over the length-scale parameters will severely deteriorate the high-dimensional optimization performance.
-As shown in Fig 12 and 13, after placing the prior Gamma(3,6) over each length-scale parameter, both the ultimate and runtime optimization performance drops substantially across many tasks, such as Mopta08 (124), SVM (388) and Ackley(150, 150), no matter using ADAM or LBFGS to conduct training.
-To figure out the reason, we further investigated the surrogate learning performance.
-Following the same setting as in Fig.
-3, we examined the prediction alignment after applying the prior Gamma(3,6) over the length-scale parameters.
+We observe that the prior distribution Gamma(3, 6) places a large probability mass around its mean 0.5, strongly regularizing the length-scale parameters toward this value.
+According to our theoretical analysis (see Section 4), this introduces a significant risk of gradient vanishing in high-dimensional settings, potentially leading to failure in GP training.
+To verify the risk, we examined the alignment between the prediction and the ground-truth across the test examples from Mopta08 (124) and DNA (180).
 As depicted in Fig.
-11 (third and fourth column), this concentrated prior leads to a severe prediction deviation from the ground-truth, indicating failure in learning the target function.
-This is consistent with the report in the SaasBO paper [6] (see Fig.1 of that paper).
-The quantitative results shown in Table 8 and 9 further confirm the substantially worse prediction accuracy as compared with GP training with a diffuse prior.
-In addition, we notice that the previous work[21] have also tested standard BO using Gamma(3,6) as the prior over the length-scale parameters.
-In their high-dimensional optimization experiments (see Fig.
-5 of that paper), standard BO performs badly.
-We used our setting (SBO-1 and SBO-2) to re-run one experiment that maximizes a high-dimensional polynomial function.
-It turns out that both SBO-1 and SBO-2 can surpass the best method I-BNN reported by that paper.
-Together these results have shown that a concentrated prior can severely limit GP’s capacity for high-dimensional function estimation, leading to inferior performance in high-dimensional optimization.
+12 (third column), the concentrated prior leads to a severe prediction deviation from the ground-truth, indicating failure in learning the target function.
+This observation aligns with findings reported in the SaasBO paper [6] (see Fig.1 of that paper).
+We further investigated the impact of this prior within the optimization loop.
+As shown in Fig.
+9, after placing the prior Gamma(3,6) over each length-scale parameter, the BO performance drops substantially across many tasks, such as Mopta08 (124), SVM (388) and Ackley(150, 150), no matter using ADAM or LBFGS to conduct training.
 Figure 11: Probability density curves. Gamma(2, 0.15) and Gamma(1.1, 0.5) are the priors for the amplitude and noise variance, respectively.
 Figure 12: Alignment between the prediction (y-axis) and ground-truth function values (x-axis).
-Task MLE MAP-Uniform(1E-3, 30) MAP-Gamma(3, 6) MAP-Gamma(3, 6)
-LBFGS ADAM
-Mopta08(124) 0.034$\pm$0.006 0.035$\pm$0.005 0.837$\pm$0.112 0.856$\pm$0.114
-Rover(60) 0.327$\pm$0.078 0.323$\pm$0.075 0.78$\pm$0.154 0.819$\pm$0.115
-DNA(180) 0.149$\pm$0.038 0.166$\pm$0.049 1.014$\pm$0.097 1.026$\pm$0.101
-SVM(388) 0.034$\pm$0.01 0.039$\pm$0.01 1.057$\pm$0.066 1.06$\pm$0.067
-Ackley(150, 150) 0.017$\pm$0.008 0.026$\pm$0.007 0.019$\pm$0.004 0.019$\pm$0.005
-Ackley(300, 150) 0.023$\pm$0.006 0.07$\pm$0.047 0.431$\pm$0.09 0.459$\pm$0.092
-Hartmann6(300, 6) 0.017$\pm$0.013 0.178$\pm$0.325 0.881$\pm$0.045 0.906$\pm$0.034
-Rosenbrock(100, 100) 0.084$\pm$0.013 0.081$\pm$0.012 0.9$\pm$0.156 0.925$\pm$0.172
-Rosenbrock(300, 100) 0.09$\pm$0.064 0.116$\pm$0.078 0.971$\pm$0.222 1.001$\pm$0.217
-Stybtang(200, 200) 0.043$\pm$0.018 0.054$\pm$0.024 0.698$\pm$0.05 0.732$\pm$0.034
-Table 8: Mean Squared Error(MSE) averaged from five runs. The best results are bolded.
-Task MLE MAP-Uniform(1E-3, 30) MAP-Gamma(3, 6) MAP-Gamma(3, 6)
-LBFGS ADAM
-Mopta08(124) 0.57$\pm$0.08 0.542$\pm$0.104 -1.337$\pm$0.058 -1.715$\pm$0.212
-Rover(60) -0.975$\pm$0.216 -1.062$\pm$0.286 -$1$1.322$\pm$0.073
-DNA(180) -0.256$\pm$0.261 -0.388$\pm$0.35 -1.426$\pm$0.054 -2.026$\pm$0.183
-SVM(388) 0.581$\pm$0.119 0.504$\pm$0.128 -1.453$\pm$0.038 -2.095$\pm$0.129
-Ackley(150, 150) 0.428$\pm$0.222 0.14$\pm$0.218 0.451$\pm$0.088 0.445$\pm$0.087
-Ackley(300, 150) 0.232$\pm$0.155 -0.257$\pm$0.528 -1.006$\pm$0.091 -1.037$\pm$0.087
-Hartmann6(300, 6) 1.2$\pm$0.169 0.67$\pm$0.867 -1.344$\pm$0.029 -1.618$\pm$0.128
-Rosenbrock(100, 100) 0.64$\pm$0.081 0.606$\pm$0.1 -1.308$\pm$0.083 -1.598$\pm$0.42
-Rosenbrock(300, 100) 0.518$\pm$0.225 0.347$\pm$0.325 -1.411$\pm$0.134 -1.967$\pm$0.413
-Stybtang(200, 200) 0.574$\pm$0.164 0.421$\pm$0.228 -1.251$\pm$0.017 -1.41$\pm$0.067
-Table 9: Test log likelihood averaged from five runs. The best results are bolded.
-Figure 12: Ultimate optimization performance using a diffuse prior, Uniform(0.001, 30), vs.
-a concentrated Gamma prior, Gamma(3,6), over the length scale parameters.
-SBO-1 and SBO-2 use Uniform(0.001, 30).
-Figure 13: Runtime optimization performance using a diffuse prior, Uniform(0.001, 30), vs.
-a concentrated Gamma prior, Gamma(3,6), over the length scale parameters.
-SBO-1 and SBO-2 use Uniform(0.001, 30).
-B.2 Alternative Acquisition Functions
-Next, we examined two other popular acquisition functions used in BO: Expected Improvement (EI) and Thompson sampling (TS).
+C.2.3 Fully Bayesian Training
+Third, we investigated whether a fully Bayesian training approach would yield substantial improvements in BO.
+In particular, we can use MCMC to perform posterior inference of the GP parameters, and then integrate them out during the prediction.
+We conducted another round of BO experiments.
+For each step, we used NUTS to train GP with ARD Matérn.
+We adopted the same settings as used in SaasBO-NUTS [6]: where the mean parameter is assigned the standard Gaussian prior $\mathcal{N}(0,1)$, amplitude $\text{Gamma}(2,0.15)$, and noise variance $\text{Gamma}(0.9,10)$.
+We then assigned a uniform prior over each length-scale, $\text{Uniform}(0.001,30)$.
+The number of warm up steps was set to 512, and after the warm up stage, we ran 256 steps, and selected the samples at every 16-th step as the posterior samples.
+We show the running and final optimization performance in Fig 7 , and Fig.
+8.
+We also show the average running time per optimization step in Table 4 and Table 5.
+It can be observed that while NUTS contributes to a slight improvement in several tasks, the enhancement is marginal.
+However, running NUTS is much more computationally costly compared to doing MAP and MLE.
+For example, on DNA, NUTS takes 59x and 40x running time of MAP and MLE, respectively, on average in each step.
+Therefore, for practical purposes, using MAP or MLE for GP training appears sufficient.
+It is computationally efficient, and the sacrifice of optimization performance is minimal compared to much more costly training methods like NUTS.
+C.3 Different Choice of $\lambda$ in UCB
+Figure 13: Runtime optimization performance with different exploration level $\lambda$ in UCB.
+We examined how the exploration parameter $\lambda$ in UCB influences the BO performance in high-dimensional optimization.
+To this end, we used the Maérn kernel and varied $\lambda$ from {0.5, 1.0, 1.5, 2.0}.
+We conducted MLE training via ADAM optimization.
+We show the ultimate and runtime optimization performance in Fig.
+13, respectively.
+It can be seen that, in most benchmarks (Mopta08, Rover, NAS201, DNA, etc.), the performance with different $\lambda$’s is close.
+However, on DNA, SVM and Ackley, the performance of $\lambda=0.5$ is typically worse than the bigger choice of $\lambda$.
+This might be due to that those problems are more challenging (e.g., Ackley has many local minima), a larger exploration level ($\lambda$) is needed to identify better optimization outcomes.
+C.4 Alternative Acquisition Functions
+Figure 14: Final performance with different acquisition functions.
+Figure 15: Runtime performance with different acquisition functions.
+Finally, we evaluated three other acquisition functions for high-dimensional optimization: Expected Improvement (EI), log-EI and Thompson sampling (TS).
 At each step, EI computes the expected improvement upon the best sample so far,
@@
 We used BOTorch to maximize EI through L-BFGS.
+The most recent work [1] introduced log-EI to overcome the numerical challenges in optimizing the original EI acquisition function.
+Accordingly, we also tested log-EI using the BoTorch implementation.
 To conduct TS at each step, we used Sobol Sequence [28] to sample 3K inputs in the domain, and jointly sampled the 3K function values from the GP posterior.
@@
 We used BOTorch implementation of both methods.
-We used ARD Matérn kernel and MLE training at each step.
+We used the Matérn kernel and MLE training at each step.
 The ultimate and runtime optimization performance for using EI, TS-3K and TS-10K, are shown in Fig.
@@
 On Ackley(150, 150) and Ackley(300, 150), the performance of EI and TS are even among the worst.
-The limitation of EI might arise from the gradient vanishing during the optimization, as pointed by the recent work [1].
-A remedy is to use log-EI as the acquisition function.
+The failure of EI might arise from the gradient vanishing during the optimization, as pointed by the recent work [1].
 The failure of the TS might be due to the high-dimension of the input space.
 A few thousand candidates might not be sufficient to cover the whole space; as a result, the selected queries are inferior in indicating the best direction to optimize the target function.
-Figure 14: Ultimate performance with different acquisition functions. Note that SBO-1 and SBO-2 use UCB.
-Figure 15: Runtime performance with different acquisition functions. Note that SBO-1 and SBO-2 use UCB.
-B.3 Using Square Exponential (SE) kernels
-Third, we considered another widely used kernel in BO, namely, the SE kernel.
-We tested with ARD SE and SE kernel using a single length-scale parameter.
-We used MLE training and UCB $(\lambda=1.5)$ to run BO.
-As depicted in Fig.
-17 and 18, with smaller optimization dimensions, such as Mopta08 (124), NAS201 (3) and DNA (180), the performance of SE is among the best.
-However, with higher dimensions, e.g., above 200, the performance of SE drops substantially, such as in SVM (388) and Ackley(300, 150).
-We argue that the inferior performance might be due to that the vanilla SE kernel is not powerful enough to enable GP to well capture high-dimensional functions.
-Consider a single length-scale parameter.
-we observe that the value of the SE kernel is determined by the squared distance between the inputs,
-$\kappa_{\text{SE}}({\bf x},{\bf x}^{\prime})=a\exp\left(-\eta\|{\bf x}-{\bf x}^{\prime}\|^{2}\right).$ (7)
-With the growth of the input dimension $d$, the kernel function value decays much faster than the increase of the distance: a squared decay coupling with the exponential decay.
-The learning can be easily driven to adapt to a sparse, diagonal-dominant covariance matrix, which roughly treats each data point as independent.
-This can make the learning fail to capture the correlation between function values, and result in inferior learning performance.
-By contrast, the value of the Matérn kernel (see (1)) is dominated by the distance rather than the squared distance.
-Therefore, with the growth of the distance, the kernel function value decays much slower than the SE kernel.
-It is easier for the optimization to adjust the length-scale parameter so as to avoid “diagonal-like” covariance matrices.
-Consequently, the correlation between function values can be better captured in the high dimensional input space.
-To verify our argument, we examined the GP learning performance with ARD SE kernel and ARD Matérn kernel.
-We used Rosenbrock(d,d) and Hartmann6(d, 6) as test cases.
-The former is a “full-rank” function, namely, all the $d$ input variables are relevant to the function value.
-The latter has a “low-rank” structure in that only the first 6 variables out of the $d$ input variables determines the function value.
-For each test case, we increased the input dimension from 50 to 300.
-We randomly generated 500 training examples and 100 test examples.
-We then trained each method with MLE, and examined the mean square error (MSE) and the test log likelihood.
-We repeated the training and test procedure for five times, and report the average MSE, average test log likelihood, and the standard deviations in Fig.
-16.
-As we can see, in both test cases, when the input dimension $d$ does not exceed 150, the performance of using SE and Matérn-5/2 is close, in terms of both MSE and log likelihood.
-The two curves almost overlap.
-However, for $d>150$ in both test scenarios, the MSE significantly escalates when employing the SE kernel, accompanied by a substantial decrease in the test log likelihood.
-These observations suggest a failure of the SE kernel in effectively learning the target function.
-In comparison, employing the Matérn kernel maintains a consistently favorable prediction performance even when $d>150$.
-On the Rosenbrock dataset, the MSE gradually increases while the test log likelihood decreases, following a roughly linear trend.
-This outcome is expected as, with the rise in input dimension $d$, more training examples are necessary to sustain comparable performance.
-Furthermore, on the Hartmann6 dataset, little variation is observed in both MSE and test log likelihood with increasing the input dimension when using the Matérn kernel.
-This suggests that the GP has effectively discerned the pertinent input variables and mitigated the influence of irrelevant ones.
-In summary, these results collectively underscore the advantage of employing Matérn kernels for learning high-dimensional functions.
-It is worth noting that the most recent work [12] designs a log-normal prior over each length-scale parameter to overcome the learning issue with the ARD SE kernel.
-Figure 16: Learning high-dimension functions with different kernels. We trained on 500 points, and tested on 100 points. Results shown with mean and standard deviation over 5 runs.
-Figure 17: Ultimate optimization performance with different kernels. Note that SBO-1 and SBO-2 use Matérn-5/2 kernel.
-Figure 18: Runtime optimization performance with different kernels. Note that SBO-1 and SBO-2 use Matérn-5/2 kernel.
-B.4 Different Choice of $\lambda$ in UCB
-Finally, we examined how the exploration parameter $\lambda$ in UCB (see (4)) influences the performance in high-dimensional optimization.
-To this end, we used ARD Maérn kernel and varied $\lambda$ from {0.5, 1.0, 1.5, 2.0}.
-We conducted MLE training via ADAM optimization.
-We show the ultimate and runtime optimization performance in Fig.
-19 and 20, respectively.
-It can be seen that, in most benchmarks (Mopta08, Rover, NAS201, DNA, etc.), the performance with different $\lambda$’s is close.
-However, on DNA, SVM and Ackley, the performance of $\lambda=0.5$ is typically worse than the bigger choice of $\lambda$.
-This might be due to that those problems are more challenging (e.g., Ackley has many local minima), a larger exploration level ($\lambda$) is needed to identify better optimization outcomes.
-Figure 19: Ultimate optimization performance with different exploration level $\lambda$ in UCB. Note that SBO-1 and SBO-2 use $\lambda=1.5$.
-Figure 20: Runtime optimization performance with different exploration level $\lambda$ in UCB. Note that SBO-1 and SBO-2 use $\lambda=1.5$.
+The use of log-EI performed excellently, achieving results comparable to those obtained with UCB.
+Appendix D Limitation
+Our current work has identified gradient vanishing as a critical failure mode of SBO in high-dimensional settings.
+We have proposed an effective, robust initialization method to mitigate this issue.
+However, we have not yet analyzed whether there is an upper limit to the input dimension for which SBO can function properly, or what that limit might be.
+To answer these questions, we will continue our investigation in future work, both theoretically and empirically.

### v5

@@
-Standard Gaussian Process Can Be Excellent for High-Dimensional Bayesian Optimization
+Standard Gaussian Process is All You Need for High-Dimensional Bayesian Optimization
 Abstract
@@
 To address this gap, we present a systematic investigation.
-First, through a comprehensive evaluation across eleven widely used benchmarks, we found that while the popular Square Exponential (SE) kernel often leads to poor performance, using Matérn kernels enables standard BO to consistently achieve top-tier results, frequently surpassing methods specifically designed for high-dimensional optimization.
+First, through a comprehensive evaluation across twelve benchmarks, we found that while the popular Square Exponential (SE) kernel often leads to poor performance, using Matérn kernels enables standard BO to consistently achieve top-tier results, frequently surpassing methods specifically designed for high-dimensional optimization.
 Second, our theoretical analysis reveals that the SE kernel’s failure primarily stems from improper initialization of the length-scale parameters, which are commonly used in practice but can cause gradient vanishing in training.
 We provide a probabilistic bound to characterize this issue, showing that Matérn kernels are less susceptible and can robustly handle much higher dimensions.
-Third, we propose a simple robust initialization strategy that dramatically improves the performance of the SE kernel, bringing it close to state-of-the-art methods, without requiring any additional priors or regularization.
+Third, we propose a simple robust initialization strategy that dramatically improves the performance of the SE kernel, bringing it close to state-of-the-art methods, without requiring additional priors or regularization.
 We prove another probabilistic bound that demonstrates how the gradient vanishing issue can be effectively mitigated with our method.
 Our findings advocate for a re-evaluation of standard BO’s potential in high-dimensional settings.
+The code is released at https://github.com/XZT008/Standard-GP-is-all-you-need-for-HDBO.
 1 Introduction
 Many applications require optimizing complex functions, allowing queries only for the function values, possibly with noises, yet without any gradient information.
-Bayesian Optimization (BO) [27, 21] is a popular approach for addressing such challenges.
-BO typically employs Gaussian process (GP) [32] as a probabilistic surrogate.
+Bayesian Optimization (BO) (Snoek et al., 2012; Mockus, 2012) is a popular approach for addressing such challenges.
+BO typically employs Gaussian process (GP) (Williams & Rasmussen, 2006) as a probabilistic surrogate.
 It iteratively approximates the target function, integrates the posterior information to maximize an acquisition function for generating new inputs at which to query, then updates the GP model with new examples, and concurrently approach the optimum.
-Despite numerous successes, there has been a widespread belief that BO with standard GP regression, referred to as standard BO, is limited to low-dimensional optimization problems [8, 23, 6, 22, 18, 31, 19].
-It is commonly thought that the number of optimization variables should not exceed 15 or 20, as beyond this threshold, BO is prone to failure [8, 23].
+Despite numerous successes, there has been a widespread belief that BO with standard GP regression, referred to as standard BO, is limited to low-dimensional optimization problems (Frazier, 2018; Nayebi et al., 2019; Eriksson & Jankowiak, 2021; Moriconi et al., 2020; Letham et al., 2020; Wang et al., 2016; Li et al., 2016).
+It is commonly thought that the number of optimization variables should not exceed 15 or 20, as beyond this threshold, BO is prone to failure (Frazier, 2018; Nayebi et al., 2019).
 A continuous line of research is dedicated to developing novel high-dimensional BO methods.
 The fundamental strategy is to impose strong structural assumptions into surrogate modeling so as to avoid directly dealing with high-dimensional inputs.
-One class of methods assumes a decomposition structure within the functional space [16, 24, 10, 33], where the target function is expressed as the summation of a group of low-dimensional functions, each operating over a small number of variables.
+One class of methods assumes a decomposition structure within the functional space (Kandasamy et al., 2015; Rolland et al., 2018; Han et al., 2021; Ziomek & Ammar, 2023), where the target function is expressed as the summation of a group of low-dimensional functions, each operating over a small number of variables.
 Another family of methods assumes that the inputs are intrinsically low-rank.
-These methods either project the original inputs into a low-dimensional space [31, 23, 18] or use sparse-inducing priors to trim down massive input variables [6].
+These methods either project the original inputs into a low-dimensional space (Wang et al., 2016; Nayebi et al., 2019; Letham et al., 2020) or use sparse-inducing priors to trim down massive input variables (Eriksson & Jankowiak, 2021).
 Subsequently, GP surrogates are built with the reduced input dimensions.
@@
   • Empirical Results.
-We investigated BO with standard GP across eleven widely-used benchmarks, including six synthetic and five real-world high-dimensional optimization tasks.
-The number of variables ranged from 30 to 388.
-We compared standard BO with nine state-of-the-art high-dimensional BO methods, and performed extensive evaluations and ablation studies.
-Surprisingly, while the popular SE kernel often led to poor performance, switching to the ARD Matérn kernel enabled standard BO to consistently achieve the best (or near-best) optimization performance across all the tasks.
+We investigated BO with standard GP across eleven widely used benchmarks and one novel benchmark, encompassing six synthetic and six real-world high-dimensional optimization tasks.
+The number of variables ranged from 30 to 1,003.
+We compared standard BO with nine state-of-the-art high-dimensional BO methods, and performed extensive evaluations.
+Surprisingly, while the popular SE kernel often led to poor performance, switching to the ARD Matérn kernel enabled standard BO to nearly always achieve the best (or near-best) optimization performance.
 When successful, standard BO seems more flexible in accommodating various structures within the target functions.
@@
 2 Standard Bayesian Optimization
-Consider maximizing a $d$-dimensional objective function $f:\mathcal{X}\subset\mathbb{R}^{d}\rightarrow\mathbb{R}$.
-The form of $f(\cdot)$ is unknown.
+Consider maximizing a $d$-dimensional objective function $f:\mathcal{X}\subset\mathbb{R}^{d}\rightarrow\mathbb{R}$, where the function form is unknown.
 We can query the function value at any input ${\bf x}\in\mathcal{X}$, possibly with noises, but no gradient information is accessible.
@@
 This information is then integrated to compute an acquisition function, which measures the utility of querying at any new input location given the current function estimate.
-The utility typically encapsulates an exploration-exploitation trade-off.
 By maximizing the acquisition function, BO identifies a new input location at which to query, ideally closer to the optimum.
@@
 BO begins by querying at a few (typically randomly selected) input locations, and trains an initial surrogate model.
-Then the iterative procedure repeats until convergence or a predefined stopping criterion is met.
-The standard BO adopts GP regression [32] for surrogate modeling.
+The iterative procedure repeats until convergence or a stopping criterion is met.
+The standard BO adopts GP regression (Williams & Rasmussen, 2006) for surrogate modeling.
 A GP prior is placed over the target function, $f\sim\mathcal{GP}\left(m({\bf x}),\kappa({\bf x},{\bf x}^{\prime})\right)$, where $m({\bf x})$ is the mean function, which is often set as a constant, and $\kappa(\cdot,\cdot)$ is the covariance function, which is usually chosen as a Mercer kernel function.
@@
 Given training inputs ${\bf X}=[{\bf x}_{1},\ldots,{\bf x}_{N}]^{\top}$ and (noisy) outputs ${\bf y}=[y_{1},\ldots,y_{N}]^{\top}$, let us denote the function values at ${\bf X}$ as ${\bf f}=[f({\bf x}_{1}),\ldots,f({\bf x}_{N})]^{\top}$, which according to the GP prior, follow a multi-variate Gaussian distribution, $p({\bf f})=\mathcal{N}({\bf f}|{\bf m},{\bf K})$, where ${\bf m}=[m({\bf x}_{1}),\ldots,m({\bf x}_{N})]^{\top}$, ${\bf K}$ is the kernel matrix on ${\bf X}$, and each $[{\bf K}]_{ij}=\kappa({\bf x}_{i},{\bf x}_{j})$.
-Then one can employ a Gaussian likelihood for the observations ${\bf y}$, and obtain
+Then one can employ a Gaussian likelihood for the observations ${\bf y}$, and
 $p({\bf y}|{\bf X})=\mathcal{N}({\bf y}|{\bf m},{\bf K}+\sigma^{2}{\bf I}),$ (3)
 where $\sigma^{2}$ is the noise variance.
-To estimate the GP parameters, e.g., the length-scales $\bm{\ell}$, one can maximize the marginal likelihood or use MCMC sampling.
-The predictive distribution (given the GP parameters) at any new input ${\bf x}^{*}$, is a closed-form, conditional Gaussian distribution, $p\left(f({\bf x}^{*})|{\bf y}\right)=\mathcal{N}\left(f({\bf x}^{*})|\mu({\bf x}^{*}),v({\bf x}^{*})\right)$, where $\mu({\bf x}^{*})=m({\bf x}^{*})+\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}({\bf y}-{\bf m})$, and $v({\bf x}^{*})=\kappa({\bf x}^{*},{\bf x}^{*})-\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}\kappa({\bf X},{\bf x}^{*})$.
+To estimate the GP parameters, e.g., the length-scales $\bm{\ell}$, one can maximize the marginal likelihood (3).
+The predictive distribution at any new input ${\bf x}^{*}$, is conditional Gaussian, $p\left(f({\bf x}^{*})|{\bf y}\right)=\mathcal{N}\left(f({\bf x}^{*})|\mu({\bf x}^{*}),v({\bf x}^{*})\right)$, where $\mu({\bf x}^{*})=m({\bf x}^{*})+\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}({\bf y}-{\bf m})$, and $v({\bf x}^{*})=\kappa({\bf x}^{*},{\bf x}^{*})-\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}\kappa({\bf X},{\bf x}^{*})$.
 In each iteration, given the current GP surrogate model, an acquisition function is maximized to identify the next input location at which to query.
 One commonly used acquisition function is the upper confidence bound (UCB), defined as $\text{UCB}({\bf x})=\mu({\bf x})+\lambda\sqrt{v({\bf x})},\;\;{\bf x}\in\mathcal{X}$, where $\lambda$ represents the exploration level.
-There have been other popular acquisition functions, such as Expected Improvement (EI) [15], Thompson sampling (TS) [25], the recently proposed log-EI [1], among others.
+There have been other popular acquisition functions, such as Expected Improvement (EI) (Jones et al., 1998), Thompson sampling (TS) (Russo et al., 2018), the recently proposed log-EI (Ament et al., 2024), among others.
 3 High Dimensional Bayesian Optimization
@@
 There can be a variety of choices for the group number $M$ and each variable group ${\bf x}^{j}$.
-In [16], the input ${\bf x}$ is partitioned into non-overlapped groups.
+In (Kandasamy et al., 2015), the input ${\bf x}$ is partitioned into non-overlapped groups.
 After every a few BO iterations, a new partition is selected from a set of random partitions.
 The selection is based on the model evidence.
-In [24], the variable groups can overlap, and are represented as the maximum cliques on a dependency graph.
+In (Rolland et al., 2018), the variable groups can overlap, and are represented as the maximum cliques on a dependency graph.
 A Gibbs sampling method was developed to learn the structure of the dependency graph.
-Han et al., [10] proposed Tree-UCB, which restricted the dependency graph to be tree-structured so as to boost the efficiency of structure learning.
-However, the more recent work [33] points out that, learning the group structure through relatively small data can be misleading.
+Han et al.
+(2021) proposed Tree-UCB, which restricts the dependency graph to be tree-structured so as to boost the efficiency of structure learning.
+However, the more recent work (Ziomek & Ammar, 2023) points out that, learning the group structure through relatively small data can be misleading.
 A wrongly learned structure can cause queries to be stuck at local optima.
-The work proposes RDUCB that randomly decomposes ${\bf x}$ into a tree structure, and surprisingly works better than those learned structures through various methods.
+This work proposes RDUCB that randomly decomposes ${\bf x}$ into a tree structure, and surprisingly works better than those learned structures through various methods.
 Structural Assumption in Input Space.
@@
 Many of them introduce Low-dimensional Embeddings.
-Wang et al., [31] proposed a method named REMBO, building the GP surrogate in a low-dimensional embedding space, $\mathcal{Z}\subset\mathbb{R}^{d_{\text{emb}}}$ where $d_{\text{emb}}\ll d$.
+Wang et al.
+(2016) proposed a method named REMBO, building the GP surrogate in a low-dimensional embedding space, $\mathcal{Z}\subset\mathbb{R}^{d_{\text{emb}}}$ where $d_{\text{emb}}\ll d$.
 The acquisition function is optimized in the embedding space.
@@
 If the recovered ${\bf x}$ is out of the domain $\mathcal{X}$, it will be clipped as a boundary point.
-To avoid clipping, Nayebi et al., [23] proposed HESBO, which randomly duplicates ${\bf z}$ (and/or flip the signs) to form ${\bf x}$.
+To avoid clipping, Nayebi et al.
+(2019) proposed HESBO, which randomly duplicates ${\bf z}$ (and/or flip the signs) to form ${\bf x}$.
 This way, the optimization of ${\bf z}$ can respect the same bounding constraints as in $\mathcal{X}$.
 This essentially imposes further constraints on the embedding space.
-A more flexible solution, ALEBO, was proposed in  [18].
+A more flexible solution, ALEBO, was proposed in  (Letham et al., 2020).
 ALEBO also uses a random matrix ${\bf A}$ to recover the input ${\bf x}$.
@@
 Another direction is to impose Partial Variable Dependency by triming down the input variables.
-The recent start-of-the-art, SaasBO [6], uses a horse-shoe prior [4] over the length-scale parameter for each variable.
+The recent start-of-the-art, SaasBO (Eriksson & Jankowiak, 2021), uses a horse-shoe prior (Carvalho et al., 2009) over the length-scale parameter for each variable.
 As the horse-shoe prior induces strong sparsity, a massive number of variables can be pruned, substantially reducing the input dimension.
-For effective training with the horse-shoe prior, SaasBO used the No-U-Turn Sampler (NUTS) [12].
-While one can also use MAP estimation, the performance is often much worse.
 4 Theoretical Analysis
@@
 Specifically, we consider at the beginning, all the length-scale parameters are set to the same initial value, $\ell_{1}=\ldots=\ell_{d}=\ell_{0}$.
-From the marginal likelihood (3), we can derive the gradient w.r.t each length-scale,
+From the marginal likelihood (3), we derive the gradient w.r.t each length-scale,
 $\frac{\partial\log p({\bf y}|{\bf X})}{\partial\ell_{k}}=\frac{1}{2}{\rm tr}\left({\bf A}\cdot\frac{\partial{\bf K}}{\partial\ell_{k}}\right),$ (5)
@@
 When using the SE kernel as specified in (1), each entry of $\frac{\partial{\bf K}}{\partial\ell_{k}}$ is as follows: $\left[\frac{\partial{\bf K}}{\partial\ell_{k}}\right]_{ii}=0$ and for $i\neq j$,
-$\left[\frac{\partial{\bf K}}{\partial\ell_{k}}\right]_{ij}=\frac{\partial\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})}{\partial\ell_{k}}=\frac{2a}{\ell_{k}}\exp(-\frac{\|{\bf x}_{i}-{\bf x}_{j}\|^{2}}{\ell_{0}^{2}})\frac{(x_{ik}-x_{jk})^{2}}{\ell_{k}^{2}}\leq\frac{2a}{\ell_{k}}\cdot\frac{\rho^{2}}{e^{\rho^{2}}},$ (6)
+$\left[\frac{\partial{\bf K}}{\partial\ell_{k}}\right]_{ij}=\frac{\partial\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})}{\partial\ell_{k}}=\frac{2a}{\ell_{k}}\exp(-\frac{\|{\bf x}_{i}-{\bf x}_{j}\|^{2}}{\ell_{k}^{2}})\frac{(x_{ik}-x_{jk})^{2}}{\ell_{k}^{2}}\leq\frac{2a}{\ell_{k}}\cdot\frac{\rho^{2}}{e^{\rho^{2}}},$ (6)
 where $x_{ik}$ and $x_{jk}$ are the $k$-th element in ${\bf x}_{i}$ and ${\bf x}_{j}$, respectively, and $\rho=\frac{\|{\bf x}_{i}-{\bf x}_{j}\|}{\ell_{0}}$.
 Intuitively, as the input dimension $d$ increases, the squared distance between input vectors grows rapidly.
-Consequently, the factor $\frac{\rho^{2}}{e^{\rho^{2}}}$ in (6) can quickly fall below the machine epsilon $\xi$, causing the gradient vanishing issue [3, 11].
+Consequently, the factor $\frac{\rho^{2}}{e^{\rho^{2}}}$ in (6) can quickly fall below the machine epsilon $\xi$, causing the gradient vanishing issue (Bengio et al., 1994; Hanin, 2018).
 That is, the length-scale cannot be effectively updated according to $\partial\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})/\partial\ell_{k}$, due to the limitations of the numerical precision.
@@
 The inferior prediction accuracy of the GP model further contributes to poor performance of the BO procedure.
-Since $\tau_{\text{SE}}$ is small, gradient vanishing can occur even when $d$ is not large.
-For example, for $\ell_{0}=0.5$ and $\ell_{0}=1$, we have $p(\rho>\tau_{\text{SE}})>0.99$ when $d\geq 205$ and $d\geq 473$, respectively.
+Since $\tau_{\text{SE}}$ is small, gradient vanishing can occur before $d$ becomes extremely large.
+For example, for $\ell_{0}=0.5$ and $\ell_{0}=1$, the probability $p(\rho>\tau_{\text{SE}})$ exceeds $0.99$ when $d\geq 205$ and $d\geq 473$, respectively.
 Why Matérn kernels perform much better?
@@
 We trained GP with both SE and Matérn kernels, using length-scale initialization $\ell_{0}$ from $\{0.1,0.5,0.693,1.0,\sqrt{d}\}$.
-Note $\ell_{0}=0.693=\text{SoftPlus}(0)$ is a popular initialization choice, used as the default choice in influential GP/BO libraries GPyTorch [9] and BOTorch [2].
+Note $\ell_{0}=0.693=\text{SoftPlus}(0)$ is a popular initialization choice, used as the default choice in influential GP/BO libraries GPyTorch (Gardner et al., 2018) and BOTorch (Balandat et al., 2020).
 Training and testing were repeated 20 times for each $d$ and $\ell_{0}$.
-We then evaluated (1) the average relative $L_{2}$ difference between the length-scale vector before and after training, namely, $\|\bm{\ell}_{\text{trained}}-\bm{\ell}_{\text{init}}\|/\|\bm{\ell}_{\text{init}}\|$, (2) the average gradient norm at the first training step, and (3) the average test Mean-Squared-Error (MSE).
+We then evaluated (A) the average relative $L_{2}$ difference between the length-scale vector before and after training, namely, $\|\bm{\ell}_{\text{trained}}-\bm{\ell}_{\text{init}}\|/\|\bm{\ell}_{\text{init}}\|$, (B) the average gradient norm at the first training step, and (C) the average test Mean-Squared-Error (MSE).
 As shown in Fig.
@@
 Alternative method.
-Recent work [13] proposes an intuitive approach to address the learning challenge of the SE kernel in the high-dimensional spaces by constructing a log-normal prior over each length scale, $p(\ell_{k})=\text{LogNormal}(\ell_{k}|\mu_{0}+\frac{\log(d)}{2},\sigma_{0})$.
+Recent work (Hvarfner et al., 2024) proposes an intuitive approach to address the learning challenge of the SE kernel in the high-dimensional spaces by constructing a log-normal prior over each length scale, $p(\ell_{k})=\text{LogNormal}(\ell_{k}|\mu_{0}+\frac{\log(d)}{2},\sigma_{0})$.
 This prior regularizes the length-scale to be on the order of $\sqrt{d}$ during training.
@@
 However, our method, directly motivated by our analysis, is even simpler — requiring no additional prior construction or associated parameters, and has also shown effective in addressing the gradient vanishing issue.
-The comparison with  [13] is given in Section 6.
+The comparison with  (Hvarfner et al., 2024) is given in Section 6.
 6 Comprehensive Evaluation
 6.1 Experimental Settings
-For a comprehensive evaluatio of the standard BO, we employed eleven benchmarks, of which six are synthetic and five are real-world problems. For clarity, all the tasks aim for function maximization.
-$\text{Fun}(d,d^{\prime})$ Structure
-Ackley(300, 150) Partial variable dependency
-Rosenbrok(300, 100) Partial variable dependency
-Nonoverlap additive decomposition
-Hartmann6(300,6) Partial variable dependency
-Rosenbrock(100, 100) Nonoverlap additive decomposition
-Stybtang(200, 200)
-Ackley(150, 150) None
-Table 3: Synthetic benchmarks and the structures within the target function: $d$ is the input dimension, and $d^{\prime}$ is the number of effective variables to compute the target function.
+For a comprehensive evaluation of the standard BO, we employed twelve benchmarks, of which six are synthetic and six are real-world problems. For clarity, all the tasks aim for function maximization.
 Synthetic Benchmarks.
@@
 The ground-truth of the target function values is computed with the first $d^{\prime}$ variables.
-The benchmarks and the structures within the target functions are summarized in Table 3.
+The benchmarks and the structures within the target functions are summarized in Appendix Table 3.
 Real-World Benchmarks.
-We employed the following real-world benchmark problems: Mopta08 (124)  [14], a vehicle design problem.
+We employed the following real-world benchmark problems: Mopta08 (124)  (Jones, 2008), a vehicle design problem.
 The objective is to minimize the mass of a vehicle with respect to 124 design variables, subject to 68 performance constraints.
-We followed [6] to encode the constraints as a soft penalty, which is added into the objective.
-SVM (388) [6], a hyper-parameter tuning problem that optimizes 3 regularization parameters and 385 length-scale parameters for a kernel support vector machine regression model.
-Rover (60), a rover trajectory planning problem from [30].
+We followed (Eriksson & Jankowiak, 2021) to encode the constraints as a soft penalty, which is added into the objective.
+SVM (388) (Eriksson & Jankowiak, 2021), a hyper-parameter tuning problem that optimizes 3 regularization parameters and 385 length-scale parameters for a kernel support vector machine regression model.
+Rover (60), a rover trajectory planning problem from (Wang et al., 2018).
 The goal is to find an optimal trajectory determined by the location of 30 waypoints in a 2D environment.
-DNA (180) [26], a hyper-parameter optimization problem that optimizes 180 regularization parameters for weighted lasso regression on an DNA dataset [20].
-Prior analysis [26] shows that only around 43 variables are relevant to the target function.
-NAS201 (30)  [5], a neural architecture search problem on CIFAR-100 dataset.
-Methods. We tested with three versions of standard BO:
+DNA (180) (vSehić et al., 2022), a hyper-parameter optimization problem that optimizes 180 regularization parameters for weighted lasso regression on an DNA dataset (Mills, 2020).
+Prior analysis (vSehić et al., 2022) shows that only around 43 variables are relevant to the target function.
+NAS201 (30)  (Dong & Yang, 2020), a neural architecture search problem on CIFAR-100 dataset.
+Humanoid Standup (1003): A novel trajectory optimization benchmark based on a humanoid simulator that uses the MuJoCo physics engine (Todorov et al., 2012).
+The problem dimension is 1,003.
+The details are provided in Appendix Section B.2.
+Methods. We tested with four versions of standard BO:
   • SBO-Matérn: GP regression with ARD Matérn-5/2 kernel as specified in (2), the length-scale initialization $\ell_{0}=\text{SoftPlus}(0)=0.693$, as the default choice in the influential GP/BO libraries GPyTorch and BOTorch.
   • SBO-SE: GP regression with ARD SE kernel as specified in (1), the length-scale initialization $\ell_{0}=\text{SoftPlus}(0)=0.693$.
-  • SBO-SE(RI): GP regression with ARD SE kernel, using our proposed robust initialization $\ell_{0}=c\sqrt{d}$ where $d$ is the input dimension and $c=1$
+  • SBO-Matérn (RI): GP regression with ARD Matérn-5/2 kernel, using our proposed robust initialization $\ell_{0}=c\sqrt{d}$ where $d$ is the input dimension and $c=1$.
+  • SBO-SE (RI): GP regression with ARD SE kernel, using our robust initialization with $c=1$.
 We used UCB as the acquisition function where the exploration level $\lambda$ was set to 1.5.
-We used GPyTorch for GP training.
-To expose the possible gradient vanishing issue , we used Uniform(0.001, 30) over each length-scale.
-The noise variance was assigned a diffused Gamma prior, Gamma(1.1, 0.05), and the amplitude the prior Gamma(2.0, 0.15).
-We ran L-BFGS to obtain MAP estimation.
-The maximum number of iterations was 15K and tolerance level 2E-9.
-We compared with six state-of-the-art high-dimensional BO methods, including Tree UCB [10], RDUCB [33], HESBO [23], ALEBO [18], SaasBO [6], and TURBO [7].
+We used GPyTorch for GP training and BOTorch for Bayesian optimization.
+To ensure efficiency, the GP was trained via point estimation of the length-scale and noise variance parameters, with the optimizer selected from L-BFGS, Adam, or RMSProp.
+Further details are provided in Appendix Section B.3.
+We compared with six state-of-the-art high-dimensional BO methods, including Tree UCB (Han et al., 2021), RDUCB (Ziomek & Ammar, 2023), HESBO (Nayebi et al., 2019), ALEBO (Letham et al., 2020), SaasBO (Eriksson & Jankowiak, 2021), and TURBO (Eriksson et al., 2019).
 The first five have been introduced in Section 3.
@@
 TURBO searches for trust-regions in the input domain, learns a set of local GP models over each region, and then uses an implicit bandit strategy to decide which local optimization runs to continue.
-In addition, we tested the most recent method VBO (Vanilla BO[13] that levels up the performance of the standard BO by constructing a log-normal prior over the length-scale parameters (see Section 5).
+In addition, we tested the most recent method VBO (Vanilla BO(Hvarfner et al., 2024) that levels up the performance of the standard BO by constructing a log-normal prior over the length-scale parameters (see Section 5).
 We used the original implementation and default (recommended) settings for the competing methods.
@@
 The details of all the competing methods are given in Appendix Section B.3.
-For each optimization task, we randomly queried the target function for 20 times to collect the initial data.
-We tested TURBO with a single trust region and five trust regions, denoted by TURBO-1 and TURBO-5 respectively.
+For each optimization task, we randomly queried the target function 20 times to collect the initial data, except for Humanoid-Standup, where we collected 50 initial data points.
+We tested TURBO with a single trust region and five trust regions, denoted as TURBO-1 and TURBO-5 respectively.
 For HESBO and ALEBO, we varied $d_{\text{emb}}\in\{10,20\}$ and denote the choice as HESBO/ALEBO-{10, 20}.
 For SaasBO, we performed tests using both NUTS and MAP estimation for surrogate training, denoted by SaasBO-NUTS and SaasBO-MAP, respectively.
-For a reliable comparison, we ran every method for ten times.
+For a reliable comparison, we ran each method ten times, ensuring that all methods used the same randomly collected initial dataset in each run (these datasets varied across different runs).
+Additionally, we conducted ten extra runs for all the standard BO methods and VBO to further validate their performance.
+Figure 3: Optimization performance in synthetic benchmarks.
+Figure 4: Optimization performance in real-world problems.
 6.2 Optimization Performance
@@
 Due to space limit, we supplement the BNN-based BO results in Appendix Section C.1.
-Following [6], we used violin plots that illustrate the average of the obtained maximum function value over the ten runs (the middle bar), the best (the top bar), the worst (the bottom bar), and the distribution of the obtained maximum (the shaded region).
-On average, SBO-Matérn, SBO-SE(RI) and VBO consistently deliver top-tier results, often outperforming the other methods by a large margin, e.g., on benchmarks Ackley (300, 150), DNA, and SVM.
-It is noteworthy that SBO-SE also performed exceptionally well when the input dimension $d\leq 180$, as seen in benchmarks Rosenbrock (100, 100), Mopta08, NAS, and DNA.
-However, as $d$ increases, e.g., Ackley (300, 150), Hartman6 (300, 6), Rosenbrock (300, 100), StybTang(200, 200) and SVM (388), the performance of SBO-SE dropped dramatically due to the gradient vanishing issue (see our theoretical analysis and numerical experiments in Section 4).
-After applying our robust initialization, the performance of SE kernels, namely, SBO-SE (RI) immediately recovered to top-tier levels, closely matching both SBO-Matérn and VBO.
-On the other hand, though SBO-Matérn used a small length-scale initialization ($\ell_{0}=0.693$), the Matérn kernel’s structure has made it far less susceptible to the gradient vanishing issue, enabling it to handle much higher dimensions than the SE kernel.
-Free from this issue, SBO-Matérn consistently achieved top-tier results.
+We used violin plots that illustrate the average of the obtained maximum function value over the runs (the middle bar), the best (the top bar), the worst (the bottom bar), and the distribution of the obtained maximum (the shaded region).
+On average, SBO-Matérn (RI), SBO-SE (RI) and VBO consistently deliver top-tier results, often outperforming the other methods by a large margin, e.g., on benchmarks Ackley (300, 150), DNA, and SVM.
+It is noteworthy that SBO-Matérn, despite using a small length-scale initialization ($\ell_{0}=0.693$), also achieved top-tier performance across all benchmarks except for Humanoid-Standup, where the optimization dimension exceeds 1,000.
+In contrast, SBO-SE failed when the input dimension $d\geq 150$ due to the gradient vanishing issue, e.g., Ackley(150, 150), Rossenbrock(300, 100), DNA, and SVM (see our theoretical analysis and numerical experiments in Section 4).
+The result verifies that the Matérn kernel is far less susceptible to gradient vanishing and performs much more robustly for high-dimensional problems than the SE kernel.
+However, when the dimension is extremely high (e.g., over 1K), the small length-scale initialization can still cause vanishing gradients and degraded optimization performance.
+After applying our robust initialization, both kernels achieved top-tier performance consistently, closely matching VBO.
 Runtime Performance.
 Next, we examined the runtime performance.
-Fig.
-5 reports the average maximum function value obtained at each step and the standard deviation across all runs.
+Appendix Fig.
+5 and 6 report the average maximum function value obtained at each step and the standard deviation across all runs.
 Since we have 16 methods for comparison, showing the standard deviation for every method makes the figure cluttered.
 To ensure clarify, we show the curves and standard deviation of SBO, VBO, and the best, worst, and median performed remaining baselines so far at each step.
-After an initial stage, SBO-Matérn, SBO-SE (RI), and VBO consistently produced superior queries, achieving larger function values.
+After an initial stage, SBO-Matérn (RI), SBO-SE (RI), and VBO consistently produced superior queries, achieving larger function values.
+So did SBO-Matérn in all the cases except Humanoid-Standup.
 Their performance curves are generally above those of competing methods, reflecting a more efficient optimization of the target function.
 This trend is particularly noticeable in benchmarks such as Ackley (300, 150), Rosenbrock (300, 100), and SVM.
-When $d>180$, however, the performance of SBO-SE declines, with its curve positioned toward the bottom due to the gradient vanishing issue.
-In benchmarks such as Rosenbrock (100, 100) and DNA (180), the curve for VBO is predominantly above those of SBO-Matérn and SBO-SE (RI), indicating even more efficient optimization.
-However, on Ackley (150, 150) and Ackley (300, 150), VBO’s curve is mostly below that of SBO-Matérn and SBO-SE (RI).
+The performance of SBO-SE degraded when $d\geq 150$, with its curve positioned toward the bottom due to the gradient vanishing issue.
+In benchmarks such as Rosenbrock (100, 100) and DNA (180), the curve for VBO is predominantly above those of SBO-Matérn (RI) and SBO-SE (RI), indicating even more efficient optimization.
+However, on Ackley (150, 150), Ackley (300, 150) and Humanoid-Standup (1003), VBO’s curve is mostly below that of SBO-Matérn (RI) and SBO-SE (RI), and exhibits worse performance.
 For the remaining benchmarks, their curves are close and often overlap.
@@
 Additional Results.
-We conducted extensive additional evaluations and ablation studies to examine the impact of alternative GP training methods, exploration parameters in UCB, and other acquisition functions on SBO performance in high-dimensional settings.
-We found that MLE achieves optimization results comparable to the MAP training employed in our evaluation.
-However, using concentrated priors on the length-scale parameters (concentrated on small mean values) — such as the popular Gamma(3,6) prior that is the default choice in BOTorch — often leads to training failure and poor performance, aligning with our theoretical analysis.
-Both MLE and MAP perform similarly to Markov Chain Monte Carlo (MCMC) training in terms of optimization outcomes, but MCMC is significantly slower.
-UCB is notably robust to variations in the exploration parameter.
-Additionally, log-EI facilitates strong performance in high-dimensional optimization, whereas EI and TS often lead to subpar results.
+We conducted extensive additional evaluations to examine exploration parameters in UCB, and other acquisition functions on SBO performance in high-dimensional settings.
+We found that UCB is notably robust to variations in the exploration parameter.
+Additionally, log-EI facilitates strong performance in high-dimensional optimization, whereas EI and Thompson Sampling (TS) often lead to subpar results.
 Detailed results and discussions are provided in Appendix Section C.
-Figure 3: Optimization performance in synthetic benchmarks. The results were summarized from 10 runs.
-Figure 4: Optimization performance in real-world problems. The results were summarized from 10 runs.
 7 Conclusion
@@
 Our empirical evaluation demonstrates that, once this problem is mitigated, standard BO can achieve state-of-the-art performance in high-dimensional optimization.
+Acknowledgement
+This work has been supported by NSF CAREER Award IIS-2046295, NSF OAC-2311685, and Margolis Foundation.
 Appendix A Proofs
@@
 Appendix B More Experiment Details
+$\text{Fun}(d,d^{\prime})$ Structure
+Ackley(300, 150) Partial variable dependency
+Rosenbrok(300, 100) Partial variable dependency
+Nonoverlap additive decomposition
+Hartmann6(300,6) Partial variable dependency
+Rosenbrock(100, 100) Nonoverlap additive decomposition
+Stybtang(200, 200)
+Ackley(150, 150) None
+Table 3: Synthetic benchmarks and the structures within the target function: $d$ is the input dimension, and $d^{\prime}$ is the number of effective variables to compute the target function.
 B.1 Definition of Synthetic Functions
@@
 where ${\bf x}\in[-2.048,2.048]^{D}$. We set $[c_{1},...,c_{D}]$ as an evenly spaced sequence in $[-2,2]$, where $c_{1}=-2$ and $c_{D}=2$. The optimum is at $x^{\dagger}=[c_{1}+1,...,c_{D}+1]$.
-Ackley. We used the Ackley function defined in [29],
+Ackley. We used the Ackley function defined in (Surjanovic & Bingham, ),
 $f({\bf x})=-20\text{ exp}\left(-0.2\sqrt{\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}}\right)-\text{exp}\left(\frac{1}{d}\sum_{i=1}^{d}\text{cos}(2{\pi}x_{i})\right)+20+\text{exp}(1),$
@@
 $f({\bf x})=-\sum_{i=1}^{4}{\alpha}_{i}{\exp}\left(-\sum_{j=1}^{6}A_{ij}(x_{j}-P_{ij})^{2}\right),$
-where each $x_{i}\in[0,1]$, ${\bf A}=[A_{ij}]$ and ${\bf P}=[P_{ij}]$ are two pre-defined matrices as in [29]. The global optimum is at ${\bf x}^{\dagger}=[0.20169,0.150011,0.476874,0.275332,0.6573]$.
+where each $x_{i}\in[0,1]$, ${\bf A}=[A_{ij}]$ and ${\bf P}=[P_{ij}]$ are two pre-defined matrices as in (Surjanovic & Bingham, ).
+The global optimum is at ${\bf x}^{\dagger}=[0.20169,0.150011,0.476874,0.275332,0.6573]$.
 Figure 5: Runtime performance: Maximum Function Value Obtained vs.
-Number of Steps in all real-world and synthetic benchmarks.
-The results were averaged from 10 runs.
+Number of Steps in all synthetic benchmarks.
+For cleaner view, we plotted the full results only for SBO and VBO.
+For the remaining methods, we show best, median, and worst result at each step.
+Figure 6: Runtime performance: Maximum Function Value Obtained vs.
+Number of Steps in all real-world benchmarks.
 For cleaner view, we plotted the full results only for SBO and VBO.
@@
 The shaded region depicts the standard deviation.
-B.2 Implementation
+B.2 Humanoid-Standup Benchmark
+We created a novel trajectory optimization benchmark in a Humanoid Standup task based on the MuJoCo physics engine (Todorov et al., 2012).
+The action for the environment is determined by 17 parameters (corresponding to 17 motors in the humanoid).
+We set the trajectory length to 59.
+Therefore, the dimension of the problem is $59\times 17=1,003$.
+The goal is to find a trajectory $\tau=({\bf a}_{1},\ldots,{\bf a}_{59})$ of motor actions that maximize the reward, for a given initial state.
+This is an instance of optimal control and planning, which is a classical problem in reinforcement learning.
+It is noteworthy that the recent work of Hvarfner et al.
+(2024) also created a BO benchmark (named “Humanoid") based on the MuJoCo engine, but the problem setting is different.
+Their benchmark seeks to optimize a linear policy, represented by a $376\times 17$ parameter matrix that linearly maps the humanoid state to an action at each step.
+In contrast, our benchmark makes no assumptions about an underlying policy and directly optimizes the entire trajectory, making it policy-free.
+B.3 Implementation
   • ALEBO. We used ALEBO implementation shared by the Adaptive Experimentation (AX) Platform (version 0.2.2). The source code is at https://github.com/facebook/Ax/blob/main/ax/models/torch/alebo.py.
@@
 The SaasBO implementation available to the public does not include the version using MAP estimation.
-We therefore implemented this method based on the original paper [6].
+We therefore implemented this method based on the original paper (Eriksson & Jankowiak, 2021).
 All the hyperparameter settings follow exactly the same as the original paper.
@@
 […]
-B.3 Computing Environment
+  • GP Training in Standard BO.
+For efficiency, we trained GP via point estimation.
+The optimizer was chosen from L-BFGS, Adam, and RMSProp.
+For Adam and RMSProp, we set the initial learning rate to 0.1 and 0.01, respectively.
+The maximum number of epochs was set to 1,500 for both methods.
+We maximize the log marginal likelihood of the GP regression model.
+The positiveness of the length-scale and noise variance parameters are ensured via SoftPlus transform.
+For L-BFGS, we found that maximizing the log marginal likelihood often incurs numerical issues.
+To achieve numerical stability, we employed the prior Uniform(0.001, 30) over each length-scale, a diffused Gamma prior over the noise variance Gamma(1.1, 0.05), and the prior Gamma(2.0, 0.15) over the amplitude.
+The maximum number of iterations was set to 15K and tolerance level 2E-9.
+B.4 Computing Environment
 We conducted the experimental investigations on a large computer cluster equipped with Intel Cascade Lake Platinum 8268 chips.
-Mopta08 (124) Rover (60) NAS201 (30) DNA (180) SVM (388)
-ARD Matérn-MAP 7.39 9.99 7.81 9.77 13.55
-ARD Matérn-MLE 11.23 9.84 5.76 14.34 32.16
-ARD Matérn-NUTS 492.62 404.17 393.62 575.03 768.29
-SaasBO-MAP 46.53 41.36 42.35 51.33 67.37
-SaasBO-NUTS 985.76 375.06 229.07 1739.56 2269.50
-Table 4: Average running time (in seconds) per optimization step in real-world benchmark tasks.
-Rosenbrock Ackley Stybtang Rosenbrock Ackley Hartmann6
-(100,100) (150, 150) (200, 200) (300, 100) (300, 150) (300, 6)
-ARD Matérn-MAP 8.52 36.26 21.42 16.36 25.90 51.56
-ARD Matérn-MLE 8.85 12.94 13.67 12.63 9.99 8.81
-ARD Matérn-NUTS 857.21 850.23 1120.28 1276.42 1049.78 948.25
-SaasBO-MAP 70.17 70.69 66.35 100.79 85.86 85.72
-SaasBO-NUTS 1294.90 1621.60 1511.99 1938.53 925.80 2204.54
-Table 5: Average running time (in seconds) per optimization step on synthetic benchmark tasks.
 Method Structural Assumption
@@
 Table 4: BO methods and their structural assumption about the target function.
-Figure 7: Optimization performance using MLE training, MAP training with a uniform prior over the length-scale parameters, and NUTS on SVM (388) and DNA (180).
-The results were summarized from 10 runs.
-Figure 8: Optimization performance using MLE training, MAP training with a uniform prior over the length-scale parameters, and NUTS on the rest of benchmarks. The results were summarized from 10 runs.
 Appendix C Additional Results
@@
 We used the Matérn kernel.
+The comparison was performed across all the synthetic and real-world benchmarks except Humanoid-Standup (1003).
 The results are reported in Fig.
-6.
+7.
 Across all the benchmarks, the performance of the BNN based methods is worse than the median of GP based methods.
 In cases, such as SVM, Hartmann6 (300, 6) and Rosenbrock, the performance of the BNN based methods is close to or overlapping with the performance of the worst GP baselines.
-This might be due to that under low data regimes, it is more challenging for a neural network to capture the shape of the target function and quantify the uncertainty, especially for high-dimensional targets.
-Interestingly, even using a single length-scale parameter, the standard BO, namely SBO-Single — as long as it does not encounter gradient vanishing — still consistently shows superior performance, which closely matches ARD Matérn across most tasks, except for Ackley(300, 150), SBO-Single struggled in finding right locations to query, leading to poor performance.
-C.2 Alternative Ways to GP Training
-Figure 9: Final optimization performance using a diffuse prior, Uniform(0.001, 30), vs. a concentrated Gamma prior, Gamma(3,6), over the length scale parameters.
-Figure 10: Runtime optimization performance using a uniform prior, Uniform(0.001, 30), vs. a concentrated Gamma prior, Gamma(3,6), over the length scale parameters.
-C.2.1 Maximum Likelihood Estimation (MLE)
-In the main paper, we report the SBO performance when training GP via MAP estimation.
-We then examined another popular GP training approach, MLE training.
-We used ADAM optimization [17] [^2] We also experimented with using L-BFGS for MLE training.
-However, L-BFGS frequently encounters numerical issues and often fails to reliably complete model training.
-This is consistent with findings reported in [13].
-The issue likely arises because, without constraints or regularization on parameters like the length-scales and amplitude, L-BFGS tends to explore too aggressively, leading to numerical instability.
-to perform MLE, with learning rate $0.1$ and the maximum number of epochs was set to $500$.
-We employed ARD Matérn kernel.
-The ultimate and running optimization performance is shown in Fig.
-9 and 10.
-As we can see, the BO performance of using MAP and MLE training is close, and in many cases significantly overlaps, such as in DNA, SVM, Ackley (150, 150) and StybTang (200, 200).
-C.2.2 MAP Training with Alternative Priors
-Second, we examined MAP training with alternative prior choices over the length-scale parameters.
-In the main paper, we use a uniform prior, Uniform(0.001, 30).
-In practice, another commonly used choice is a concentrated Gamma prior, e.g., Gamma(3, 6), which is the default choice in the popular GP/BO libraries GPyTorch and BOTorch.
-To see how it is concentrated, in Fig.
-11 we compare its probability density curve with other priors, including the aforementioned uniform prior and the diffuse Gamma priors over the amplitude and noise variance.
-We observe that the prior distribution Gamma(3, 6) places a large probability mass around its mean 0.5, strongly regularizing the length-scale parameters toward this value.
-According to our theoretical analysis (see Section 4), this introduces a significant risk of gradient vanishing in high-dimensional settings, potentially leading to failure in GP training.
-To verify the risk, we examined the alignment between the prediction and the ground-truth across the test examples from Mopta08 (124) and DNA (180).
-As depicted in Fig.
-12 (third column), the concentrated prior leads to a severe prediction deviation from the ground-truth, indicating failure in learning the target function.
-This observation aligns with findings reported in the SaasBO paper [6] (see Fig.1 of that paper).
-We further investigated the impact of this prior within the optimization loop.
-As shown in Fig.
-9, after placing the prior Gamma(3,6) over each length-scale parameter, the BO performance drops substantially across many tasks, such as Mopta08 (124), SVM (388) and Ackley(150, 150), no matter using ADAM or LBFGS to conduct training.
-Figure 11: Probability density curves. Gamma(2, 0.15) and Gamma(1.1, 0.5) are the priors for the amplitude and noise variance, respectively.
-Figure 12: Alignment between the prediction (y-axis) and ground-truth function values (x-axis).
-C.2.3 Fully Bayesian Training
-Third, we investigated whether a fully Bayesian training approach would yield substantial improvements in BO.
-In particular, we can use MCMC to perform posterior inference of the GP parameters, and then integrate them out during the prediction.
-We conducted another round of BO experiments.
-For each step, we used NUTS to train GP with ARD Matérn.
-We adopted the same settings as used in SaasBO-NUTS [6]: where the mean parameter is assigned the standard Gaussian prior $\mathcal{N}(0,1)$, amplitude $\text{Gamma}(2,0.15)$, and noise variance $\text{Gamma}(0.9,10)$.
-We then assigned a uniform prior over each length-scale, $\text{Uniform}(0.001,30)$.
-The number of warm up steps was set to 512, and after the warm up stage, we ran 256 steps, and selected the samples at every 16-th step as the posterior samples.
-We show the running and final optimization performance in Fig 7 , and Fig.
-8.
-We also show the average running time per optimization step in Table 4 and Table 5.
-It can be observed that while NUTS contributes to a slight improvement in several tasks, the enhancement is marginal.
-However, running NUTS is much more computationally costly compared to doing MAP and MLE.
-For example, on DNA, NUTS takes 59x and 40x running time of MAP and MLE, respectively, on average in each step.
-Therefore, for practical purposes, using MAP or MLE for GP training appears sufficient.
-It is computationally efficient, and the sacrifice of optimization performance is minimal compared to much more costly training methods like NUTS.
-C.3 Different Choice of $\lambda$ in UCB
-Figure 13: Runtime optimization performance with different exploration level $\lambda$ in UCB.
+This might be due to that under low data regimes, it is more challenging for a neural network to capture the landscape of the target function and quantify the uncertainty, especially for high-dimensional targets.
+Interestingly, even using a single length-scale parameter, the standard BO, namely SBO-Single — as long as it does not encounter gradient vanishing — still consistently shows superior performance, which often closely matches ARD Matérn, except for Ackley(300, 150), SBO-Single struggled in finding right locations to query, leading to poor performance.
+C.2 Different Choice of $\lambda$ in UCB
+Figure 8: Runtime optimization performance with different exploration level $\lambda$ in UCB. In the main paper and Figure 5 and 6, SBO was compared with other methods using $\lambda=1.5$.
 We examined how the exploration parameter $\lambda$ in UCB influences the BO performance in high-dimensional optimization.
 To this end, we used the Maérn kernel and varied $\lambda$ from {0.5, 1.0, 1.5, 2.0}.
-We conducted MLE training via ADAM optimization.
-We show the ultimate and runtime optimization performance in Fig.
-13, respectively.
+We show the runtime optimization performance in Figure 8.
 It can be seen that, in most benchmarks (Mopta08, Rover, NAS201, DNA, etc.), the performance with different $\lambda$’s is close.
-However, on DNA, SVM and Ackley, the performance of $\lambda=0.5$ is typically worse than the bigger choice of $\lambda$.
+However, on SVM and Ackley, the performance of $\lambda=0.5$ is typically worse than the bigger choice of $\lambda$.
 This might be due to that those problems are more challenging (e.g., Ackley has many local minima), a larger exploration level ($\lambda$) is needed to identify better optimization outcomes.
-C.4 Alternative Acquisition Functions
+C.3 Alternative Acquisition Functions
 Figure 9: Final performance with different acquisition functions.
@@
 We used BOTorch to maximize EI through L-BFGS.
-The most recent work [1] introduced log-EI to overcome the numerical challenges in optimizing the original EI acquisition function.
+The most recent work (Ament et al., 2024) introduced log-EI to overcome the numerical challenges in optimizing the original EI acquisition function.
 Accordingly, we also tested log-EI using the BoTorch implementation.
-To conduct TS at each step, we used Sobol Sequence [28] to sample 3K inputs in the domain, and jointly sampled the 3K function values from the GP posterior.
+To conduct TS at each step, we used Sobol Sequence (Sobol’, 1967) to sample 3K inputs in the domain, and jointly sampled the 3K function values from the GP posterior.
 We then selected the input with the largest function value as the next query.
@@
 We used BOTorch implementation of both methods.
-We used the Matérn kernel and MLE training at each step.
+We used the Matérn kernel.
 The ultimate and runtime optimization performance for using EI, TS-3K and TS-10K, are shown in Fig.
-14 and 15, respectively.
+9 and 10, respectively.
 As we can see, in most cases, such as Mopta08 (128), SVM (388), DNA (180) and Rosenbrock (100, 100), both EI and TS perform worse than UCB, and also worse than the best baseline.
 On Ackley(150, 150) and Ackley(300, 150), the performance of EI and TS are even among the worst.
-The failure of EI might arise from the gradient vanishing during the optimization, as pointed by the recent work [1].
+The failure of EI might arise from the gradient vanishing during the optimization, as pointed by the recent work (Ament et al., 2024).
 The failure of the TS might be due to the high-dimension of the input space.
@@
 The use of log-EI performed excellently, achieving results comparable to those obtained with UCB.
-Appendix D Limitation
+Appendix D Length-scale Initialization Strategies Across Different Gaussian Process Libraries
+In this section, we have investigated the initialization strategy of a (non-exhaustive) list of GP training libraries.
+Most libraries employ a small constant initialization while a few others created the initialization based on the distances between the training data points.
+D.1 Constant Initialization
+D.2 Data-Dependent Initialization
+  • DiceKriging [^10] https://github.com/cran/DiceKriging initializes the length-scale randomly based on the range of the input features.
+The initial length scale for each feature is sampled from a uniform distribution, with a lower bound of $1\times 10^{-10}$ and an upper bound set to twice the maximum difference of the variable values in the training dataset.
+  • hetGP [^11] https://github.com/cran/hetGP initializes the length-scale to the 50th percentile of the pairwise distances between the training data points.
+  • GPVecchia [^12] https://github.com/katzfuss-group/GPvecchia initializes the length-scale to be proportional to the mean of the pairwise distances between the training data points.
+Appendix E Limitation
 Our current work has identified gradient vanishing as a critical failure mode of SBO in high-dimensional settings.
@@
 To answer these questions, we will continue our investigation in future work, both theoretically and empirically.
+In addition, we empirically found that the choice of optimizer can also impact GP training performance in high-dimensional settings.
+For instance, L-BFGS tends to be less stable than ADAM, while RMSProp can outperform ADAM in certain cases.
+In future work, we will continue exploring initialization strategies to enhance the robustness of GP training across different optimizers.